### v2

# Fighting Exponentially Small Gaps by Counterdiabatic Driving

Abstract
We investigate the efficiency of approximate counterdiabatic driving (CD) in accelerating adiabatic passage through a first-order quantum phase transition. First, we analyze a minimal spin-glass bottleneck model that is analytically tractable and exhibits both an exponentially small gap at the transition point and a change in the ground state that involves a macroscopic rearrangement of spins. Using the variational Floquet-Krylov expansion to construct CD terms, we find that while the formation of excitations is significantly suppressed, achieving a fully adiabatic evolution remains challenging. Extending our investigation to realistic NP-hard spin-glass problems—specifically, the $3$-regular Max Cut and $3$-XORSAT—we find again that local CD expansions lead to negligible improvements in the final ground state fidelity. These results highlight the limited impact of local CD methods in overcoming the bottlenecks associated with first-order quantum phase transitions. To address this limitation, we propose an alternative method, termed quantum brachistochrone counterdiabatic driving (QBCD), which employs the approximate full CD connecting the ground state and the first excited state at a single parameter value close to the critical point. In the minimal spin-glass model, QBCD enables exponentially faster adiabatic evolution than the local strategies. To bring experimental realization within reach for realistic NP-hard problems, we further enhance the QBCD Hamiltonian by sparsifying it to the density of the quadratic local expansion. This reduces the non-locality exponentially while preserving finite ground-state fidelity at driving times exponentially shorter than in the local strategies and in the counterdiabatic optimized local driving (COLD). These findings highlight that targeting a single avoided crossing with an exponentially reduced counterdiabatic correction based on approximate numerical knowledge is sufficient to bypass spin-glass bottlenecks, offering a compelling path forward for near-term quantum devices.
I Introduction
The breaking of adiabatic dynamics in many-body spin systems arises in a wide variety of scenarios ranging from condensed matter physics to quantum simulation and quantum computing. According to the adiabatic theorem, the driving of a quantum system in a time shorter than the inverse of the spectral gaps generally induces diabatic transitions. Nonadiabatic dynamics limit the preparation of novel ground-state phases of matter in quantum simulation. Excitations further account for errors in adiabatic quantum computation, quantum annealing, and quantum optimization algorithms. They are also responsible for quantum friction in finite-time thermodynamics, limiting the scaling of quantum heat engines and refrigerators, among other quantum devices.
Enforcing adiabaticity in many-body quantum systems is thus broadly desirable. Yet, it faces the need to overcome small gaps in the spectrum of the driven system. By way of example, hard instances of combinatorial problems might be difficult to solve by quantum annealers because of the presence of gaps on the driving path that are exponentially small in the system size [1, *Young2010First, 3, 4, 5, *Altshuler2010Anderson]. Such gaps render quantum adiabatic evolution unfeasible in practice since the adiabatic theorem requires the driving time $T$ to be larger than the inverse square of the minimal gap $\Delta_{\min}$ [7, 8]. The physical origin of these gaps remains a topic of active debate. Some authors have used perturbation theory to link them to first-order quantum phase transitions [3, 4], suggesting they arise from Anderson localization  [5, 6, 9]. Others argue that the gaps have a non-perturbative origin [10] or result from the clustering of solutions typical in spin-glass phases [11]. Despite the debate, large-scale numerical simulations have consistently demonstrated that certain models display exponentially small gaps [1, *Young2010First, 12, 13].
Facing the presence of small gaps and their dependence on the system size has led to reformulations of the adiabatic theorem in a many-body setting [14, 15], and the notion of quasi-adiabatic continuation used to determine the existence and stability of phases of matter [16, *Hastings2007Quasi, *Hastings2010Locality, *Hastings2010Quasi, 20, 21, 22].
These efforts have been accompanied by a closely related quest for fast quantum control of many-body systems, tackling the presence of small gaps on the adiabatic path. In the context of adiabatic quantum computing, tailoring the mixer or problem Hamiltonian has been proposed to remove the exponential gaps from the path [23, 24, *Choi2011Different, *Choi2011Different2, 27, 28, 29]. Alternative methods involve additional “catalyst” Hamiltonians, i.e., operators that act only at intermediate times to amplify the gap [30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. Other strategies rely on time discretization combined with the use of classical optimization, leading to the quantum approximate optimization algorithm (QAOA) [40, 41], and the use of quantum optimal control to tailor non-linear parameter driving [42, 43, 44, 45, 46].
The problem of speeding up the adiabatic evolution was also addressed in other contexts, such as the application of quantum control in chemistry and physics [47]. Seminal works [48, *Demirplak2005Assisted, *Demirplak2008Consistency, 51] showed that adiabatic evolution for a generic Hamiltonian can be achieved in any finite time, provided an auxiliary counterdiabatic driving (CD) control term is implemented. The CD term is constructed in such a way as to cancel all diabatic transitions, realizing exactly the adiabatic evolution of the reference uncontrolled Hamiltonian. While its explicit form is familiar from early proofs of the adiabatic theorem [52] and can be obtained directly given the spectral decomposition of the instantaneous Hamiltonian, the CD Hamiltonian is generically non-local and involves multiple-body interaction terms  [53, 54, 55, 56]. However, it can also be obtained variationally via controlled expansions in increasingly non-local terms [56, 57, 58, 59, 60, 61], with exact adiabatic dynamics being recovered when all the terms in the expansion are retained. As a result, contrary to catalysts and ad-hoc deformations of the driving Hamiltonian, CD provides a coherent framework to address the problem of exponentially small gaps. CD has been shown to boost the performance of QAOA, motivating its use in quantum optimization  [62, 63, 64, 65, 66, 67, 68]. It should be noted that the required CD terms cannot generally be realized in analog quantum annealers and simulators. However, they can be implemented by digital [53, 56, 62, 63, 66, 69, 70, 67, 68] or hybrid digital-analog quantum simulation [71, 72]. Within such approaches, the implementation is generally facilitated by the locality of the interactions.
Despite a surge of interest in harnessing CD for many-body interacting systems [53, 54, 56, 55, 73, 74, 58, 57, 59, 75, 76, 62, 63, 64, 65, 66, 77, 60, 61, 78, 67], crucial aspects remain to be elucidated. Early studies pointed out the interplay between CD and the Kibble-Zurek (KZ) scaling at second-order quantum phase transitions, which describes the breakdown of adiabaticity while crossing critical point in terms of universal equilibrium critical exponents [79, 80, 81, 82, 83, 84]. It was found that non-local CD terms are needed to suppress the KZ scaling and restore adiabaticity [53, 84, 55], which is at odds with the quest for approximate CD terms with locality tailored for implementations. The breaking of adiabaticity described by KZ mechanism in first-order phase transitions has been theoretically and experimentally studied [85, 86], while the use of bias fields in digitized quantum optimization assisted by CD has only recently been put forward [87]. The performance of CD in a problem with an exponentially small gap and frustration is yet to be elucidated. Some works [88, 89] have addressed the CD-assisted annealing of ferromagnetic $p$-spin models, showing that CD improves the final fidelity while still remaining exponentially small in the system size. The effect of frustration, however, being at the basis of the exponential difficulty of solving NP-hard problems, was not considered. Furthermore, while in certain $p$-spin models, simple antiferromagnetic catalysts can yield a polynomial scaling of the gap [37], this cannot happen if the system is frustrated [39].
Several vital questions thus arise: How efficiently can approximate counterdiabatic (CD) driving improve adiabatic passage through exponentially small gaps between frustrated energy levels? Can alternative strategies be developed to overcome these regimes more effectively? Most importantly, what additional, minimal information beyond local terms is required to achieve exponential speedups in adiabaticity? And how far can the required non-locality and the evaluation range of parameters be restricted?
In this work, we investigate whether local, approximate counterdiabatic (CD) driving can effectively suppress excitations and accelerate adiabatic dynamics in many-body quantum systems exhibiting exponentially small gaps and frustration. To this end, we evaluate the performance using three experimentally relevant metrics: the density of non-adiabatic excitations, the ground-state gap, and the final ground-state fidelity[90, 91, 92, 93, 94, 95, 96, 97, 98]. We find that while local approximate CD can significantly reduce excitations away from the adiabatic limit, its impact diminishes near the adiabatic timescale. To address this limitation, we introduce quantum brachistochrone counterdiabatic driving (QBCD), which yields a remarkable improvement by relying solely on approximate knowledge of the ground and first excited states at a single point in parameter space. We further validate these findings on two realistic NP-hard problems, demonstrating the superiority of an advanced, sparsified QBCD method—featuring exponentially reduced non-locality—over both the local expansion schemes and the counterdiabatic local optimized driving (COLD) approach.
I.1 Summary of results
To address the questions above, we study a recently introduced minimal model of a spin-glass bottleneck [99]. The model is an Ising-like spin chain that displays both a small gap and frustration: not only does the gap to the first excited state scale exponentially with the system size $L$, but also $O(L)$ spins need to be flipped at the avoided crossing to track adiabatically the ground state. These features are due to a simple localization mechanism reminiscent of the formation of bound states for delta-function potentials on the real line. The model can be mapped onto a chain of free fermions, which allows a complete analytical understanding of the small gap formation [99] and the simulation of large systems to access the asymptotic scaling. Although this model exhibits most of the essential features of realistic spin glass problems, we further extend the investigation to two realistic NP-hard problems, the Max Cut and the $3$-XORSAT problems [2], in which spin glass behavior and an exponentially small gap emerges for numerically accessible system sizes, $L\gtrsim 10$.
First of all, we present a thorough analysis of the model properties in Sec. II. While most properties were already derived in the original reference [99], we elucidate the role of a chiral symmetry that was previously mistaken for a regular symmetry and provide a detailed derivation of the free-fermionic representation of the model. Then, after briefly reviewing counterdiabatic driving (CD) in Sec. III.1, we compute the free-fermionic representation of approximate CD for the model under consideration in Sec. III.2.
In Sec. IV, we present our main results for the minimal spin glass model. We show that local, approximate CD terms obtained without prior knowledge of the spectrum provide reasonable speed-up only for a given regime of driving times. We do so with two complementary approaches. First, we consider the effect of CD terms on the minimal gap along the adiabatic path, i.e.,  we study the performance of CD as a gap amplification technique. Since the CD Hamiltonian depends on the total evolution time, defining a gap requires care. A solution is presented in Sec. IV.1, showing that if the problem Hamiltonian displays a minimal gap $\Delta_{\min}\sim e^{-\alpha L}$, then adding CD terms leads to $\Delta_{\min{},\mathrm{CD}}\sim e^{-\alpha_{\mathrm{CD}}L}$ with $\alpha_{\mathrm{CD}}$ being smaller than $\alpha$ but remaining finite. Therefore, while CD accelerates the driving protocol exponentially, the scaling remains exponentially hard. Second, in Sec. IV.2, we confirm that CD does not help the system overcome the smallest gap by directly simulating the time evolution of the system. We consider the number of kinks at the end of the schedule, displaying relevant suppression only for the short-time evolutions. Meanwhile, significant improvement for slower schedules can only be achieved for a sufficiently large number of expansion terms due to the macroscopic rearrangement of the ground states around the critical point. In the language of free fermions, local CD maps to local fermion hopping terms, while a non-local population transfer is needed to track adiabatically the ground state.
In Sec. V, we introduce a novel approach to the bottleneck problem by attacking directly the first-order phase transition. The QBCD Hamiltonian is constructed using approximate knowledge of the eigenstates at a single parameter value at the transition. In the minimal spin glass model, the QCDB can be understood as driving the localized edge states associated with the minimal gap. We show that this approach achieves exponentially shorter adiabatic time scales by exploiting only an exponentially small fraction of the full counterdiabatic Hamiltonian. Combined with its system-size-independent energetic cost, these features support near-term experimental implementation.
We then test the generality of our findings in two realistic spin glass models representing NP-hard optimization problems, the Max Cut and the $3$-XORSAT. First, we demonstrate the same limited efficiency of the first and second-order local expansion via the ground state gap and the final fidelity. Next, we introduce a yet further enhanced version of the QBCD Hamiltonian, by sparsifying it to the density of the first order local expansion. Remarkably, this approach with an exponentially reduced non-locality provides the same level of shortening of the adiabatic time-scale as the full one. These features significantly ease the challenges of implementation, making the approach promising for both experimental platforms and large-scale classical simulations. Finally, we also test the efficiency of the COLD approach and find that while it offers a noticeable improvement over local expansion at small system sizes, this advantage gradually decreases as $L$ increases. We show that the sparsified QBCD clearly outperforms the COLD protocol. In addition, the numerical implementation of the COLD protocol is orders of magnitude more demanding than that of the other approaches, restricting its applicability to very small system sizes ($L\lesssim 10$).
A closing discussion in relation to previous findings follows in Sec. VI.
II Minimal model of spin-glass bottleneck
Consider the quantum many-body Hamiltonian [99]
$H[\lambda]=-(1-\lambda)\sum_{j=1}^{L}\sigma_{j}^{x}-\lambda\sum_{j=1}^{L}J_{j}\sigma_{j}^{z}\sigma_{j+1}^{z},$ (1)
where $\lambda\in[0,1]$ is the driving parameter, and the chain has an odd number of sites $L=2\ell+1$, with a periodic boundary. We use square brackets to denote the dependence on the parameter $\lambda$, while we use parentheses for the other functional dependencies, e.g., of $\lambda(t)$ on time. The couplings are set to
$J_{j}=\begin{cases}J&j=\ell,\ell+1\\ -J^{\prime}&j=2\ell+1\\ 1&\text{otherwise,}\end{cases}$ (2)
where it is assumed
$0<J^{\prime}<J<1,\qquad J^{2}<J^{\prime}.$ (3)
From the pictorial representation in Fig. 1a, it can be seen that Eq. (1) describes an Ising chain with uniform ferromagnetic couplings, except for three weaker couplings placed in diametrically opposite positions: two weak ferromagnetic ones ($J$) next to each other, and one weaker antiferromagnetic one ($-J^{\prime}$) on the other side. The condition $J^{2}<J^{\prime}$ in Eq. (3) is more technical, and is explained in App. B.4.
The Hamiltonian in Eq. (1) has $\mathbb{Z}_{2}$ fermion parity symmetry (the fermionic nature of the parity will become clear in Sec. II.2), due to the possibility of flipping all the spins along $z$:
$\Pi_{\mathrm{F}}=\prod_{j=1}^{L}\sigma_{j}^{x},\qquad\big{[}\Pi_{\mathrm{F}},H[\lambda]\big{]}=0.$ (4)
In the following, it will be convenient to restrict to the even parity sector, which is the one dynamically accessible starting from the ground state at $\lambda=0$.
The model has also $\mathbb{Z}_{2}$ reflection parity symmetry $\Pi_{\mathrm{R}}$, defined by the action
$\Pi_{\mathrm{R}}\sigma_{j}^{\alpha}\Pi_{\mathrm{R}}=\sigma_{L-j+1}^{\alpha},\qquad\alpha=x,y,z,$ (5)
and highlighted with a grey dashed line in Figs. 1a and 2b. One can check that $\big{[}\Pi_{\mathrm{R}},H[\lambda]\big{]}=0$. It will prove convenient to use a modified version of this symmetry to exploit the free fermionic nature of the chain (see Sec. II.2).
Figure 1: (a) Ising chain of Eqs. (1)–(2). The chain is $L=2\ell+1$ sites long, with uniform ferromagnetic bonds (green lines), except for two neighboring, weakly ferromagnetic bonds (light green lines) and a single antiferromagnetic bond (red line). The dashed gray line indicates the presence of the $\mathbb{Z}_{2}$ reflection symmetries, Eqs. (5) and (10). (b) The ground state at $\lambda=1$ (fixing $L=9$) has only the antiferromagnetic bond frustrated (dotted line). The other degenerate ground state with opposite fermion parity is not shown. (c)–(d) Two degenerate first excited states at $\lambda=1$ (with $L=9$), characterized by only one weak ferromagnetic bond frustrated (dotted line). Two other degenerate first excited states are obtained by flipping all the spins. (e) The restriction to the even fermion parity sector makes the ground state nondegenerate and the first excited manifold two-dimensional. The positive fermion parity states are obtained by means of symmetric combinations, as shown here for the ground state.
II.1 Simple limits
In order to gain an understanding of the model, we analyze the lowest energy states at the beginning ($\lambda=0$) and at the end ($\lambda=1$) of the driving schedule. When $\lambda=0$, the ground state is simply the product state of spins pointing along the $x$ direction, |$\ket{E_{0}[\lambda=0]}=\prod_{j}(\ket{\uparrow}_{j}+\ket{\downarrow}_{j})/\sqrt{2}$. It has an even fermionic parity: $\Pi_{\mathrm{F}}\ket{E_{0}[\lambda=0]}=\ket{E_{0}[\lambda=0]}$, since the state on each site is even itself. The $L$ degenerate first-excited states are obtained by flipping one of the spins.
At the end of the driving schedule ($\lambda=1$), because of the presence of a single antiferromagnetic bond in a closed loop geometry, the ground state is frustrated. The condition $10<J^{\prime}<J<$ 1, Eq. (3), implies that the frustrated bond is at $j=2\ell+1$, see Fig. 1b, and thus the ground-state energy is $E_{0}=-(L-3)-2J+J^{\prime}$. Moreover, because of the fermion parity $\Pi_{\mathrm{F}}$, the ground state is doubly degenerate: the degeneracy is removed by restricting to the even fermion parity sector and taking the symmetric combination |$\ket{E_{0}^{+}}$ as in Fig. 1e.
Also the four degenerate first-excited states are easy to determine: they display a single frustrated bond, i.e.,  the weak ferromagnetic link $j=\ell$ or $\ell+1$, and have energies $E_{1}=-(L-3)-J^{\prime}$; see Fig. 1c–d for a pictorial representation (the two other degenerate states are obtained by flipping all the spins). The restriction to the even parity sector implies that symmetric combinations need to be taken, similarly to what is shown in Fig. 1e for the ground state.
From the exact computation reported below, it turns out that at an intermediate value of $\lambda$ the ground state undergoes an exponentially small avoided crossing with the first excited state. In particular, the ground state changes from a dressed version of the states in Fig. 1c–d ($\lambda<\lambda_{c}$) to a dressed version of the state in Fig. 1b ($\lambda>\lambda_{c}$). The ground states before and after the avoided crossing are macroscopically different, as the flipping of half of the chain is involved. Therefore, the Hamiltonian  (1) features a first-order transition similar to the ones that take place in spin-glass models.
The rest of this section is devoted to the analytical description of the avoided crossing of the model. While the computation was already performed in the original work [99], some steps were either missing or wrong (while the final results are substantially correct). For this reason, and to make the manuscript self-contained, we detail the solution and the salient features of the model.
II.2 Fermionic representation and reflection parity symmetry
By means of a Jordan-Wigner transformation, the spin chain Eq. (1) can be fermionized into a chain of Dirac fermions, as detailed in App. A.1. The fermion parity operator $\Pi_{\mathrm{F}}$ in Eq. (4) is transformed into
$\Pi_{\mathrm{F}}=\prod_{j=1}^{L}\left(1-2c_{j}^{\dagger}c_{j}^{\phantom{\dagger}}\right),$ (6)
which counts the parity of the total number of fermions. Then, the Hamiltonian is transformed into
$H[\lambda]=\sum_{j=1}^{L}\left[-\lambda\tilde{J}_{j}\left(c_{j}^{\dagger}-c_{j}^{\phantom{\dagger}}\right)\left(c_{j+1}^{\dagger}+c_{j+1}^{\phantom{\dagger}}\right)\right.\\ \left.+2(1-\lambda)c_{j}^{\dagger}c_{j}^{\phantom{\dagger}}\right]-(1-\lambda)L,$ (7)
where the couplings $\tilde{J}_{j}$ are the same as $J_{j}$, except that the sign of the last coupling is flipped: $\tilde{J}_{L}=-J_{L}$. This is to take into account the even fermionic parity of the ground state sector; see App. A.1 for a thorough discussion.
Because of the particle non-conservation entailed by the Dirac Hamiltonian (7), it is convenient to pass to the Bogoliubov basis, where excitations are conserved in the adiabatic limit. However, the matrix representation of Eq. (7) is not analytically diagonalizable to the best of our knowledge (see also App. E.1). It is thus convenient to perform two additional transformations [99]. First, each Dirac fermion is split into a couple of Majorana fermions, as illustrated in Fig. 2a, according to:
$\gamma_{2j-1}:=\frac{1}{\sqrt{2}}\left(c_{j}^{\phantom{\dagger}}+c_{j}^{\dagger}\right),\quad\gamma_{2j}:=\frac{i}{\sqrt{2}}\left(c_{j}^{\phantom{\dagger}}-c_{j}^{\dagger}\right),$ (8)
that obey commutation relations $\{\gamma_{k},\gamma_{k^{\prime}}\}=\delta_{k,k^{\prime}}$. The Hamiltonian acquires the form
$H[\lambda]=-2i\sum_{j=1}^{L}\big{[}\lambda\tilde{J}_{j}\gamma_{2j}\gamma_{2j+1}+(1-\lambda)\gamma_{2j-1}\gamma_{2j}\big{]},$ (9)
see also Fig. 2a and App. A.2. Second, it is convenient to define a modified reflection symmetry $\widetilde{\Pi}_{\mathrm{R}}$, instead of $\Pi_{\mathrm{R}}$ in Eq. (5) (for a detailed explanation see App. A.3). Defining $\widetilde{\Pi}_{\mathrm{R}}$ by the action
$\widetilde{\Pi}_{\mathrm{R}}\gamma_{k}\widetilde{\Pi}_{\mathrm{R}}=\gamma_{2L-k+1},$ (10)
one can check that $\widetilde{\Pi}_{\mathrm{R}}^{2}=\mathbb{I}$ and
$\widetilde{\Pi}_{\mathrm{R}}H[\lambda]\widetilde{\Pi}_{\mathrm{R}}=-H[\lambda],$ (11)
i.e., that $\widetilde{\Pi}_{\mathrm{R}}$ is a chiral symmetry for $H[\lambda]$. This means that $H[\lambda]$ is not split into sectors and does not acquire a block diagonal form, but simply the spectrum of $H[\lambda]$ is symmetric with respect to 00, with eigenstates coming in pairs [100].
Figure 2: (a) Majorana representation of the chain for $L=5$. Each Dirac site is doubled into two Majorana sites, according to Eq. (8), and thus the length of the Majorana chain becomes $2L$ (i.e., $10$ in the picture). Notice that the sign of the coupling $J_{L}$ between the first and last Majorana fermions is flipped to take into account the even fermionic parity; the bond is dotted in the figure to highlight the fact. (b) Symmetrized Dirac representation of the chain. Each Majorana operator $\gamma_{k}$ is coupled with its symmetric partner $\gamma_{2L-k+1}$, according to Eqs. (12): in the picture, this is shown as a color-code matching of the sites. The chain goes back to a length of $L$ because the sites to the right of the symmetry cut (dashed gray line) become redundant.
The presence of the chiral symmetry $\widetilde{\Pi}_{\mathrm{R}}$ suggests to define the symmetrized operators illustrated in Fig. 2b,
$\Gamma^{\pm}_{k}$ =±$=\frac{\pm i\gamma_{k}+\gamma_{2L-k+1}}{\sqrt{2}},\quad$ $k=1,3,\dots,L,$ (12a)
$\Gamma^{\pm}_{k}$ $=\frac{\gamma_{k}\pm i\gamma_{2L-k+1}}{\sqrt{2}},\quad$ $k=2,4,\dots,L-1,$ (12b)
where the factors ±$\pm i$ where chosen so that
$\{\Gamma_{k}^{+},\Gamma_{k^{\prime}}^{-}\}=\delta_{k,k^{\prime}},\quad\{\Gamma_{k}^{+},\Gamma_{k^{\prime}}^{+}\}=\{\Gamma_{k}^{-},\Gamma_{k^{\prime}}^{-}\}=0,$ (13a)
$\widetilde{\Pi}_{\mathrm{R}}\Gamma_{k}^{\pm}\widetilde{\Pi}_{\mathrm{R}}=\pm i\Gamma_{k}^{\mp},\quad(\Gamma^{\pm}_{k})^{\dagger}=\Gamma^{\mp}_{k}.$ (13b)
The equations above show that $\Gamma_{k}^{\pm}$ represent a new set of Dirac fermions, in terms of which the Hamiltonian takes the convenient form
$H[\lambda]=$ $2\lambda\sum_{j=1}^{\ell}J_{j}\left(\Gamma_{2j}^{+}\Gamma_{2j+1}^{-}+\Gamma_{2j+1}^{+}\Gamma_{2j}^{-}\right)$ (14)
$-2(1-\lambda)\sum_{j=1}^{\ell}\left(\Gamma_{2j-1}^{+}\Gamma_{2j}^{-}+\Gamma_{2j}^{+}\Gamma_{2j-1}^{-}\right)$
$-\lambda J_{L}\left(2\Gamma_{1}^{+}\Gamma_{1}^{-}-1\right)-(1-\lambda)\left(2\Gamma_{L}^{+}\Gamma_{L}^{-}-1\right),$
as shown in App. A.4. Above, the $J_{j}$’s have been restored since the last coupling is isolated from the others, and there is no need to keep the $\tilde{J}_{j}$’s. One can check again that $\widetilde{\Pi}_{\mathrm{R}}H[\lambda]\widetilde{\Pi}_{\mathrm{R}}=-H[\lambda]$, as follows immediately from Eqs. (13a)–(13b). Introducing the vectors
$\vec{\Gamma}^{+}=\begin{pmatrix}\Gamma^{+}_{1}&\Gamma^{+}_{2}&\dots&\Gamma^{+}_{L}\end{pmatrix},\qquad\vec{\Gamma}^{-}=\big{(}\vec{\Gamma}^{+}\big{)}^{\dagger},$ (15)
and the effective single-particle Hamiltonian
$\mathcal{H}[\lambda]=\begin{pmatrix}-\lambda J_{L}&\lambda-1&&&&&\\ \lambda-1&0&\lambda J_{1}&&&&\\ &\lambda J_{1}&0&\lambda-1&&&\\ &&\lambda-1&0&\lambda J_{2}&&\\ &&&\lambda J_{2}&\ddots&\ddots&\\ &&&&\ddots&0&\lambda J_{\ell}\\ &&&&&\lambda J_{\ell}&\lambda-1\end{pmatrix},$ (16)
the matrix representation of the Hamiltonian (14) becomes
$H[\lambda]=2\,\vec{\Gamma}^{+}\mathcal{H}[\lambda]\vec{\Gamma}^{-}+\lambda J_{L}+(1-\lambda).$ (17)
From the explicit matrix above, one can see that the diagonalization of $H[\lambda]$ has been reduced to the diagonalization of a simpler one-particle hopping model with open boundary conditions, described by the matrix $\mathcal{H}[\lambda]$. Here and throughout the rest of the paper, we use calligraphic letters to refer to the operators acting on the linear space of the effective Hamiltonian $\mathcal{H}[\lambda]$, and a modified ket notation $|\psi)$ for the corresponding single-particle modes.
II.3 Description of the avoided crossing
The matrix $\mathcal{H}[\lambda]$ in Eq. (16) cannot be diagonalized analytically for general $J_{j}$’s. However, using the $J_{j}$’s given in Eq. (2), the entire spectrum can be obtained with a Fourier ansatz: the effective hopping model is translation invariant in the bulk. Thus, the eigenstates must have a plane (or evanescent) wave structure in the bulk, and the effect of the inhomogeneous boundaries can be taken into account by simply fixing the boundary conditions. This is reminiscent of the solution of the Schrödinger equation on a line with a piecewise constant (or delta-function) potential. The computation is carried on in detail in App. B, while here, only the main features are stated.
Figure 3: (a) Sketch of the boundary-localized eigenmodes $|\psi_{\mathrm{L}})$,$|\psi_{\mathrm{R}})$ of the effective hopping Hamiltonian (16). (b) Spectrum of the hopping Hamiltonian (16) for $\ell=8$, $J=0.5$, $J^{\prime}=0.27$. The many-body ground state of the $\Gamma$ Hamiltonian in Eq. (14) is formed by populating all the negative-energy modes at $\lambda=0$ (solid lines) and leaving the positive-energy modes empty (dashed lines). The two localized modes $\epsilon_{\mathrm{L}}$ (blue) and $\epsilon_{\mathrm{R}}$ (orange) undergo an avoided crossing at $\lambda=\lambda_{\mathrm{c}}$: while $\epsilon_{\mathrm{R}}$ contributes to the ground state for $\lambda<\lambda_{\mathrm{c}}$ (continuous portion of the orange line), it is replaced by $\epsilon_{\mathrm{L}}$ for $\lambda>\lambda_{\mathrm{c}}$ (the continuous portion of the blue line). Since the spatial overlap between $|\psi_{\mathrm{L}})$ and $|\psi_{\mathrm{R}})$ is exponentially small in $L$, also the gap $\Delta_{\mathrm{min}}$ remains exponentially small according to Eq. (19).
The spectrum of the effective hopping model, Eq. (16), is essentially composed of plane wave eigenmodes. Under the condition $J^{2}<J^{\prime}$. However, a couple of boundary modes are present; see also Eq. (107): one, call it $|\psi_{\mathrm{L}})$, is localized around the left end $j=0$ of the $\Gamma$ chain, and a second one, denoted by $|\psi_{\mathrm{R}})$, extends around the right end $j=L$. Their spatial decay is regulated by the rates $\kappa_{\mathrm{L}}$ and $\kappa_{\mathrm{R}}$ as
$\psi_{\mathrm{L},j}\sim e^{-\kappa_{\mathrm{L}}j/2},\qquad\psi_{\mathrm{R},j}\sim e^{-\kappa_{\mathrm{R}}(L-j)/2};$ (18)
see Fig. 3a for a sketch. The corresponding eigenvalues $\epsilon_{\mathrm{L}}$ and $\epsilon_{\mathrm{R}}$ undergo an avoided crossing at a value of the driving parameter $\lambda=\lambda_{\mathrm{c}}$, with a gap
$\Delta_{\mathrm{min}}\propto e^{-\alpha L},\qquad\alpha=\frac{1}{2}\ln\left[\frac{J^{\prime}(1-J^{2})}{J^{2}-(J^{\prime})^{2}}\right].$ (19)
Notice that $\alpha>0$, thanks to the conditions Eq. (3). The exponentially small gap $\Delta_{\mathrm{min}}$ hinders adiabatic driving because the many-body ground state of the spin chain contains the single-particle orbital $|\psi_{\mathrm{R}})$ for $\lambda\lesssim\lambda_{\mathrm{c}}$, while it contains the single-particle orbital $|\psi_{\mathrm{L}})$ for $\lambda\gtrsim\lambda_{\mathrm{c}}$, see Fig. 3b. Therefore, the minimal gap along the adiabatic path is equal to $\Delta_{\mathrm{min}}$. At the avoided crossing, it is necessary to bring an effective particle from one end of the $\Gamma$ chain to the other end: in the language of spins, this corresponds to flipping an extensive number of them.
III Counterdiabatic driving
III.1 Brief review of counterdiabatic driving
The adiabatic theorem of quantum mechanics guarantees that the quantum dynamics, generated by a time-dependent Hamiltonian $H[\lambda(t)]$, tracks its instantaneous eigenstates, provided that variations in time are slow enough. The adiabaticity condition reads
$\dot{\lambda}\,\frac{|\bra{m[\lambda(t)]}\partial_{\lambda}H[\lambda(t)]\ket{n[\lambda(t)]}|}{(E_{n}[\lambda(t)]-E_{m}[\lambda(t)])^{2}}\ll 1,$ (20)
where |$\ket{n[\lambda]}$ are the instantaneous eigenstates and $E_{n}[\lambda]$ the corresponding energies. If the condition above is met, then transitions between eigenstates are suppressed [52]; see as well Refs. [7, 8] for refinements on the adiabaticity condition. Yet, a perfect adiabatic tracking can be obtained at any finite driving rate, if the dynamics is generated by the modified counterdiabatic driving (CD) Hamiltonian [48, *Demirplak2005Assisted, *Demirplak2008Consistency, 51]
$H_{\mathrm{CD}}(t)=H[\lambda(t)]+\partial_{t}\lambda\cdot H_{1}[\lambda(t)],$ (21)
with
[…]
where the last equation assumes that the spectrum of $H[\lambda(t)]$ is nondegenerate [51]. The first term on the right-hand side of $H_{1}[\lambda]$ is the generator of parallel transport familiar from proofs of the adiabatic theorem. The second term, being diagonal in the instantaneous eigenbasis, provides the correct Berry phase associated with a perfectly adiabatic trajectory [101].
The challenge in utilizing the CD Hamiltonian, such as the one in Eq. (22), comes from the fact that the terms involved are highly non-local and multiple-body, being |$\outerproduct{\partial_{\lambda}n}{n}$ | similar in structure to a projector [53]. For this reason, a growing body of works has been investigating the possibility of truncating the CD Hamiltonian to few-body local terms while retaining the suppression of diabatic transitions [53, 54, 56, 55, 58, 59, 76, 77]. A convenient framework to this end relies on the Floquet-Krylov expansion of the CD term [59, 60, 61]. Using the equivalent representation
[…]
with $\epsilon=0^{+}$, one can expand the integrand in powers of $x$ and find, upon integration,
$H_{1}[\lambda]=i\sum_{n=1}^{\infty}\alpha_{n}\underbrace{[H,\cdots,[H}_{2n-1\text{ times}},\partial_{\lambda}H]\cdots].$ (25)
Here, $\alpha_{n}=(-1)^{n}\epsilon^{-2n}$ is formally diverging, but allowing the $\alpha_{n}$’s to be free variational parameters, one obtains a controlled expansion of the CD term with increasingly non-local operators [59]. To fix the parameters $\alpha_{n}$, one may either employ a QAOA-inspired procedure, in which they are updated recursively by using as cost function the energy at the end of the driving process [102, 62, 63, 64], or an a priori method like the minimization of the action [58]
[…]
where
$G(H_{1}):=\partial_{t}H[\lambda(t)]-i[H[\lambda(t)],H_{1}],$ (27)
with respect to the parameters $\alpha_{n}$. We choose to follow the second route since it allows for greater analytical control and does not rely on heavy numerical optimization.
III.2 Approximate counterdiabatic driving for the spin-glass bottleneck model
Figure 4: The local approximate counterdiabatic Hamiltonians $H_{1}^{(1)}$, Eq. (28), and $H_{1}^{(2)}$, Eq. (30), take a simple form in the $\Gamma$ free fermionic representation, Eq. (14): they are a longer-range hopping terms, with fine-tuned coefficients $i\alpha_{j},i\beta_{j}$. Higher-order variational ansatzes yield a similar structure with longer-range couplings.
Following Refs. [58, 59], consider the lowest-order approximant to the CD term in the Floquet-Krylov expansion. At the first order, the commutator $[H,\partial_{\lambda}H]$, with $H$ in Eq. (1), contains terms $\sigma_{j}^{y}\sigma_{j+1}^{z}+\sigma_{j}^{z}\sigma_{j+1}^{y}$, motivating the variational ansatz
$H_{1}^{(1)}[\lambda]=\sum_{j=1}^{L}\alpha_{j}\left(\sigma_{j}^{y}\sigma_{j+1}^{z}+\sigma_{j}^{z}\sigma_{j+1}^{y}\right).$ (28)
The coefficients $\alpha_{j}$ can be found by minimizing the action (26). This yields the equations
$\left[8(1-\lambda)^{2}+\lambda^{2}\left(J_{j-1}^{2}+2J_{j}^{2}+J_{j+1}^{2}\right)\right]\alpha_{j}\\ +2\lambda^{2}J_{j}\left(J_{j-1}\alpha_{j-1}+J_{j+1}\alpha_{j+1}\right)=-J_{j},$ (29)
with periodic boundary conditions implied, which can be solved via linear algebra techniques with a cost polynomial in $L$. Going to the second order, from the computation of the nested commutator $[H,[H,[H,\partial_{\lambda}H]]]$, one determines the variational ansatz $H_{1}^{(1)}[\lambda]+H_{1}^{(2)}[\lambda]$, where
$H_{1}^{(2)}[\lambda]=\sum_{j=1}^{L}\beta_{j}\left(\sigma_{j}^{y}\sigma_{j+1}^{x}\sigma_{j+2}^{z}+\sigma_{j}^{z}\sigma_{j+1}^{x}\sigma_{j+2}^{y}\right).$ (30)
Again, the coefficients $\{\alpha_{j},\beta_{j}\}$ can be determined by solving a $2L\times 2L$ linear system (Eq. (125)). Generalizing, one needs to solve a $nN\times nN$ linear system for the $n$-th order term of the Floquet-Krylov expansion. Thus, for finite $n$, the CD term can be determined efficiently with the help of a classical computer.
At all orders of the Floquet-Krylov expansion of the CD Hamiltonian, Eq. (25), the approximate CD term for the spin-glass bottleneck model admits a free-fermionic representation. The full derivation is reported in App. C, and it leads to
$H_{1}^{(1)}[\lambda]$ $=-2i\sum_{j=1}^{L}\tilde{\alpha}_{j}\left[c^{\dagger}_{j}c^{\dagger}_{j+1}+c_{j}c_{j+1}\right],$ (31)
$H_{1}^{(2)}[\lambda]$ $=-2i\sum_{j=1}^{L}\tilde{\beta}_{j}\left[c^{\dagger}_{j}c^{\dagger}_{j+2}+c_{j}c_{j+2}\right].$ (32)
Notice that the CD terms take the form of longer-range hoppings. A similar longer-range-hopping form is acquired in the $\Gamma_{j}^{\pm}$ basis as well, see Eqs. (123) and (129). Thus, CD helps the driving by preventing localization with two-fermion longer-range hopping terms; see also Sec. VI for further discussion.
IV Results
Here, we present our main results regarding the behavior of CD in the quantum driving of the spin-glass bottleneck model. While the Hamiltonian in Eq. (1) could be studied fully analytically, the introduction of the approximate CD terms forces one to resort to numerical methods: the variational equations (29) and (125) do not seem to have an explicit analytical solution in the general case. Nevertheless, the simple structure of the model, together with the mappings to $c$ and $\Gamma$ fermions performed above, allows for a thorough understanding of the effect of CD on the driving process.
IV.1 Counterdiabatic driving and gap amplification
According to the adiabatic theorem, adiabaticity can be preserved upon reducing the driving time, provided that the gap is increased by the same factor. However, the construction of the CD term outlined in Sec. III.1 indicates that CD is more than a mere gap amplification method: the operators involved are constructed in such a way to transport a state along its adiabatic path, and they do so by enforcing parallel transport while accounting the corresponding quantum phase. By contrast, catalysts and gap-amplifying terms  just enhance the dephasing due to highly oscillatory terms $e^{i(E_{m}-E_{n})t}$ for $m\neq n$, making the adiabatic approximation hold to a better degree [30, 31, 32, 33, 34, 35, 36, 37, 38, 39].
Notice that using a schedule $\lambda(t)$ with vanishing derivative at the endpoints, the eigenstates of $H_{\mathrm{CD}}[\lambda(t)]$, Eq. (21), coincide with those of $H[\lambda(t)]$ at the beginning ($\lambda(0)=0$) and the end ($\lambda(T)=1$) of the protocol. Therefore, when using approximate CD, one can apply the adiabatic theorem to the dynamics generated by $H_{\mathrm{CD}}[\lambda(t)]$ and try to infer the success of the driving process from the minimal gap $\Delta_{\mathrm{min,CD}}$ along the path (by construction, the exact CD term yield success with probability 1 and this study would not make sense there). Noticing that the approximate $H_{\mathrm{CD}}$ depends on the final time $T$, one cannot blindly use the adiabaticity condition to estimate the time $T_{\mathrm{ad}}$ needed for adiabaticity to hold. Nevertheless, it can be used to determine self-consistently $T_{\mathrm{ad}}$, if the dependence $\Delta_{\mathrm{min,CD}}=\Delta_{\mathrm{min,CD}}(T)$ is known.
Figure 5: Gap amplification from counterdiabatic driving (CD). (a) For any driving rate, the minimal gap $\Delta_{\mathrm{min,CD}}$ along the CD-assisted adiabatic path remains exponentially small in the system size, both for the 1st order (Eq. (28), filled markers) and 2nd order (Eq. (30), empty markers) variational ansatzes. The bare minimal gap $\Delta_{\mathrm{min}}$ is shown for comparison with black crosses. (b) The CD-assisted gap $\Delta_{\mathrm{min,CD}}$, despite being exponentially small in the system size, is still exponentially larger than the bare gap $\Delta_{\mathrm{min}}$: all the log-log fits (lines) are compatible with a growth $\Delta_{\mathrm{min,CD}}\sim\Delta_{\mathrm{min}}e^{c(T)L}$, with $c(T)=\alpha-\alpha_{\mathrm{CD}}(T)$.
In Fig. 5, the numerical results for the gap $\Delta_{\mathrm{min,CD}}(T)$ of the CD Hamiltonian are shown. The cubic ramp $\lambda(t)=3(t/T)^{2}-2(t/T)^{3}$, that satisfies $\dot{\lambda}(0)=\dot{\lambda}(T)=0$, is used. From Fig. 5a, one can see that, for large system sizes and long driving times, the minimum of the gap remains exponentially small in the system size: $\Delta_{\mathrm{min,CD}}\propto e^{-\alpha_{\mathrm{CD}}L}$. From Fig. 5b, however, one can notice that $\Delta_{\mathrm{min,CD}}$ is still exponentially larger than $\Delta_{\mathrm{min}}$, as a function of the system size $L$ and for every fixed time $T$. In other words, $\alpha_{\mathrm{CD}}<\alpha$. Also, there is no appreciable difference in the values of $\alpha_{\mathrm{CD}}$ extracted from either the first- or second-order expansions; only the prefactor of the exponential is larger for the second-order one.
From plots like that in Fig. 5a, the exponential decay rate $\alpha_{\mathrm{CD}}$ of $\Delta_{\mathrm{min,CD}}$ is extracted for many values of $T$, and the values are reported in Fig. 6. The fits are performed on the first-order expansion gaps; quantitatively similar results follow from using the second-order expansion gaps. One can see that already the first-order ansatz is sufficient to provide an exponential speed-up for moderate driving times, $T\lesssim 25$, as $\alpha_{\mathrm{CD}}$ is substantially smaller than $\alpha$. Increasing the driving time, however, $\alpha_{\mathrm{CD}}$ approaches $\alpha$: higher order CD terms would be required to maintain the same gap amplification.
The conclusion above can be put on quantitative grounds by estimating the adiabatic timescale for approximate CD with the inverse of the minimal gap, $T_{\mathrm{ad,CD}}\sim\Delta^{-2}_{\mathrm{min,CD}}(T)$. Using $\Delta_{\mathrm{min,CD}}(T)\simeq\Delta_{0}e^{-\alpha_{\mathrm{CD}}(T)L}$, with $\alpha_{\mathrm{CD}}(T)\simeq\alpha-\delta/T^{2}$ that is shown to hold up to high precision in Fig. 6b ($\Delta_{0}$ and $\delta$ are two fitting coefficients), one obtains
$T_{\mathrm{ad,CD}}\approx T_{\mathrm{ad}}\,e^{-2\delta L/T^{2}_{\mathrm{ad,CD}}}\approx T_{\mathrm{ad}}\left(1-\frac{2\delta L}{T^{2}_{\mathrm{ad}}}\right),$ (33)
where in the last step we expanded up to first order the exponential and approximated $T_{\mathrm{ad,CD}}$ with $T_{\mathrm{ad}}$. One can see that the solution $T_{\mathrm{ad,CD}}$ is significantly reduced for moderate driving times, and it converges to the bare value, $T_{\mathrm{ad}}$, upon increasing $T$. Let us remark that both this empirical power law and the coefficients are expected to change in favour of smaller adiabatic times for larger order expansions.
Figure 6: The approximate CD path displays a minimal gap $\Delta_{\mathrm{min,CD}}\propto e^{-\alpha_{\mathrm{CD}}(T)L}$. (a) Rate $\alpha_{\mathrm{CD}}$ as a function of the driving time $T$: the value of $\alpha_{\mathrm{CD}}$ (dots) saturates at the analytical value $\alpha$ in the absence of CD in Eq. (19) (solid lines), already for moderately short times. (b) The approach to the asymptotic value is well fitted by a power law: $\alpha-\alpha_{\mathrm{CD}}(T)\sim T^{-2}$.
IV.2 Performance of counterdiabatic driving
Above, the adiabatic theorem was used to predict the performance of approximate CD. It was found that the gap remains exponentially small in the system size for all driving times, with the gain of CD being less and less significant as the driving time increases. In order to quantify precisely to what extent CD helps the driving process, however, we next integrate the time-dependent Schrödinger equation and see whether the success probability is actually increased.
As before, a cubic ramp $\lambda(t)=3(t/T)^{2}-2(t/T)^{3}$, satisfying $\dot{\lambda}(0)=\dot{\lambda}(T)=0$, is employed. Upon initializing the system in the paramagnetic ground state of $H[\lambda=0]$, we consider the evolution with the approximate CD Hamiltonians $H[\lambda(t)]+\dot{\lambda}(t)H^{(1)}[\lambda(t)]$ (first order ansatz) or $H[\lambda(t)]+\dot{\lambda}(t)[H^{(1)}[\lambda(t)]+H^{(2)}[\lambda(t)]]$ (second order ansatz). Because of the free-fermionic nature of the model under consideration, we evolve the single-particle operators $(c_{j}^{\phantom{\dagger}},c_{j}^{\dagger})$ or $\Gamma_{k}^{\pm}$ by means of Bogoliubov-de Gennes equations, and then compute the desired expectation values at the end of the protocol with the appropriate contractions; see App. E for all the details. This technique allows one to reach much larger system sizes, which are needed in the present case to extract the asymptotic scalings. Incidentally, this is the reason why we employ the chain in Eq. (1) instead of examples of NP-hard Ising Hamiltonians [103].
[…]
The results of our numerical simulations to assess the performance of the driving are presented in Figs. 7. First, we compute the kink number in Fig. 7, i.e., the number of frustrated bonds at the end of the process. It is given by the expectation value on the final state of the operator
$K:=\frac{1}{2}\sum_{j=1}^{L}\left[1-\mathrm{sign}(J_{j})\sigma_{j}^{z}\sigma_{j+1}^{z}\right].$ (34)
[…]
The findings of Fig. 7 agree with the known results about the Kibble-Zurek (KZ) density of defects formed upon crossing a quantum phase transition [79, 80, 81, 82, 83]. The collapse in Fig. 7b shows that the defect density scales as $\sim T^{-\nu/(1+z\nu)}=T^{-1/2}$, where $\nu=1$ is the correlation length exponent of the transverse-field Ising model and $z=1$ is its dynamical exponent. The presence of the three weak links manifests only in the breakdown of the KZ scaling at large times, with the formation of the plateau. The breakdown of the KZ scaling at short times, due to the introduction of approximate CD terms, confirms the early results of Ref. [53] in this setting. Thus, our findings demonstrate that CD improves fast driving at short times but does not help significantly to overcome the exponentially small gap, as further discussed in Sec. VI.
V Quantum Brachistochrone Counterdiabatic Driving (QBCD): exponential speedup
In this section, we show how the efficiency of the driving can be improved beyond the low-order CD expansions by canceling the transitions between the ground state and the first excited state around the critical point $\lambda=\lambda_{c}$.
V.1 Comparison to the local approach in the spin glass toy model
Above, the first- and second-order ansatzes were used without requiring any knowledge of the instantaneous spectral properties. Here, an approximate knowledge of the spectrum at a single point in parameter space is assumed, and much better results are obtained. This approach is reminiscent of reverse quantum annealing, where an approximate ground state of the problem Hamiltonian is used to improve the performance of the adiabatic algorithm [104, 105, 106]. In addition, this method differs from the early approaches to CD in many-body systems, where the exact CD term was truncated according to the locality of the operators involved [53, 56]. Here, instead, the complete CD Hamiltonian is employed, but with an approximate knowledge of the eigenvalues and eigenvectors. The truncation is, in turn, performed in the number of parameter values at which it is evaluated. Although in the worst-case scenario, this does not reduce the computational cost of the search for the final ground state when compared to diagonalizing the final Hamiltonian, the use of additional numerical methods such as reverse annealing or quantum Monte Carlo to extract approximate knowledge of the critical point are feasible strategies to keep the numerical cost under control. The simplest version of this approach is to take a single parameter value at which the approximate knowledge is required.
The idea is to focus on the bottleneck of the driving dynamics and use approximate left- and right-localized edge states, which are responsible for the exponentially closing gap, to construct the CD term at the critical point. By multiplying such a term with the time-dependent derivative of the schedule, it is ensured that the spectrum is not altered at the endpoints $\lambda=0$ and 1 (see App. C.4). In the language of the one-particle effective model, Eqs. (16)–(17), the QBCD Hamiltonian reads
$\mathcal{H}^{\mathrm{QBCD}}=i\,\frac{(\psi_{\mathrm{R}}|\partial_{\lambda}\mathcal{H}[\lambda_{c}]|\psi_{\mathrm{L}})}{\Delta_{\min}}\big{[}|\psi_{\mathrm{R}})(\psi_{\mathrm{L}}|-|\psi_{\mathrm{L}})(\psi_{\mathrm{R}}|\big{]},$ (35)
where $\lambda_{c}$ denotes the point in parameter space where the gap is minimal and equal to $\Delta_{\min}$, $|\psi_{\mathrm{L,R}})$ denote the left/right localized edge states in the $L$-dimensional linear space of the effective chain. Equation (35), inspired by CD, shares the structure of the Hamiltonian in the quantum brachistochrone problem for the time-optimal evolution within this two-level subspace [107, 108, 109], motivating the term QBCD. Upon contracting with the $\Gamma$ operators, one obtains the full expression of the QBCD Hamiltonian. The gap and the left and right eigenvectors are obtained via analytical approximations valid in the close neighborhood of the critical point (see App. C.4 Notice also that $\mathcal{H}^{\mathrm{QBCD}}$ takes the form of a $\sigma^{y}$ Pauli matrix in the reduced subspace of the left- and right-localized single-particle states.
Figure 8: Gap amplification from the QBCD term, Eq. (35). Comparing with Fig. 5, one can see that $H^{\mathrm{QBCD}}$ is able to enlarge the minimal gap on the adiabatic path by a significant amount for all system sizes and driving times.
The QBCD Hamiltonian, Eq. (35), cancels exactly the transition between the ground and first excited states only at the gap-closing point; however, it also provides an efficient CD term in the neighborhood thereof. This feature can be inferred from the results shown in Fig. 8 and 9: both the gap $\Delta_{{\min},\mathrm{QBCD}}$ and the ground state fidelity are significantly increased with respect to the bare Hamiltonian. Here, the fidelity is captured by the kink number, since a value $K<1$ can come only from the ground state. In particular, one can see from Fig. 9a that for values of $T$ for which the bare Hamiltonian exhibits a kink number plateauing very close to 1, the QBCD term is able to lower it to $\expectationvalue{K}_{T}\approx 0.5$. This implies that the ground state is found at the end of the protocol with a finite probability of $\approx 0.5$, which is significantly larger than the one found with the bare protocol.
The effects of QBCD remain remarkably robust against deviations from the true critical point $\lambda_{c}$. This robustness is especially valuable in real-world applications, where experimental noise and fluctuations often hinder the precise identification of critical parameters. To test this, we studied the behavior of the minimal gap by varying the estimated critical point $\lambda^{*}_{c}$ within the range $\lambda_{c}\in[0.85\lambda^{*}_{c},1.15\lambda^{*}_{c}]$.
Beyond analytical approximations (App. C.4), we extended our analysis by performing a fully numerical search for the ground state energy gap and corresponding eigenstates using only 40 sampling points. This approach requires significantly less computational effort than simulating the full QBCD protocol, yet still captures the essential features needed for its construction. As such, it broadens its applicability to a wide class of spin-glass systems, even in the absence of analytical control over the critical properties. Figure 10 shows the performance of the resulting QBCD Hamiltonian, constructed from the corresponding eigenstates $\left\lvert\psi_{R}\right)$ and $\left\lvert\psi_{L}\right)$, and the numerically determined minimal gap $\Delta_{\mathrm{min,QBCD}}$ at the shifted $\lambda_{c}$. Apart from verifying that the QBCD can achieve a polynomial decreasing gap with respect to $L$, this scaling also remains stable up to a constant factor across deviations up to approximately $10\%$ from the true critical point.
Interestingly, for the analytically approximated QBCD, the minimal gap even increases slightly with small shifts away from the critical point. Although the gap eventually decreases—for example, at $\lambda_{c}=1.125,\lambda_{c}^{*}$—the numerically obtained QBCD Hamiltonian still achieves a two-order-of-magnitude improvement over local expansions for system sizes as large as $L=600$. Constructing the QBCD using approximate numerical or experimental methods that demand far less computational effort than running the full simulation, thus suffices to realize exponential speedups in adiabatic protocols.
Figure 9: Performance of the driving assisted by the QBCD term, Eq. (35). The expectation value of the kink number on the final state plateaus for intermediate driving times, but the actual value of $\expectationvalue{K}_{T}$ is reduced when employing $H^{\mathrm{QBCD}}_{1}$ (dashed lines) with respect to the bare evolution (solid lines).
Figure 10: Robustness of the minimal gap scaling against deviations between the applied QBCD Hamiltonian and that taken exactly at $\lambda_{c}$. Red dashed lines correspond to the approximate analytical results, while the blue dotted ones follow from a pure numerical approach. The black solid line corresponds to the analytical approach at $\lambda_{c}=\lambda^{*}_{c}$. The curves clearly indicate a polynomial decay rather than an exponential one.
Finally, let us investigate the efficiency of the variational and QBCD approaches presented from the perspective of energy resources. While reducing external resources is essential for experimental realizations [110, 111], the energetic costs determine the speed at which the assisted process can stay adiabatic [50, 112, 113]. The Hilbert-Schmidt norm of the Hamiltonians, defined as $\|H_{1}\|^{2}=\mathrm{Tr}[H^{\dagger}_{1}\,H_{1}]$, provides a natural measure to quantify the cost of CD [50, 53, 60, 61]. It can be shown (see App. D) that the energy costs of the low-order expansions grow as
$\left\|H_{1}^{(1,2)}\right\|^{2}\sim L$ (36)
for the first- and second-order variational expansions, respectively. The QB approach surpasses the variational method in this aspect as well. As shown in App. D, the numerator and the denominator in Eq. (35) grow as (to leading order in $\ell$)
$\Delta_{\min}\simeq\ell(\lambda-1)e^{-(\ell-1)\kappa}(1+e^{-2\mu})\\ +\ell\lambda\left(e^{-\ell\kappa-2\mu}+e^{-(\ell-2)\kappa}\right)$ (37)
and
$(\psi_{\mathrm{R}}|\partial_{\lambda}\mathcal{H}[\lambda_{c}]|\psi_{\mathrm{L}})\approx\\ \ell\left(e^{-(\ell-1)\kappa}(1+e^{-2\mu})+e^{-\ell\kappa-2\mu}+e^{-(\ell-2)\kappa}\right).$ (38)
The two equations above imply a constant norm to leading order:
$\left\|\mathcal{H}^{\mathrm{QBCD}}\right\|^{2}\sim O(1)\,.$ (39)
This fact is a consequence of the exponentially localized nature of the two single-particle states involved in the avoided crossing and their appearance both in the gap and the transition matrix element.
To also incorporate the time dependence in the computation, in Fig. 11 the quantities $\frac{1}{T}\int_{0}^{T}\mathrm{d}t\,\|H_{1}\|^{2}$ are presented. It is clearly demonstrated that $\|H_{1}^{(1)}\|$ and $\|H_{1}^{(2)}\|$ increase with the system size, while $\|\mathcal{H}^{\mathrm{QBCD}}\|$ remains independent of $L$ to leading order. Thus, the QBCD approach not only achieves significantly larger fidelities to the target state, but it does so with a constant energy cost.
Figure 11: Energetic cost of implementing the counterdiabatic terms, as measured by the average Hilbert-Schmidt norm of the Hamiltonian along the driving path. While the first- and second-order variational terms have a cost that grows with the system size, the QBCD approach of Eq. (35) converges to an $L$-independent value.
Let us stress that the specific features of the model and its free-fermionic representation were used solely to access large system sizes, to compare the purely numerical construction with analytical approximations, and to also test analytically its energetic costs, while the general effects of the QBCD are expected to hold across a wide range of spin glass problems as detailed in the following sections.
V.2 QBCD in the $3$-regular Max Cut problem
As demonstrated above, the QBCD Hamiltonian achieves finite ground-state fidelities at exponentially shorter driving times compared to the first- and second-order local CD expansions, within the minimal spin-glass bottleneck model. This is evidenced by the polynomial closing of the minimum energy gap and by the finite fidelities in the exponentially long plateaus, with vanishing occupation of the ground state with the local CD expansion. Moreover, the QBCD approach proves remarkably robust: its performance remains stable even under significant deviations from the exact critical point. It also retains its effectiveness when constructed through a purely numerical procedure, which requires computational resources that are subleading compared to those needed for simulating the full time evolution. This numerically accessible construction not only improves practicality but also brings experimental realization within closer reach. Furthermore, the QBCD protocol maintains a bounded energy scale, further underscoring its advantage compared to the linearly diverging energy cost of the local expansions.
To this end, we investigate the performance of the local CD expansion and the QBCD method on a realistic NP-hard problem Hamiltonian, the $3$-regular Max Cut [114, 115]. The $3$-regular Max Cut problem describes the optimization task of dividing a $3$-regular random graph into two subgraphs such that the number of connecting edges is maximal. The solution for a given graph is equivalent to finding the ground state of the following Hamiltonian,
$H=\sum_{c=1}^{M}\frac{1+\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}}{2},$ (40)
where $c$ denotes the edge connecting the two spins representing the vertices of the graph. Edge configurations are distributed uniformly such that all spins appear in exactly three interaction terms without repetition, $c_{1}\neq c_{2}$, implying a total $M=3L/2$ number of quadratic terms. In Ref. [114], it was shown that the ground state energy gap of the Max Cut exhibits an exponential decay even for intermediate system sizes, $L=8,\dots,18$ for problem instances with two-fold ground state degeneracy. Correspondingly, we also restrict the exposition to these instances and consider the time-evolution initiated from the fully paramagnetic ground state governed by the Hamiltonian
$H(t)=-\lambda\sum_{c=1}^{M}\frac{1+\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}}{2}-(1-\lambda)\sum_{j=1}^{L}\,\sigma^{x}_{j},$ (41)
where the time-dependence of the linear protocol was omitted for brevity, $\lambda(t)=t/T,\quad t\in[0,T]$.
First, we probe the efficiency of the local variational expansion. Similar to Sec. III.2, we consider two and three-body terms in which all interactions are restricted to spins connected by the graph edges. Allowing also for all local couplings to be optimized, the first and second order CD corrections read
$H^{(1)}_{1}=\sum_{c=1}^{M}\alpha_{c}(\sigma^{z}_{c_{1}}\sigma^{y}_{c_{2}}+\sigma^{z}_{c_{1}}\sigma^{y}_{c_{2}}),$ (42)
$H^{(2)}_{1}=\sum_{c,c^{\prime}=1}^{M}\delta_{c_{2},c^{\prime}_{1}}\beta_{c,c^{\prime}}(\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{y}_{c^{\prime}_{2}}+\sigma^{y}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}).$ (43)
In the second-order expansion, $H^{(2)}_{1}$, three-body terms run over connected spins in the interaction graph of the Max Cut problem, indicated by the constraint $\delta_{c_{2},c^{\prime}_{1}}$ and the adapted order of indices, $c_{1}<c_{2}=c^{\prime}_{1}<c^{\prime}_{2}$. The corresponding minimization equation with the first order expansion is given by (see App. F)
$\left[32(1-\lambda)^{2}+12\lambda^{2}\right]\alpha_{c}+2\lambda^{2}\sum_{c^{\prime}}\alpha_{c^{\prime}}\left[\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})+(1-\delta_{c_{1},c^{\prime}_{1}})\delta_{c_{2},c^{\prime}_{2}}+2\delta_{c_{1},c^{\prime}_{2}}+2\delta_{c_{2},c^{\prime}_{1}}\right]=-2(\lambda-2).$ (44)
The local parameters for the second order term (for the detailed derivation see App. F) are obtained by minimizing the action, Eq. (26), with respect to both $\alpha_{c}$ and $\beta_{c,c^{\prime}}$,
$\left[32\lambda^{2}+16(1-\lambda)^{2}\right]\alpha_{c}+(1-\lambda)^{2}\sum_{c^{\prime}}\alpha_{c^{\prime}}\left(\delta_{c_{2},c^{\prime}_{1}}+\delta_{c_{2},c^{\prime}_{2}}(1-\delta_{c_{1},c^{\prime}_{1}})+\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})+\delta_{c_{1},c^{\prime}_{2}}\right)$ (45)
$+8(1-2\lambda)\sum_{c^{\prime}}(\beta_{c,c^{\prime}}\delta_{c^{\prime}_{1},c_{2}}+\beta_{c^{\prime},c}\delta_{c^{\prime}_{2},c_{1}})=-2(2-3\lambda),$
$16((1-\lambda)^{2}+\lambda^{2})\beta_{c,c^{\prime}}\delta_{c_{2},c^{\prime}_{1}}+\lambda^{2}\sum_{c^{\prime\prime}}\beta_{c^{\prime\prime},c^{\prime}}\left[(1-\delta_{c^{\prime\prime}_{1},c_{1}})\delta_{c^{\prime\prime}_{2},c_{2}}+(1-\delta_{c^{\prime\prime}_{2},c_{2}})\delta_{c^{\prime\prime}_{1},c_{1}}\right]=2(2\lambda-1)\delta_{c_{2},c^{\prime}_{1}}(\alpha_{c}+\alpha_{c^{\prime}}).$ (46)
Similar to the case of the minimal spin glass model, the low-order local CD expansion provides a negligible improvement in shortening the adiabatic time scale. Besides the ground state gap, we investigate the ground state fidelity, which is a more suitable measure of adiabaticity for the achievable system sizes, $L=6,\dots,18$. As shown in Fig. 12, the ground state gap increases only by a constant factor, even for short driving times, $T=2$. This increase is slightly more pronounced in the second-order expansion compared to the slow exponential decay observed in the bare process without CD. Additionally, our findings also reveal that the relative difference from the gap without CD follows a universal curve decaying approximately as $(\Delta^{(1,2)}_{\mathrm{min,CD}}-\Delta_{\mathrm{min}})/\Delta_{\mathrm{min}}\sim T^{-2}$, as shown in Fig. 13. Notably, this result further demonstrates the robustness of the predictions drawn from the minimal spin glass model, Fig. 6. Regarding the final ground state fidelities, even weaker efficiency is observed for the first two orders of the local expansion. While for long driving times, minimal improvement is observed, for faster processes, the final fidelity becomes even smaller than in the bare process. These observations are displayed in Fig. 14.
Next, we test the performance of the QBCD Hamiltonian, constructed by the ground state gap, $\Delta$, and the many-body ground ($\lvert\mathrm{GS}\rangle$) and first excited states ( $\lvert\mathrm{ES}\rangle$) near the critical point
$H_{\mathrm{QBCD}}=$ (47)
$-i\frac{\langle\mathrm{GS}\lvert\partial_{\lambda}H\rvert\mathrm{ES}\rangle}{\Delta}\left(\lvert\mathrm{ES}\rangle\langle\mathrm{GS}\rvert-\lvert\mathrm{GS}\rangle\langle\mathrm{ES}\rvert\right).$
The approximate knowledge of these quantities is obtained via the same purely numerical procedure employed for the minimal spin glass model, based on a search over only $40$ parameter values. While in the minimal model this choice had limited computational relevance, for the realistic spin glass problem, it clearly requires only subleading numerical costs compared to the subsequent QBCD-assisted simulation.
To further improve its feasibility for near-term experimental realizations, we sparsify the QBCD Hamiltonian to the density of the quadratic local expansion by keeping only the $L(L+1)/2$ largest absolute valued matrix elements. Although the exact QBCD construction involves only two states, selected out of the $\sim 4^{L}$ possible energy level pairs of the full CD driving, the sparsification further reduces its non-locality exponentially, bringing its experimental realization within reach. Moreover, the computational cost of constructing this sparsified QBCD Hamiltonian grows only polynomially with the system size, making it accessible for classical large-scale algorithms as well [116, 117].
To this end, the performance of the sparsified QBCD is also tested via the ground state gap and the final fidelity. As shown in Fig. 12, the ground state gap is exponentially amplified, exhibiting an approximately polynomial decay even at intermediate driving times, $T=5,20$, whereas it only approaches the behavior without CD driving close to the fully adiabatic timescales. The fidelity exhibits similar features, remaining finite with the sparsified QBCD even at short driving times, $T=1,5$, while the fidelity achieved using the local CD approach becomes exponentially small under the same conditions. In the latter case, the fidelity only becomes finite close to adiabaticity, but with negligible improvements compared to the bare process. Whereas, in this regime, the sparsified QBCD performs more efficiently, achieving approximately unit fidelity.
Figure 12: Ground state energy gaps for the Max Cut problem without CD, with first-order and second-order local expansions (dotted lines), and with the sparsified QBCD (solid lines). Local methods yield only a constant-factor improvement over the exponentially small gap of the bare protocol, as shown for short driving times ($T=2$). The sparsified QBCD leads to exponential gap amplification across all timescales ($T=2,5,20,80$).
Figure 13: Relative difference between ground-state gaps without CD and with the local CD approaches. The improvement achieved by first- (solid lines, filled dots) and second-order (dashed lines, empty dots) local expansions decays approximately quadratically with system size. Notably, the amplification of the gap vanishes at driving times just preceding the onset of finite fidelity.
Figure 14: Ground-state fidelity for the Max Cut problem. Comparison of the second-order local CD (dashed lines) and no CD (dotted lines) approaches, with the sparsified QBCD method (solid lines). Local strategies offer negligible improvement even in slow processes, while the sparsified QBCD achieves finite fidelity already at short driving times, where the other methods yield exponentially small values.
Finally, we compare the performance of the sparsified QBCD and the local expansion with the COLD approach [77]. Within this scheme, one introduces an additional local operator, $\mathcal{O}_{\mathrm{opt}}$, to the problem Hamiltonian with optimizable additional parameters,
$H_{\mathrm{COLD}}(t,\boldsymbol{\gamma})=$ (48)
$-\lambda\sum_{c=1}^{M}\frac{1+\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}}{2}-(1-\lambda)\sum_{j=1}^{L}\,\sigma^{x}_{j}+f(t,\boldsymbol{\gamma})\mathcal{O}_{\mathrm{opt}}.$
For computational convenience and following the strategy of Ref. [77], we choose the optimized operator as the sum of longitudinal spin operators, $\mathcal{O}_{\mathrm{opt}}=\sum_{j}\sigma^{z}_{j}$. The time- and parameter-dependent coefficient is represented by its Fourier decomposition,
$f(t,\boldsymbol{\gamma})=\sum_{k=1}^{N_{k}}\gamma_{k}\sin\left(\pi\frac{t}{T}\right),$ (49)
with $N_{k}$ denoting the number of modes. This form ensures that the CD term vanishes at the initial and final times, i.e., $f(t=0,T,\boldsymbol{\gamma})=0$, thereby preserving the original Hamiltonian at the endpoints. The Fourier coefficients, $\gamma_{k}$ are numerically optimized within the local CD expansion framework, where the final fidelity of the evolved state serves as the objective function. Since the presence of the longitudinal field allows for linear Pauli operators in the lowest order CD expansion, we consider the second-order local CD approach
$H^{(2)}_{\mathrm{COLD}}=\sum_{j}\epsilon_{j}\sigma^{y}_{j}+\sum_{c}\alpha_{c}(\sigma^{z}_{c_{1}}\sigma^{y}_{c_{2}}+\sigma^{y}_{c_{1}}\sigma^{z}_{c_{2}}).$ (50)
The minimization principle is written in terms of this second-order local CD term, together with the optimizable operator,
[…]
$G(H^{(2)}_{\mathrm{COLD}})=$ $\partial_{t}H(t)+\partial_{t}f(t,\boldsymbol{\gamma})\mathcal{O}_{\mathrm{opt}}$
$-i[H(t),H^{(2)}_{\mathrm{COLD}}]-if(t,\boldsymbol{\gamma})[\mathcal{O}_{\mathrm{opt}},H^{(2)}_{\mathrm{COLD}}].$
The corresponding trace minimization and the resulting system of equations for the sets of $\alpha_{c}\left(\lambda,f(t,\boldsymbol{\gamma})\right)$ and $\epsilon_{j}\left(\lambda,f(t,\boldsymbol{\gamma})\right)$ are detailed in App. G. To perform the numerical optimization, we employed MATLAB R2023’s built-in genetic algorithm to search for the optimal set of Fourier coefficients $\gamma_{k}$ that maximize the final fidelity. The optimization was repeated $100$ times to ensure the convergence of the results. As a first clear sign of the inferiority of the COLD approach compared to the QBCD, the procedure proved to be computationally demanding, restricting our analysis to systems with up to $L=10$ spins—only almost half the size studied in the QBCD framework. Due to similar numerical restrictions, the number of Fourier modes was also limited to $N_{k}=4$. Despite these constraints, our setting already surpasses that of Ref. [77], by using twice the number of spins and one extra Fourier mode, allowing for a more systematic exploration of its performance in spin glasses. Crucially, the fact that such modest extensions already impose severe computational bottlenecks underscores a central message: the practical implementability and efficiency of the QBCD method is fundamentally superior to what the COLD approach can offer. They also highlight the significant advantage of the QBCD approach—not only in terms of algorithmic scalability but also in terms of its greater feasibility for experimental realization.
To this end, we benchmarked the performance of the second-order COLD method against both the sparsified QBCD protocol and the second-order local CD expansion, Eq. (43), using the final ground state fidelity as the primary measure for $L=6,8,10$ spins. Given the small system size, averaging over $10$ random graph instances was sufficient to obtain well-converged results for the average fidelity. As shown in Fig. 15, the COLD approach clearly outperforms the local expansion, particularly at intermediate annealing times, reflecting the benefits of its variationally optimized structure. Although the fidelity decreased rapidly in faster processes, it converged to the local expansion and the bare process as the adiabatic limit is reached.
Notably, the advantage of QBCD becomes most prominent modest driving regime, $T\lesssim 15$, where it continues to maintain fidelities close to unity. In contrast, both local and the COLD approaches begin to deteriorate, highlighting the onset of nonperturbative spin-glass effects that resist purely local counterdiabatic corrections. These results compellingly demonstrate that even in small systems, the spin-glass bottleneck can severely limit the effectiveness of local methods, regardless of how the driving protocol is optimized.
Note that the relatively small system size allows optimization effects to be more visible. In this regime, the COLD protocol—despite being limited to two-body operators—can slightly outperform the second-order local expansion with three-body terms. However, this advantage starts to fade away as the system size increases. The emergence of stronger spin-glass behavior and a growing number of diabatic transitions lowers the performance of COLD toward that of the second-order local expansion. Meanwhile, the superiority of QBCD becomes increasingly pronounced, as it continues to capture essential global features—through its exponentially suppressed non-locality—that remain out of reach for purely local approaches.
Figure 15: Fidelity as a function of the driving time for the sparsified QBCD, COLD approach, second order local expansion, and for the bare process, for various system sizes. (a) Even for $L=6$ QBCD provides a sizable improvement at intermediate times. (b) The advantage of QBCD becomes more apparent also at longer driving times for $L=8$. (c) For $L=10$ even close to adiabaticity, the QBCD clearly outperforms the COLD, while the latter gets closer to the second-order expansion and the bare process results.
V.3 Demonstration on the $3$-regular $3$-XORSAT
Finally, we also test the QBCD method on the $3$-regular $3$-XORSAT corresponding to optimization problems whose solutions can be represented as the ground state of a spin glass Hamiltonian with cubic spin interactions. In particular, the $3$-XORSAT problem can be formulated in terms of linear equations modulo $2$ with each of them containing three variables with the constraint that each variable appears in exactly three equations without repetition,
$\underline{\underline{A}}\cdot\underline{x}=\underline{y}\qquad(\text{mod}\,2).$ (53)
Here $\underline{\underline{A}}$ is an $L\times L$ matrix with each row containing only three non-zero elements corresponding to the prescription of the problem. As the possible outcomes can only be 00 or $1$, the assigned variables can take the same values as well, $x_{i}=0,1$. The corresponding Hamiltonian is given by
$H_{P}=\sum_{c}\frac{1+J_{c}\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{c_{3}}}{2},$ (54)
with alternating randomly as $J_{c}=\pm 1$. Following the strategy of Refs. [114, 13] we only consider such instances where the ground state is non-degenerate, allowing for a better characterization of the energy gap at the critical point. In these instances, one can drop the alternating coupling constant, and it is satisfactory to consider the Hamiltonian instead
$H_{P}=\sum_{c}\frac{1-\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{c_{3}}}{2},$ (55)
where the summation runs over $M=L$ equations. The corresponding quantum annealing Hamiltonian is written as
$H(t)=-\lambda\sum_{c}\frac{1-\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{c_{3}}}{2}-(1-\lambda)\sum_{j=1}^{L}\sigma^{x}_{j},$ (56)
exhibiting an exponentially small gap around $\lambda_{c}=1/2$, where exponentially long driving times are required to reach near adiabatic dynamics. To this end, we probe the performance of both the QBCD Hamiltonian and the first-order locally optimized variational CD ansatz. Similar to the Max Cut, we investigate the final ground state fidelity and the minimal energy gap at the critical point. The first-order CD expansion is given by
$H^{(1)}_{1}=\sum_{c}\alpha_{c}\left(\sigma^{y}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{c_{3}}+\sigma^{z}_{c_{1}}\sigma^{y}_{c_{2}}\sigma^{z}_{c_{3}}+\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{y}_{c_{3}}\right),$
with variational coefficients obtained by the solutions of the following equations (see App. H)
$18\lambda^{2}\alpha_{c}+96(1-\lambda)^{2}+2\lambda^{2}\sum_{c^{\prime}}\alpha_{c^{\prime}}\left[3\delta_{c_{1},c^{\prime}_{2}}+3\delta_{c_{2},c^{\prime}_{1}}+2\delta_{c_{1},c^{\prime}_{3}}+2\delta_{c_{3},c^{\prime}_{1}}\right]+2\lambda^{2}\sum_{c^{\prime}\neq c}\alpha_{c^{\prime}}\left[\delta_{c_{1},c^{\prime}_{1}}+\delta_{c_{2},c^{\prime}_{2}}+\delta_{c_{3},c^{\prime}_{3}}\right]$
$=-6(1-\lambda)-\frac{2}{3}\lambda\sum_{c^{\prime}\neq c}\left[\delta_{c_{1},c^{\prime}_{1}}+\delta_{c_{2},c^{\prime}_{2}}+\delta_{c_{3},c^{\prime}_{3}}+\delta_{c_{1}c^{\prime}_{2}}+\delta_{c_{1}c^{\prime}_{3}}+\delta_{c_{2}c^{\prime}_{3}}+\delta_{c_{2}c^{\prime}_{1}}+\delta_{c_{3}c^{\prime}_{1}}+\delta_{c_{3}c^{\prime}_{2}}\right].$ (58)
Within the QBCD approach, we follow the same approach as in the case of the Max Cut and determine with a mesh of solely $40$ points the minimal gap, $\Delta$ and the corresponding eigenvectors, $\lvert\mathrm{GS}\rangle$ and $\lvert\mathrm{ES}\rangle$ for the ground state and the first excited state, respectively. This method requires around two orders of magnitude fewer computational resources than simulating the annealing process. The corresponding QBCD Hamiltonian is written as
$H_{\mathrm{QBCD}}=-i\frac{\langle\mathrm{GS}\lvert\partial_{\lambda}H\rvert\mathrm{ES}\rangle}{\Delta}\left(\lvert\mathrm{ES}\rangle\langle\mathrm{GS}\rvert-\lvert\mathrm{GS}\rangle\langle\mathrm{ES}\rvert\right).$
Additionally, we also sparsified the full QBCD to the same density as that of the leading order local CD expansion, Eq. (V.3), reducing exponentially its non-locality and the computational resources of its implementation on quantum computers and in classical algorithms.
To this end, we use again the ground state gap and the final ground state fidelity as measures of adiabaticity. As shown in Fig. 16, even for short annealing times, $T=5$, the local expansion approach could only amplify the exponential decay of that ground state gap by a constant prefactor. On the other hand, the sparsified QBCD method achieves exponentially larger gap sizes with an approximately polynomial decay even for much longer annealing times, $T=10,20,80$. Its effects only vanish once adiabaticity sets in, already in the process without CD. Quite remarkably, the relative difference from the ground state gap of the bare process exhibits the same $L$-independent quadratic decay with respect to $T$ as in the case of the Max Cut, as demonstrated in Fig. 17. Similar trends are observed for the final fidelity, as shown in Fig. 18. The sparsified QBCD achieves significantly higher fidelity at short annealing times ($T=2$) than both the local expansion and the bare process, even in cases when the latter are run for much longer durations ($T=40$). At these longer time scales, the fidelity for QBCD remains close to unity, while the dynamics becomes essentially insensitive to the presence of the local CD expansion. In this regime, the local expansion together with the bare process still exhibits an exponentially decaying fidelity.
Figure 16: Ground state energy gap for the $3$-XORSAT model. Local CD expansion shows only negligible amplification even at short driving times (orange squares) $T=5$, compared to the bare process (red line). The sparsified QBCD exhibits an exponential amplification of the gap, as long as adiabaticity is not reached in the bare process.
Figure 17: Relative difference between the energy gaps without CD and with the first order ansatz decaying approximately as the minimal spin glass bottleneck independently of the system size. No essential gap amplification is present around the driving times, where even without CD, the fidelity departs from zero.
Figure 18: Ground-state fidelity for the $3$-XORSAT problem. The first-order local CD method yields a slower exponential decay at short driving times, but its advantage diminishes as the system approaches the adiabatic regime, where no significant improvement is observed. In contrast, the sparsified QBCD exponentially enhances the fidelity across all timescales, yielding finite values already at $T=2$ and approaching near-adiabatic fidelity for longer processes ($T=5,20,40$). Solid lines refer to QBCD, dashed lines to 1st order CD, dotted lines to no CD.
VI Discussion
In this study, we explored how quantum adiabatic dynamics can be enhanced to overcome spin-glass bottlenecks using various approximate counterdiabatic (CD) driving schemes. Specifically, we examined the variational local expansion, its improved version known as counterdiabatic optimized local driving (COLD). We also introduced a new scheme—quantum brachistochrone counterdiabatic driving (QBCD)—which focuses directly on the bottleneck region while relying only on the approximate knowledge of an exponentially small fraction of the full CD at a single parameter value. First, we tested the local expansion approaches on an exactly solvable minimal spin glass toy model described by an Ising chain admitting a free-fermionic description, first introduced in Ref. [99]. This model serves as a valuable testbed displaying both a ground state gap, exponentially small in the system size and two degenerate ground states that differ macroscopically from each other in terms of spin flips. The solvability of the model, however, was used only to reach larger system sizes and extract the asymptotic scalings: most of our analysis can be applied without substantial modifications to NP-hard Ising Hamiltonians. To this end, we chose to employ CD terms that can be constructed efficiently on a classical computer: we resorted to two- and three-body ansatzes with coefficients fixed from the Floquet-Krylov variational principle [58, 59]. A key finding is that, while these CD ansatzes can exponentially amplify the size of the minimal gap along the adiabatic path, such gap remains exponentially small in the system size $L$, and the probability of ending up in the ground state at the end of the protocol is correspondingly very small for times shorter than $\sim e^{\alpha L}$, with $\alpha$ given in Eq. (19).
Our analysis highlights key features of CD of frustrated spin systems. First, from the gap characterization in Sec. IV.1, it can be inferred that CD can amplify the gap exponentially in the diabatic regime, and thus substantially reduce the diabatic transitions. However, reaching the regime of driving times where diabatic transitions come solely from the exponentially small ground state gap, the first two CD terms provide negligible improvement (see Fig. 7) and higher order expansions are required. On the other hand, slowing down the driving schedule also reduces the overall strength of the CD term, given the $\partial_{t}\lambda$ prefactor in Eq. (21). These features prevent the construction of an effective CD term that ensures adiabatic dynamics in the presence of exponentially small gaps between frustrated energy levels.
In addition, CD techniques, including the variational formulation [58, 59, 60], aim at keeping the evolved state as close as possible to the instantaneous ground state. The simplicity of the model allows us to identify a single transition from one end to the other of the effective chain, Eq. (16). Thus, in crossing the first-order phase transition, the state needs to be modified extensively in order to adjust to the change in the ground state.
As an alternative and novel approach, we have introduced the QBCD Hamiltonian in Sec. V, which shares the structure of the Hamiltonian for the time-optimal evolution in the quantum brachistochrone problem for the associated subspace of localized edge states. The QBCD Hamiltonian employs only an exponentially small fraction of the full CD at a single parameter value, focusing exclusively on the ground and first excited states at the critical point. With the help of it, the success rate of the protocol increases from a vanishing probability to almost 0.5, as indicated by the average kink number at the end of the schedule. While implementing non-local, multi-body interactions on quantum hardware remains technically challenging, the structure of QBCD makes its realization feasible with current digital or digital-analog quantum simulation platforms. This feasibility arises from two key features: only an exponentially small fraction of the full counterdiabatic correction—restricted to the ground and first excited states—is required, and an approximate knowledge of this Hamiltonian at a single parameter value is sufficient. These properties substantially reduce the complexity of the implementation, placing experimental realization of QBCD within reach. Additionally, QBCD offers a practical advantage in terms of energetic cost: the Hilbert-Schmidt norm of the Hamiltonian remains finite due to the exponentially reduced contributions from higher-order multi-body terms.
In the second part of the paper, we probed the generality of our findings in two realistic spin glass models representing NP-hard problems, namely the $3$-regular Max Cut and the $3$-XORSAT problems. These models exhibit an exponentially small gap already at system sizes $L\gtrsim 10$, accessible by classical numerical simulations. Specifically, the $3$-XORSAT problem extends the validity of our results to cubic spin-glass Hamiltonians. Due to limitations in accessible system sizes, we used the final ground-state fidelity—rather than the excitation density—as a measure of adiabaticity, alongside the minimum gap. Our findings indicate that the local CD expansions (second-order expansion for Max Cut and first-order for $3$-XORSAT) display a level of effectiveness similar to that observed in the minimal spin-glass model. Only modest improvements were observed outside the adiabatic regime, in both the gap and the final fidelity. Meanwhile, for intermediate annealing times, where the fidelity departs from zero in the bare process without CD, no significant improvement was detected.
For the QBCD method, we introduced an enhanced, sparsified version that matches the density of the second-order local expansion. This sparsification significantly improves its prospects for implementations in both experimental setups and large-scale classical algorithms. Remarkably, even retaining only an exponentially small fraction of the already exponentially restricted non-local terms in the QBCD Hamiltonian yields a similar improvement in the adiabatic time scale as seen in the minimal spin-glass model.
Finally, we also compared the efficiency of the COLD approach with the local expansion and the sparsified QBCD in the Max Cut problem. Although it provided a noticeable improvement over the local expansion, its performance remained consistently inferior to that of the sparsified QBCD. In fact, even at small system sizes ($L=6$), spin-glass effects already degrade COLD’s effectiveness below that of QBCD. Moreover, the COLD method proved surprisingly difficult to implement numerically—even for modest system sizes ($L\lesssim 10$). Furthermore, as the system size increased and spin-glass effects became more pronounced, the performance of COLD gradually converged to that of the second-order local expansion. These findings underscore both the limited impact of COLD and the significant computational challenges compared to the QBCD method.
This combination of results suggests that the challenge of traversing exponentially small gaps in these spin-glass problems is primarily of entropic rather than energetic origin [118]. In other words, overcoming the bottleneck does not require access to strong control fields, but rather a minimal amount of information about where—and how—to steer the quantum state. This perspective further highlights the practical value of methods like QBCD, which leverage such targeted information with limited and implementable control resources.
Accordingly, while exploring alternative strategies beyond a priori action minimization remains an interesting direction, our results indicate that focusing on further improving the QBCD approach may ultimately be more effective for overcoming bottlenecks. Approaches based on integrating CD ansatzes with optimal control methods—such as QAOA [63, 62, 64, 65]—optimize control fields iteratively using only the final state reached by the dynamics [77]. However, these methods often suffer from limited analytical transparency and can be hampered by complex control landscapes and barren plateaus [119, 120, 121, 122]. In contrast, QBCD already demonstrates substantial improvements with minimal and experimentally realistic information. Given these advantages, it appears more promising to build upon QBCD and approximate CD strategies—potentially enhancing them with methods such as reverse annealing [104, 105, 106], bias fields [123, 87], Lyapunov control [67, 68], or targeted control engineering near critical points.
Acknowledgements.
It is a pleasure to thank Pranav Chandarana and Koushik Paul for comments on the manuscript and Oskar A. Prośniak and Vittorio Vitale for discussions. This project was supported by the Luxembourg National Research Fund (FNR Grant Nos. 17132054 and 16434093). It has also received funding from the QuantERA II Joint Programme and co-funding from the European Union’s Horizon 2020 research and innovation programme. The numerical simulations presented in this work were partly carried out using the HPC facilities of the University of Luxembourg.
Appendix A Free-fermion representation of the Hamiltonian
A.1 Jordan-Wigner transformation and Dirac fermions
Start from the Hamiltonian (1). By applying a set of one-qubit gates
$U=\frac{1}{\sqrt{2}}\begin{pmatrix}1&-1\\ 1&1\end{pmatrix},$ (60)
by virtue of the identities
$U\sigma^{x}U^{\dagger}=-\tilde{\sigma}^{z},\qquad U\sigma^{y}U^{\dagger}=\tilde{\sigma}^{y},\qquad U\sigma^{z}U^{\dagger}=\tilde{\sigma}^{x},$ (61)
one can rewrite equivalently
$H[\lambda]=\sum_{j=1}^{L}\left[-\lambda J_{j}\tilde{\sigma}^{x}_{j}\tilde{\sigma}^{x}_{j+1}+(1-\lambda)\tilde{\sigma}_{j}^{z}\right]$ (62)
(tildes are used to distinguish the two sets of Pauli matrices for later convenience). This form of the Hamiltonian is convenient for a Jordan-Wigner transformation to be applied: one passes from spins $\{\tilde{\sigma}_{j}\}$ to Dirac fermions $\{c_{j}^{\phantom{\dagger}},c_{j}^{\dagger}\}$ by means of the identities
$\tilde{\sigma}_{j}^{+}=e^{-i\pi\Sigma_{j}}c_{j}^{\dagger},\qquad\tilde{\sigma}_{j}^{-}=e^{i\pi\Sigma_{j}}c_{j},\qquad\tilde{\sigma}_{j}^{z}=2c_{j}^{\dagger}c_{j}^{\phantom{\dagger}}-1,$ (63)
where the string operator $\Sigma_{j}$ in the exponent reads
$\Sigma_{j}:=\sum_{k=1}^{j-1}c_{k}^{\dagger}c_{k}^{\phantom{\dagger}}.$ (64)
Another possible way of writing the transformation is
$c_{j}^{\dagger}=\prod_{k=1}^{j-1}\big{(}-\tilde{\sigma}_{k}^{z}\big{)}\,\tilde{\sigma}_{j}^{+},\qquad c_{j}=\prod_{k=1}^{j-1}\big{(}-\tilde{\sigma}_{k}^{z}\big{)}\,\tilde{\sigma}_{j}^{-}.$ (65)
Then, the fermionic vacuum |$\ket{0_{c}}$ corresponds to spins positively aligned along $x$-axis in the original basis (|$\ket{0_{c}}=\ket{\rightarrow}$), and the fully occupied fermionic state corresponds to the spins anti-aligned along $x$-axis (|$\ket{1_{c}}=\ket{\leftarrow}$).
Applying the Jordan-Wigner transformation to Eq. (62), one finds
$H[\lambda]$ $=\sum_{j=1}^{L}\left[-\lambda J_{j}\left(e^{-i\pi\Sigma_{j}}c_{j}^{\dagger}+e^{i\pi\Sigma_{j}}c_{j}\right)\left(e^{-i\pi\Sigma_{j+1}}c_{j+1}^{\dagger}+e^{i\pi\Sigma_{j+1}}c_{j+1}\right)+(1-\lambda)\left(2c_{j}^{\dagger}c_{j}^{\phantom{\dagger}}-1\right)\right]$ (66)
$=\sum_{j=1}^{L}\left[-\lambda J_{j}\left(c_{j}^{\dagger}e^{-i\pi c_{j}^{\dagger}c_{j}^{\phantom{\dagger}}}c_{j+1}^{\dagger}+c_{j}^{\dagger}e^{i\pi c_{j}^{\dagger}c_{j}^{\phantom{\dagger}}}c_{j+1}+c_{j}e^{-i\pi c_{j}^{\dagger}c_{j}^{\phantom{\dagger}}}c_{j+1}^{\dagger}+c_{j}e^{i\pi c_{j}^{\dagger}c_{j}^{\phantom{\dagger}}}c_{j+1}\right)+2(1-\lambda)c_{j}^{\dagger}c_{j}^{\phantom{\dagger}}\right]-(1-\lambda)L$ (67)
$=\sum_{j=1}^{L}\left[-\lambda J_{j}\left(c_{j}^{\dagger}c_{j+1}^{\dagger}+c_{j}^{\dagger}c_{j+1}^{\phantom{\dagger}}-c_{j}^{\phantom{\dagger}}c_{j+1}^{\dagger}-c_{j}c_{j+1}\right)+2(1-\lambda)c_{j}^{\dagger}c_{j}^{\phantom{\dagger}}\right]-(1-\lambda)L.$ (68)
The Hamiltonian is manifestly a quadratic form in the creation/annihilation operators. In order to fix the boundary conditions, one needs to look back at the original Hamiltonian (1). In terms of the spins, the boundary conditions are periodic: $\sigma_{j+L}\equiv\sigma_{j}$. However, for the fermionic operators, it holds $c_{j+L}=+c_{j}$ or $c_{j+L}=-c_{j}$ depending on whether there is an odd or even number of fermions, respectively. Since the initial state (ground state at $\lambda=0$) is the fermionic vacuum |$\ket{0_{c}}$, the correct sector to take is the even one.
In order to simplify things, we will not use the relation $c_{j+L}=-c_{j}$, but equivalently use $c_{j+L}=c_{j}$ and flip the sign of the last coupling $J_{L}\to-J_{L}$, while still working within the even parity sector. In order to keep track of this change, we use tildes:
$\tilde{J}_{j}:=\begin{cases}J_{j}&\text{if }j\neq L\\ -J_{L}&\text{if }j=L.\end{cases}$ (69)
Therefore, the Hamiltonian reads
$H[\lambda]$ $=\sum_{j=1}^{L}\left[-\lambda\tilde{J}_{j}\left(c_{j}^{\dagger}c_{j+1}^{\dagger}+c_{j}^{\dagger}c_{j+1}^{\phantom{\dagger}}-c_{j}^{\phantom{\dagger}}c_{j+1}^{\dagger}-c_{j}c_{j+1}\right)+2(1-\lambda)c_{j}^{\dagger}c_{j}^{\phantom{\dagger}}\right]-(1-\lambda)L$ (70)
$=\sum_{j=1}^{L}\left[-\lambda\tilde{J}_{j}\left(c_{j}^{\dagger}-c_{j}^{\phantom{\dagger}}\right)\left(c_{j+1}^{\dagger}+c_{j+1}^{\phantom{\dagger}}\right)+2(1-\lambda)c_{j}^{\dagger}c_{j}^{\phantom{\dagger}}\right]-(1-\lambda)L.$ (71)
Notice that if any operator (different from the Hamiltonian) has terms coupling $c_{L}$ with $c_{L+1}$, also there the sign must be flipped. We will use tildes to remember this fact.
A.2 Majorana fermions
From the Dirac fermion representation $\{c_{i}^{\phantom{\dagger}},c_{j}^{\dagger}\}_{j=1}^{L}$, one can pass to an equivalent Majorana fermionic representation $\{\gamma_{2j-1},\gamma_{2j}\}_{j=1}^{L}$ by effectively doubling the length of the chain. The defining relations are
$\gamma_{2j-1}:=\frac{1}{\sqrt{2}}\left(c_{j}^{\phantom{\dagger}}+c_{j}^{\dagger}\right),\qquad\gamma_{2j}:=\frac{i}{\sqrt{2}}\left(c_{j}^{\phantom{\dagger}}-c_{j}^{\dagger}\right),$ (72)
and the inverse transformation is
$c_{j}^{\dagger}=\frac{\gamma_{2j-1}+i\gamma_{2j}}{\sqrt{2}},\qquad c_{j}^{\phantom{\dagger}}=\frac{\gamma_{2j-1}-i\gamma_{2j}}{\sqrt{2}},$ (73)
leading to anticommutation relations
$\left\{\gamma_{k},\gamma_{k^{\prime}}\right\}=\delta_{k,k^{\prime}}.$ (74)
One can rewrite
$H[\lambda]=-2i\sum_{j=1}^{L}\big{[}\lambda\tilde{J}_{j}\gamma_{2j}\gamma_{2j+1}+(1-\lambda)\gamma_{2j-1}\gamma_{2j}\big{]}.$ (75)
The boundary conditions inherited by the Dirac fermions entail $\gamma_{k+2L}=-\gamma_{k}$ (even parity sector) but, as stated above, we are employing $\gamma_{k+2L}=\gamma_{k}$ and the flipped couplings $\tilde{J}_{j}$.
A.3 Reflection parity symmetry
The Hamiltonian (1), and thus its Majorana representation (75), is odd under the chiral reflection parity symmetry $\widetilde{\Pi}_{\mathrm{R}}$ defined by Eq. (10), i.e. $\widetilde{\Pi}_{\mathrm{R}}\gamma_{k}\widetilde{\Pi}_{\mathrm{R}}=\gamma_{2L-k+1}$. Indeed, the couplings satisfy $J_{j}=J_{L-j}$, and one can write
$\widetilde{\Pi}_{\mathrm{R}}H\widetilde{\Pi}_{\mathrm{R}}$ $=-2i\sum_{j=1}^{L}\big{[}\lambda\tilde{J}_{j}\gamma_{2L-2j+1}\gamma_{2L-2j}+(1-\lambda)\gamma_{2L-2j+2}\gamma_{2L-2j+1}\big{]}$ (76)
$=-2i\sum_{j=1}^{L}\big{[}\lambda\tilde{J}_{L-j}\gamma_{2L-2j+1}\gamma_{2L-2j}+(1-\lambda)\gamma_{2L-2j+2}\gamma_{2L-2j+1}\big{]}$ (77)
$=-2i\sum_{l=0}^{L-1}\big{[}\lambda\tilde{J}_{l}\gamma_{2l+1}\gamma_{2l}+(1-\lambda)\gamma_{2l+2}\gamma_{2l+1}\big{]}=-H.$ (78)
In the first line, the operator $\widetilde{\Pi}_{\mathrm{R}}$ was applied on the $\gamma$’s; in the second line, $\tilde{J}_{j}$ was exchanged with $\tilde{J}_{L-j}$; in the third line, it was set $l:=L-j$. $H$ is odd under the reflection parity $\widetilde{\Pi}_{\mathrm{R}}$, thus $\widetilde{\Pi}_{\mathrm{R}}$ is a chiral symmetry for $H$ [100].
Before moving on, it is interesting to express the chiral reflection parity symmetry also in terms of the Dirac operators:
$\widetilde{\Pi}_{\mathrm{R}}c_{j}\widetilde{\Pi}_{\mathrm{R}}$ $=\widetilde{\Pi}_{\mathrm{R}}\frac{\gamma_{2j-1}-i\gamma_{2j}}{\sqrt{2}}\widetilde{\Pi}_{\mathrm{R}}=\frac{\gamma_{2L-2j+2}-i\gamma_{2L-2j+1}}{\sqrt{2}}=-ic^{\dagger}_{L-j+1},$ (79a)
$\widetilde{\Pi}_{\mathrm{R}}c_{j}^{\dagger}\widetilde{\Pi}_{\mathrm{R}}$ $=\widetilde{\Pi}_{\mathrm{R}}\frac{\gamma_{2j-1}+i\gamma_{2j}}{\sqrt{2}}\widetilde{\Pi}_{\mathrm{R}}=\frac{\gamma_{2L-2j+2}+i\gamma_{2L-2j+1}}{\sqrt{2}}=ic_{L-j+1},$ (79b)
which shows that $\widetilde{\Pi}_{\mathrm{R}}$ also exchanges particles and holes. In terms of the spin operators, it holds instead
$\widetilde{\Pi}_{\mathrm{R}}\sigma_{j}^{x}\widetilde{\Pi}_{\mathrm{R}}=-\widetilde{\Pi}_{\mathrm{R}}\tilde{\sigma}_{j}^{z}\widetilde{\Pi}_{\mathrm{R}}=-\widetilde{\Pi}_{\mathrm{R}}c_{j}^{\dagger}c_{j}^{\phantom{\dagger}}\widetilde{\Pi}_{\mathrm{R}}=-ic_{L-j+1}(-i)c_{L-j+1}^{\dagger}=c_{L-j+1}^{\dagger}c_{L-j+1}=\tilde{\sigma}_{L-j+1}^{z}=-\sigma_{L-j+1}^{x},$ (80)
while $\widetilde{\Pi}_{\mathrm{R}}\sigma_{j}^{z}\widetilde{\Pi}_{\mathrm{R}}$ involves also the transformation of the Jordan-Wigner string. Thus, the chiral reflection parity symmetry is non-local in the spin representation of the Hamiltonian.
The non-locality of $\widetilde{\Pi}_{\mathrm{R}}$ in the spin representation is “dual” to the non-locality of $\Pi_{\mathrm{R}}$ in the fermionic representation. Indeed, the presence of the Jordan-Wigner string entails the transformation rules
$\Pi_{\mathrm{R}}c_{j}\Pi_{\mathrm{R}}$ $=\Pi_{\mathrm{R}}\prod_{k=1}^{j-1}\big{(}-\sigma_{k}^{z}\big{)}\sigma_{j}^{-}\Pi_{\mathrm{R}}=\prod_{k=1}^{j-1}\big{(}-\sigma_{L-k+1}^{z}\big{)}\sigma_{L-j+1}^{-}=(-1)^{N_{\mathrm{F}}}\prod_{k=1}^{L-j+1}\big{(}-\sigma_{k}^{z}\big{)}\sigma_{L-j+1}^{-}$
$=(-1)^{N_{\mathrm{F}}}\left(1-2c_{L-j+1}^{\dagger}c_{L-j+1}^{\phantom{\dagger}}\right)c_{L-j+1}=(-1)^{N_{\mathrm{F}}}c_{L-j+1},$ (81a)
$\Pi_{\mathrm{R}}c_{j}^{\dagger}\Pi_{\mathrm{R}}$ $=\Pi_{\mathrm{R}}\prod_{k=1}^{j-1}\big{(}-\sigma_{k}^{z}\big{)}\sigma_{j}^{+}\Pi_{\mathrm{R}}=\prod_{k=1}^{j-1}\big{(}-\sigma_{L-k+1}^{z}\big{)}\sigma_{L-j+1}^{+}=(-1)^{N_{\mathrm{F}}}\prod_{k=1}^{L-j+1}\big{(}-\sigma_{k}^{z}\big{)}\sigma_{L-j+1}^{+}$
$=(-1)^{N_{\mathrm{F}}}\left(1-2c_{L-j+1}^{\dagger}c_{L-j+1}^{\phantom{\dagger}}\right)c_{L-j+1}^{\dagger}=-(-1)^{N_{\mathrm{F}}}c_{L-j+1}^{\dagger},$ (81b)
where $N_{\mathrm{F}}:=\sum_{j=1}^{L}c_{j}^{\dagger}c_{j}^{\phantom{\dagger}}$ is the total fermion number operator. The prefactor $(-1)^{N_{\mathrm{F}}}$ is a non-local operator that makes it more difficult to define creation and annihilation operators symmetrized with respect to $\Pi_{\mathrm{R}}$. This is the reason why it is convenient to use the chiral symmetry $\widetilde{\Pi}_{\mathrm{R}}$ instead of the unitary symmetry $\Pi_{\mathrm{R}}$ when working with the Dirac fermions. A similar situation takes place in the Majorana representation. A comprehensive, modern view on the subtleties of symmetries in Ising-like chains can be found in Ref. [124].
A.4 Reflection-symmetrized representation
Let us introduce symmetrized operators, as in Eqs. (12) in the main text. The inverse transformation is
$\gamma_{k}$ $=\frac{\Gamma_{k}^{+}-\Gamma_{k}^{-}}{\sqrt{2}i},\qquad$ $k=1,3,\dots,L,$ (82a)
$\gamma_{k}$ $=\frac{\Gamma_{k}^{+}+\Gamma_{k}^{-}}{\sqrt{2}},\qquad$ $k=2,4,\dots,L-1,$ (82b)
$\gamma_{k}$ $=\frac{\Gamma_{2L-k+1}^{+}+\Gamma_{2L-k+1}^{-}}{\sqrt{2}},\qquad$ $k=L+1,L+3,\dots,2L,$ (82c)
$\gamma_{k}$ $=\frac{\Gamma_{2L-k+1}^{+}-\Gamma_{2L-k+1}^{-}}{\sqrt{2}i},\qquad$ $k=L+2,L+3,\dots,2L-1.$ (82d)
Let us split the Hamiltonian (75) into the different contributions: for $-\lambda J$ or $(1-\lambda)$, and bulk or boundary. It holds for $j=1,2,\dots,\ell$
$\tilde{J}_{j}\gamma_{2j}\gamma_{2j+1}+\tilde{J}_{L-j}\gamma_{2L-2j}\gamma_{2L-2j+1}$ $=\tilde{J}_{j}\left[\frac{\Gamma_{2j}^{+}+\Gamma_{2j}^{-}}{\sqrt{2}}\frac{\Gamma_{2j+1}^{+}-\Gamma_{2j+1}^{-}}{\sqrt{2}i}+\frac{\Gamma_{2j+1}^{+}+\Gamma_{2j+1}^{-}}{\sqrt{2}}\frac{\Gamma_{2j}^{+}-\Gamma_{2j}^{-}}{\sqrt{2}i}\right]$
$=-\frac{i\tilde{J}_{j}}{2}\left[\left(\Gamma_{2j}^{+}+\Gamma_{2j}^{-}\right)\left(\Gamma_{2j+1}^{+}-\Gamma_{2j+1}^{-}\right)-\left(\Gamma_{2j}^{+}-\Gamma_{2j}^{-}\right)\left(\Gamma_{2j+1}^{+}+\Gamma_{2j+1}^{-}\right)\right]$
$=i\tilde{J}_{j}\left(\Gamma_{2j}^{+}\Gamma_{2j+1}^{-}-\Gamma_{2j}^{-}\Gamma_{2j+1}^{+}\right),$ (83)
and
$(1-\lambda)(\gamma_{2j-1}\gamma_{2j}+\gamma_{2L-2j-1}\gamma_{2L-2j})$ $=(1-\lambda)\left[\frac{\Gamma_{2j-1}^{+}-\Gamma_{2j-1}^{-}}{\sqrt{2}i}\frac{\Gamma_{2j}^{+}+\Gamma_{2j}^{-}}{\sqrt{2}}+\frac{\Gamma_{2j}^{+}-\Gamma_{2j}^{-}}{\sqrt{2}i}\frac{\Gamma_{2j-1}^{+}+\Gamma_{2j-1}^{-}}{\sqrt{2}}\right]$
$=-\frac{i(1-\lambda)}{2}\left[\left(\Gamma_{2j-1}^{+}-\Gamma_{2j-1}^{-}\right)\left(\Gamma_{2j}^{+}+\Gamma_{2j}^{-}\right)-\left(\Gamma_{2j-1}^{+}+\Gamma_{2j-1}^{-}\right)\left(\Gamma_{2j}^{+}-\Gamma_{2j}^{-}\right)\right]$
$=i(1-\lambda)\left(\Gamma_{2j-1}^{-}\Gamma_{2j}^{+}-\Gamma_{2j-1}^{+}\Gamma_{2j}^{-}\right),$ (84)
while for the other sites
$\tilde{J}_{L}\gamma_{2L}\gamma_{1}=\tilde{J}_{L}\frac{\Gamma_{1}^{+}+\Gamma_{1}^{-}}{\sqrt{2}}\frac{\Gamma_{1}^{+}-\Gamma_{1}^{-}}{\sqrt{2}i}=\frac{i\tilde{J}_{L}}{2}\left(\Gamma_{1}^{+}\Gamma_{1}^{-}-\Gamma_{1}^{-}\Gamma_{1}^{+}\right),$ (85)
$(1-\lambda)\gamma_{L}\gamma_{L+1}=(1-\lambda)\frac{\Gamma_{L}^{+}-\Gamma_{L}^{-}}{\sqrt{2}i}\frac{\Gamma_{L}^{+}+\Gamma_{L}^{-}}{\sqrt{2}}=\frac{i}{2}(1-\lambda)\left(\Gamma_{L}^{-}\Gamma_{L}^{+}-\Gamma_{L}^{+}\Gamma_{L}^{-}\right).$ (86)
The Hamiltonian acquires the form
$H[\lambda]=2\lambda\sum_{j=1}^{\ell}\tilde{J}_{j}\left(\Gamma_{2j}^{+}\Gamma_{2j+1}^{-}-\Gamma_{2j}^{-}\Gamma_{2j+1}^{+}\right)+2(1-\lambda)\sum_{j=1}^{\ell}\left(\Gamma_{2j-1}^{-}\Gamma_{2j}^{+}-\Gamma_{2j-1}^{+}\Gamma_{2j}^{-}\right)\\ +\lambda\tilde{J}_{L}\left(\Gamma_{1}^{+}\Gamma_{1}^{-}-\Gamma_{1}^{-}\Gamma_{1}^{+}\right)+(1-\lambda)\left(\Gamma_{L}^{-}\Gamma_{L}^{+}-\Gamma_{L}^{+}\Gamma_{L}^{-}\right),$ (87)
which is the same of Eq. (14) in the main text.
A.5 Transformation of states
From the equations above, one can also find the relation between the vacuum state for the $c$’s and the one for the $\Gamma$’s. Defining |$\ket{0_{c}}$ to be the state annihilated by all $c_{j}$’s, and |$\ket{0_{\Gamma}}$ to be the one annihilated by all the $\Gamma_{k}^{-}$’s, it holds
|$\ket{0_{c}}=\prod_{k=1}^{\ell}\left(1+\Gamma_{2k-1}^{+}\Gamma_{2k}^{+}\right)\ket{0_{\Gamma}},$ (88)
as one can check by directly applying each $c_{j}$ expressed in terms of the $\Gamma_{k}^{\pm}$’s.
Appendix B Diagonalization of the effective hopping model
The purpose of this Appendix is to provide the eigendecomposition of the system Hamiltonian (1) by diagonalizing the effective 1D hopping model Eq. (16). It is convenient to rewrite the matrix as
$\mathcal{H}=\lambda\begin{pmatrix}-J_{L}&-B&&&&&\\ -B&0&J_{1}&&&&\\ &J_{1}&0&-B&&&\\ &&-B&0&J_{2}&&\\ &&&J_{2}&\ddots&\ddots&\\ &&&&\ddots&0&J_{\ell}\\ &&&&&J_{\ell}&-B\end{pmatrix},$ (89)
where $B:=(1-\lambda)/\lambda$.
As a first thing, notice that for the choice of $J_{j}$’s in Eq. (2), the model is translation invariant in bulk (step 2), while only the sites $1$, $L-1$ and $L$ experience different couplings. This fact suggests that plane waves are eigenfunctions of the system, provided the boundary conditions are fixed.
[…]
B.1 Plane waves in the bulk
Let us start by constructing the plane wave eigenvectors of Eq. (89) in the bulk. The eigenvalue equation $\mathcal{H}|\psi)=\lambda\epsilon|\psi)$ reads, in components,
$\begin{cases}-B\psi_{2j}+\psi_{2j-2}=\epsilon\psi_{2j-1}\\ \psi_{2j+1}-B\psi_{2j-1}=\epsilon\psi_{2j},\end{cases}$ (90)
where the couplings $J_{j}$ of Eq. (2) have been specified. Using the Fourier ansatz
$\begin{pmatrix}\psi_{2j-1}\\ \psi_{2j}\end{pmatrix}\sim e^{ikj}\begin{pmatrix}a_{k}\\ b_{k}\end{pmatrix},$ (91)
the bulk eigenvalue equation acquires the form
$\begin{cases}-Bb_{k}+e^{-ik}b_{k}=\epsilon_{k}a_{k}\\ e^{ik}a_{k}-Ba_{k}=\epsilon_{k}b_{k}.\end{cases}$ (92)
From the linear system above, one determines
$(e^{-ik}-B)(e^{ik}-B)=\epsilon^{2}_{k}$ (93)
and
$\frac{a_{k}}{b_{k}}=\frac{\epsilon_{k}}{e^{ik}-B}.$ (94)
The first equation fixes the dispersion relation in the bulk, and the second the relative scale of $a_{k}$ to $b_{k}$. They need to be supplemented by the boundary conditions, which fix the admissible values of $k$, and the normalization condition, which fixes the overall scale of the eigenvector. In order to simplify things, it is instrumental to work already in the thermodynamic limit $L\to\infty$, since this way, the left and right boundary conditions are completely decoupled. The two will be considered separately in the following subsections.
B.2 Left-localized state
Here, consider a semi-infinite chain starting from $j=0$ and ending at $j=+\infty$. The only boundary condition one needs to take into account is provided by the eigenvalue equation at the first site:
$J^{\prime}\psi_{1}-B\psi_{2}=\epsilon\psi_{1},$ (95)
that can be rewritten as
$J^{\prime}a_{k}-Bb_{k}=\epsilon_{k}a_{k}\qquad\implies\qquad\frac{a_{k}}{b_{k}}=\frac{B}{J^{\prime}-\epsilon_{k}}.$ (96)
When supplemented with Eqs. (93), (94) and the normalization condition, one obtains four equations for the four unknowns $a_{k}$,$b_{k}$,$k$,$\epsilon_{k}$. Here, we consider only the localized edge state, for which $k\equiv i\kappa$, $\kappa>0$. Eliminating $a_{k}/b_{k}$ and $\epsilon_{k}$ from Eq. (96) via Eqs. (93) and (94), one finds
$\frac{1-e^{\kappa}B}{1-e^{-\kappa}B}=(J^{\prime})^{2},$ (97)
which has a real positive solution
$\kappa=\ln\frac{1-(J^{\prime})^{2}+\sqrt{4B^{2}(J^{\prime})^{2}+[(J^{\prime})^{2}-1]^{2}}}{2B}.$ (98)
B.3 Right-localized state
Here, consider the half-infinite chain that starts at $j=-\infty$ and ends at $j=L$. Besides Eqs. (93) and (94) that describe the wave propagation in the bulk, the two boundary conditions at the right end read
$-B\psi_{L-2}+J_{\ell}\psi_{L}=\epsilon\psi_{L-1},\qquad J_{\ell}\psi_{L-1}-B\psi_{L}=\epsilon\psi_{L}.$ (99)
Being the last site odd, the Fourier ansatz Eq. (91) needs to be supplemented by
$\psi_{L}\sim e^{ikL}c_{k}.$ (100)
The boundary conditions Eq. (99) read, in terms of the Fourier ansatz,
$-Ba_{k}+Je^{ik}c_{k}=\epsilon_{k}b_{k},\qquad Jb_{k}e^{-ik}-Bc_{k}=\epsilon_{k}c_{k}.$ (101)
From here, one can eliminate $c_{k}$, obtaining
$\epsilon_{k}b_{k}+Ba_{k}=\frac{J^{2}b_{k}}{\epsilon_{k}+B}.$ (102)
Now, Eqs. (93), (94), (102) and the normalization condition are again four equations for the four unknowns $a_{k}$,$b_{k}$,$k$,$\epsilon_{k}$. The right-localized state will have an imaginary momentum $k\equiv-i\kappa$, $\kappa>0$, that can be determined by combining the cited equations. Since this would lead to cumbersome equations, it is convenient to utilize a workaround to determine the avoided crossing directly, as in Ref. [99].
B.4 Avoided crossing
The avoided crossing takes place at the value of $B$ such that $\epsilon_{k}$ is the same for the left ($\mathrm{L}$) and right ($\mathrm{R}$) localized states. In principle, one should impose $(\epsilon_{k})_{\mathrm{L}}\equiv(\epsilon_{k})_{\mathrm{R}}$ by using the expression (93) with the values of $\kappa_{\mathrm{L}}$ and $\kappa_{\mathrm{R}}$ determined above. However, it is convenient to notice that, because of Eq. (93), it must hold $\kappa_{\mathrm{L}}=\kappa_{\mathrm{R}}$, and from Eq. (94) follows $(a_{k}/b_{k})_{\mathrm{L}}=(b_{k}/a_{k})_{\mathrm{R}}$ (being $k_{\mathrm{R}}=-k_{\mathrm{L}}$ for the choice of $i$’s). Thus, one determines $(a_{k}/b_{k})_{\mathrm{L}}$ from Eq. (96), and equates it to $(b_{k}/a_{k})_{\mathrm{R}}$ from Eq. (102), finding
$\epsilon+B=\frac{J^{2}}{J^{\prime}}.$ (103)
Plugging above $\epsilon$ from Eq. (93), with $\kappa$ taken from Eq. (98), one determines the value of $B$ at which the crossing takes place:
$B_{\mathrm{c}}=\frac{(1-J^{2})[J^{2}-(J^{\prime})^{2}]}{J^{\prime}[1-2J^{2}+(J^{\prime})^{2}]}.$ (104)
From Eq. (103) follows
$\epsilon_{\mathrm{c}}=\frac{(J^{\prime})^{2}-J^{4}}{J^{\prime}[1-2J^{2}+(J^{\prime})^{2}]},$ (105)
and in turn from Eq. (93)
$\kappa_{\mathrm{c}}=\ln\frac{J^{\prime}(1-J^{2})}{J^{2}-(J^{\prime})^{2}}.$ (106)
From the equation above, the condition $\kappa>0$ implies for the couplings
$J^{2}<J^{\prime},$ (107)
which explains the requirement stated in Eq. (3).
The last step consists in determining the magnitude of the gap at the avoided level crossing. Since $|\psi_{\mathrm{L}})$ and $|\psi_{\mathrm{R}})$ determined above are not eigenfunctions of the chain of finite length, one can argue that
$\Delta_{\min}\simeq(\psi_{\mathrm{L}}|\mathcal{H}|\psi_{\mathrm{R}})\sim\sum_{j=1}^{L}e^{-\kappa_{c}j/2}e^{-\kappa_{c}(L-j)/2}\sim e^{-\kappa_{c}L/2},$ (108)
which corresponds to Eq. (19) in the main text. Indeed, the hybridization of $|\psi_{\mathrm{L}})$ and $|\psi_{\mathrm{R}})$ is approximately given by the off-diagonal matrix element $(\psi_{\mathrm{L}}|\mathcal{H}|\psi_{\mathrm{R}})$ in the reduced 2$\times$2 subspace spanned by $|\psi_{\mathrm{L}})$ and $|\psi_{\mathrm{R}})$ themselves. All the factors dropped in Eq. (108) are subleading with respect to the exponential term.
Appendix C Free-fermion representation of the counterdiabatic Hamiltonian
C.1 Chiral reflection parity of the CD Hamiltonian
As a first thing, let us investigate how the reflection parity symmetry acts on the CD Hamiltonian, Eq. (24). Using the fact that $\widetilde{\Pi}_{\mathrm{R}}H\widetilde{\Pi}_{\mathrm{R}}=-H$, as proven in Sec. A.3, it follows
[…]
having set $y:=-x$. Thus, $H_{1}$ is even under the chiral reflection parity, Eq. (10), while $H$ is odd. The same conclusion can be reached by noticing that in Eq. (25), there is an even number of $H$’s at every order of the expansion.
At the same time, $H_{1}$ is even under the standard reflection parity, Eq. (5), as can be checked directly from the spin representation.
C.2 First order ansatz
Let us move to the Hamiltonian (28). By applying the set of one-qubit gates defined in Eq. (60), one can rewrite
$H_{1}^{(1)}[\lambda]$ $=\sum_{j=1}^{L}\alpha_{j}\left(\tilde{\sigma}_{j}^{y}\tilde{\sigma}_{j+1}^{x}+\tilde{\sigma}_{j}^{x}\tilde{\sigma}_{j+1}^{y}\right)$ (112)
$=2i\sum_{j=1}^{L}\alpha_{j}\left(\tilde{\sigma}_{j}^{-}\tilde{\sigma}_{j+1}^{-}-\tilde{\sigma}_{j}^{+}\tilde{\sigma}_{j+1}^{+}\right).$ (113)
This form of the CD Hamiltonian is expressed in terms of Dirac fermions, Eq. (63), as
$H_{1}^{(1)}[\lambda]$ $=-2i\sum_{j=1}^{L}\tilde{\alpha}_{j}\left(e^{-i\pi\Sigma_{j}}c_{j}^{\dagger}e^{-i\pi\Sigma_{j+1}}c_{j+1}^{\dagger}\right)+\mathrm{h.c.}$ (114)
$=-2i\sum_{j=1}^{L}\tilde{\alpha}_{j}\left(c_{j}^{\dagger}e^{-i\pi c^{\dagger}_{j}c^{\phantom{\dagger}}_{j}}c_{j+1}^{\dagger}\right)+\mathrm{h.c.}$ (115)
$=-2i\sum_{j=1}^{L}\tilde{\alpha}_{j}\left[c^{\dagger}_{j}c^{\dagger}_{j+1}+c_{j}c_{j+1}\right],$ (116)
where there is a tilde on the $\alpha_{j}$’s because of the fermion parity symmetry, see above Eq. (69).
Next, one has to rewrite the Hamiltonian above in terms of Majorana fermions, Eq. (73):
$H_{1}^{(1)}[\lambda]$ $=-2i\sum_{j=1}^{L}\tilde{\alpha}_{j}\left[\frac{1}{2}(\gamma_{2j-1}+i\gamma_{2j})(\gamma_{2j+1}+i\gamma_{2j+2})+\frac{1}{2}(\gamma_{2j-1}-i\gamma_{2j})(\gamma_{2j+1}-i\gamma_{2j+2})\right]$ (117)
$=2i\sum_{j=1}^{L}\tilde{\alpha}_{j}\left[\gamma_{2j}\gamma_{2j+2}-\gamma_{2j-1}\gamma_{2j+1}\right].$ (118)
It is convenient to check the reflection parity of the expression above for $H_{1}$:
$\widetilde{\Pi}_{\mathrm{R}}H_{1}^{(1)}[\lambda]\widetilde{\Pi}_{\mathrm{R}}$ $=2i\sum_{j=1}^{L}\tilde{\alpha}_{j}\left[\gamma_{2L-2j+1}\gamma_{2L-2j-1}-\gamma_{2L-2j+2}\gamma_{2L-2j}\right]$ (119)
$=2i\sum_{j=1}^{L}\tilde{\alpha}_{L-j}\left[\gamma_{2j+1}\gamma_{2j-1}-\gamma_{2j+2}\gamma_{2j}\right]$ (120)
$=2i\sum_{j=1}^{L}\tilde{\alpha}_{L-j}\left[\gamma_{2j}\gamma_{2j+2}-\gamma_{2j-1}\gamma_{2j+1}\right],$ (121)
where in the first line Eq. (10) was used, and in the second it was changed $j\to L-j$. In order to have $\widetilde{\Pi}_{\mathrm{R}}H_{1}\widetilde{\Pi}_{\mathrm{R}}=H_{1}$, it must hold $\alpha_{j}\equiv\alpha_{L-j}$, which indeed follows from the variational equations (29).
At this point, one can pass to the symmetrized operators $\Gamma^{\pm}$. Using
$-\tilde{\alpha}_{j}\gamma_{2j-1}\gamma_{2j+1}+\tilde{\alpha}_{L-j}\gamma_{2L-2j}\gamma_{2L-2j+2}$ $=-\tilde{\alpha}_{j}\left(\Gamma_{2j-1}^{-}\Gamma_{2j+1}^{+}+\Gamma_{2j-1}^{+}\Gamma_{2j+1}^{-}\right),\qquad$ $j=1,2,\dots,\ell,$ (122a)
$\tilde{\alpha}_{j}\gamma_{2j}\gamma_{2j+2}-\tilde{\alpha}_{L-j}\gamma_{2L-2j-1}\gamma_{2L-2j+1}$ $=\tilde{\alpha}_{j}\left(\Gamma_{2j}^{+}\Gamma_{2j+2}^{-}+\Gamma_{2j}^{-}\Gamma_{2j+2}^{+}\right),$ $j=1,2,\dots,\ell-1,$ (122b)
$\tilde{\alpha}_{\ell}\gamma_{2\ell}\gamma_{2\ell+2}-\tilde{\alpha}_{\ell+1}\gamma_{2\ell+1}\gamma_{2\ell+3}$ $=\tilde{\alpha}_{\ell}\left(\Gamma_{2\ell}^{-}\Gamma_{2\ell+1}^{+}+\Gamma_{2\ell}^{+}\Gamma_{2\ell+1}^{-}\right),$ (122c)
$\tilde{\alpha}_{L}\left(-\gamma_{2L-1}\gamma_{1}+\gamma_{2L}\gamma_{2}\right)$ $=\tilde{\alpha}_{L}\left(\Gamma_{1}^{+}\Gamma_{2}^{-}+\Gamma_{1}^{-}\Gamma_{2}^{+}\right),$ (122d)
it follows
$H_{1}^{(1)}[\lambda]=2i\sum_{j=1}^{\ell}\tilde{\alpha}_{j}\left(\Gamma_{2j+1}^{+}\Gamma_{2j-1}^{-}-\Gamma_{2j-1}^{+}\Gamma_{2j+1}^{-}\right)+2i\sum_{j=1}^{\ell-1}\tilde{\alpha}_{j}\left(\Gamma_{2j}^{+}\Gamma_{2j+2}^{-}-\Gamma_{2j+2}^{+}\Gamma_{2j}^{-}\right)\\ +2i\tilde{\alpha}_{\ell}\left(\Gamma_{2\ell}^{+}\Gamma_{2\ell+1}^{-}-\Gamma_{2\ell+1}^{+}\Gamma_{2\ell}^{-}\right)+2i\tilde{\alpha}_{L}\left(\Gamma_{1}^{+}\Gamma_{2}^{-}-\Gamma_{2}^{+}\Gamma_{1}^{-}\right).$ (123)
The matrix representation reads (for $\ell=3$, thus $L=7$)
$H_{1}^{(1)}[\lambda]=2\,\vec{\Gamma}^{+}\begin{pmatrix}0&-i\alpha_{7}&-i\alpha_{1}&&&&\\ i\alpha_{7}&0&0&i\alpha_{1}&&&\\ i\alpha_{1}&0&0&0&-i\alpha_{2}&&\\ &-i\alpha_{1}&0&0&0&i\alpha_{2}&\\ &&i\alpha_{2}&0&0&0&-i\alpha_{3}\\ &&&-i\alpha_{2}&0&0&i\alpha_{3}\\ &&&&i\alpha_{3}&-i\alpha_{3}&0\\ \end{pmatrix}\vec{\Gamma}^{-}.$ (124)
C.3 Second order ansatz
When considered together the first and second-order approximate CD Hamiltonians $H_{1}^{(1)}+H_{1}^{(2)}$, one can check that the variational equations, obtained from the minimization of the action Eq. (26), are
$\left[8(1-\lambda)^{2}+\lambda^{2}\left(J_{j-1}^{2}+2J_{j}^{2}+J_{j+1}^{2}\right)\right]\alpha_{j}+2\lambda^{2}J_{j}\left(J_{j-1}\alpha_{j-1}+J_{j+1}\alpha_{j+1}\right)+4\lambda(\lambda-1)(\beta_{j-1}J_{j-1}+\beta_{j}J_{j+1})=-J_{j},$
$\left[8(1-\lambda)^{2}+\lambda^{2}\left(J_{j-1}^{2}+J_{j}^{2}+J_{j+1}^{2}+J_{j+2}^{2}\right)\right]\beta_{j}+2\lambda^{2}\left(J_{j-1}J_{j+1}\beta_{j-1}+J_{j}J_{j+2}\beta_{j+1}\right)+4\lambda(\lambda-1)(\alpha_{j+1}J_{j}+\alpha_{j}J_{j+1})=0,$ (125)
Thanks to the relation $J_{j}=J_{L-j}$, these equations are symmetric under the exchange
[…]
One can check that the second order ansatz for the CD Hamiltonian, Eq. (30), is expressed in terms of Dirac fermions as
$H_{1}^{(2)}[\lambda]=-2i\sum_{j=1}^{L}\tilde{\beta}_{j}\left[c^{\dagger}_{j}c^{\dagger}_{j+2}+c_{j}c_{j+2}\right].$ (127)
Indeed, the same computation of App. C.2 applies, and the intermediate operator $\sigma_{j+1}^{x}$ just cancels the Jordan-Wigner string. The tilde on the $\beta_{j}$’s is again to remind that the sign in front of any operator that connects the last sites of the chain to the firsts needs to be flipped. Contrary to the cases above, where only the last coupling $J_{L}$ or $\alpha_{L}$ needed to be flipped, here also $\beta_{L-1}$ needs to.
Proceeding in the same fashion of App. C.2, one finds
$H_{1}^{(2)}[\lambda]=2i\sum_{j=1}^{L}\tilde{\beta}_{j}\left[\gamma_{2j}\gamma_{2j+4}-\gamma_{2j-1}\gamma_{2j+3}\right]$ (128)
and
$H_{1}^{(2)}[\lambda]=2i\sum_{j=1}^{\ell-1}\tilde{\beta}_{j}\left(\Gamma_{2j+3}^{+}\Gamma_{2j-1}^{-}-\Gamma_{2j-1}^{+}\Gamma_{2j+3}^{-}\right)+2i\sum_{j=1}^{\ell-2}\tilde{\beta}_{j}\left(\Gamma_{2j}^{+}\Gamma_{2j+4}^{-}-\Gamma_{2j+4}^{+}\Gamma_{2j}^{-}\right)\\ +2i\tilde{\beta}_{\ell-1}\left(\Gamma_{2\ell-2}^{+}\Gamma_{2\ell+1}^{-}-\Gamma_{2\ell+1}^{+}\Gamma_{2\ell-2}^{-}\right)+2i\tilde{\beta}_{\ell}\left(\Gamma_{2\ell}^{+}\Gamma_{2\ell-1}^{-}-\Gamma_{2\ell-1}^{+}\Gamma_{2\ell}^{-}\right)\\ +2i\tilde{\beta}_{L-1}\left(\Gamma_{1}^{+}\Gamma_{4}^{-}-\Gamma_{4}^{+}\Gamma_{1}^{-}\right)+2i\tilde{\beta}_{L}\left(\Gamma_{3}^{+}\Gamma_{2}^{-}-\Gamma_{2}^{+}\Gamma_{3}^{-}\right).$ (129)
In matrix form (for $\ell=3$, thus $L=7$):
$H_{1}^{(2)}[\lambda]=2\,\vec{\Gamma}^{+}\begin{pmatrix}0&0&0&-i\beta_{6}&-i\beta_{1}&0&0\\ 0&0&i\beta_{7}&0&0&i\beta_{1}&0\\ 0&-i\beta_{7}&0&0&0&0&-i\beta_{2}\\ i\beta_{6}&0&0&0&0&0&i\beta_{2}\\ i\beta_{1}&0&0&0&0&-i\beta_{3}&0\\ 0&-i\beta_{1}&0&0&i\beta_{3}&0&0\\ 0&0&i\beta_{2}&-i\beta_{2}&0&0&0\\ \end{pmatrix}\vec{\Gamma}^{-}.$ (130)
C.4 Quantum Brachistochrone Counterdiabatic Driving (QBCD) in the minimal spin glass model
We report the details of how to compute $\mathcal{H}^{\mathrm{QBCD}}$, Eq. (35). The left- and right-localized edge states of the effective hopping model, $|\psi_{\mathrm{L}})$ and $|\psi_{\mathrm{R}})$ respectively, were derived in App. B for an infinitely long chain:
$|\psi_{\mathrm{L}})=\begin{pmatrix}a_{\mathrm{L}}&b_{\mathrm{L}}&e^{-\kappa_{\mathrm{L}}}a_{\mathrm{L}}&e^{-\kappa_{\mathrm{L}}}b_{\mathrm{L}}&e^{-2\kappa_{\mathrm{L}}}a_{\mathrm{L}}&\cdots\end{pmatrix}^{T},\qquad|\psi_{\mathrm{R}})=\begin{pmatrix}\cdots&e^{-2\kappa_{\mathrm{R}}}a_{\mathrm{R}}&e^{-\kappa_{\mathrm{R}}}b_{\mathrm{R}}&e^{-\kappa_{\mathrm{R}}}a_{\mathrm{R}}&b_{\mathrm{R}}&a_{\mathrm{R}}&c_{\mathrm{R}}\end{pmatrix}^{T}.$ (131)
Here, however, they are needed for a chain of finite length $L$. First of all, we set $a_{\mathrm{L}}=a_{\mathrm{R}}$, $b_{\mathrm{L}}=b_{\mathrm{R}}$ and $\kappa_{\mathrm{L}}=\kappa_{\mathrm{R}}$, which from App. B.4 is known to be valid at the avoided crossing. As explained in the main text, we want to obtain a CD term from the knowledge of the avoided crossing alone. Let us also parametrize the ratio of $a$ and $b$ by $e^{-\mu}\equiv a/b$. Plugging $|\psi_{\mathrm{L}})$ into the eigenvalue equation of $\mathcal{H}$, Eq. (89), one can find the relation between $\mu$, $\kappa$ and the eigenvalue $\epsilon$:
$e^{\kappa}=J^{\prime}\,e^{-\mu},\qquad\epsilon=J^{\prime}-Be^{\mu}.$ (132)
The right-end eigenvalue equation instead leads to
$B+\epsilon=\frac{J^{2}}{J^{\prime}}.$ (133)
From here, one can derive the results
$B=\frac{(1-J^{2})[J^{2}-(J^{\prime})^{2}]}{J^{\prime}[(J^{\prime})^{2}+1-2J^{2}]},\qquad\epsilon=\frac{(J^{\prime})^{2}-J^{4}}{J^{\prime}[(J^{\prime})^{2}+1-2J^{2}]},\qquad\mu=\ln\frac{J^{2}-\left(J^{\prime}\right)^{2}}{(1-J^{2})},\qquad k=\ln\frac{J^{\prime}(1-J^{2})}{[J^{2}-(J^{\prime})^{2}]},$ (134)
which are consistent with what found in App. B.4. Additionally, the last and first elements of the left and right localized states, respectively, are found to be
$c=\frac{JBe^{-\ell k}}{J^{\prime}}b\,.$ (135)
At this point, the two localized edge modes read
$|\psi_{\mathrm{L}})\simeq\frac{1}{\sqrt{Z_{\mathrm{L}}}}\begin{pmatrix}&e^{-\mu}\\ &1\\ &e^{-\kappa-\mu}\\ &e^{-\kappa}\\ &e^{-2\kappa-\mu}\\ &e^{-2\kappa}\\ &\vdots\\ &e^{-(\ell-1)\kappa-\mu}\\ &e^{-(\ell-1)\kappa}\\ &JBe^{-\ell\kappa}/J^{\prime}\end{pmatrix},\qquad|\psi_{\mathrm{R}})\simeq\frac{1}{\sqrt{Z_{\mathrm{R}}}}\begin{pmatrix}&e^{-(\ell-1)\kappa}\\ &e^{-(\ell-1)\kappa-\mu}\\ &\vdots\\ &e^{-2\kappa}\\ &e^{-2\kappa-\mu}\\ &e^{-\kappa}\\ &e^{-\kappa-\mu}\\ &1\\ &e^{-\mu}\\ &JBe^{-\mu}/J^{\prime}\end{pmatrix},$ (136)
where the normalization factors $Z_{\mathrm{L,R}}$ need not be specified for the moment.
The derivative with respect to $\lambda$ of the effective Hamiltonian is
$\partial_{\lambda}\mathcal{H}=\begin{pmatrix}J^{\prime}&1&&&&&\\ 1&0&1&&&&\\ &1&0&1&&&\\ &&1&0&1&&\\ &&&1&\ddots&\ddots&\\ &&&&\ddots&0&J\\ &&&&&J&1\end{pmatrix}.$ (137)
Correspondingly, the approximate energy gap takes the form
$\Delta_{\min}\approx(\psi_{\mathrm{R}}|\mathcal{H}|\psi_{\mathrm{L}})=J^{\prime}\lambda\,e^{-(\ell-1)\kappa-\mu}+\frac{J^{2}(\lambda-1)^{3}}{\lambda^{2}(J^{\prime})^{2}}\,e^{-\ell\kappa-\mu}+(\lambda-1)\ell e^{-(\ell-1)\kappa}(1+e^{-2\mu})\\ +(\ell-1)\lambda\left(e^{-\ell\kappa-2\mu}+e^{-(\ell-2)\kappa}\right)+J\lambda\left(\frac{J}{J^{\prime}}\frac{\lambda-1}{\lambda}\right)^{2}e^{-\ell\kappa-\mu}\,(1+e^{\kappa}),$ (138)
while the matrix element of the parametric derivative of the Hamiltonian is
$(\psi_{\mathrm{R}}|\partial_{\lambda}\mathcal{H}[\lambda_{c}]|\psi_{\mathrm{L}})=J^{\prime}\,e^{-(\ell-1)\kappa-\mu}+\frac{J^{2}(\lambda-1)^{2}}{\lambda^{2}(J^{\prime})^{2}}\,e^{-\ell\kappa-\mu}+\ell e^{-(\ell-1)\kappa}(1+e^{-2\mu})\\ +(\ell-1)\left(e^{-\ell\kappa-2\mu}+e^{-(\ell-2)\kappa}\right)+J\left(\frac{J}{J^{\prime}}\frac{\lambda-1}{\lambda}\right)^{2}e^{-\ell\kappa-\mu}\,(1+e^{\kappa}).$ (139)
With all this information, the QBCD term $\mathcal{H}^{\mathrm{QBCD}}$, Eq. (35), can be constructed. Notice that the normalization factors $Z_{\mathrm{L,R}}$ cancel out and need not be computed explicitly.
Appendix D Energy cost of the counterdiabatic Hamiltonians
In this section, we compute the trace norms of the CD Hamiltonians. Starting with the QBCD ansatz, Eq. (35), the trace norm of the square results in two projectors, leading to
$\begin{split}&\left\|\mathcal{H}^{\mathrm{QBCD}}\right\|^{2}=\frac{|(\psi_{\mathrm{R}}|\partial_{\lambda}\mathcal{H}|\psi_{\mathrm{L}})|^{2}}{\Delta_{\min}^{2}}\mathrm{Tr}\big{[}|\psi_{\mathrm{R}})(\psi_{\mathrm{L}}|+|\psi_{\mathrm{L}})(\psi_{\mathrm{R}}|\big{]}=O(1).\end{split}$ (140)
Indeed, the ratio of the matrix element and the gap is size-independent to leading order, as follows from Eqs. (138) and (139). Note that the non-locality and the exponentially small gap do not affect the overall energy cost.
Concerning the approximate first- and second-order variational CD terms, since the coefficients $\alpha_{i},\beta_{i}\sim\mathrm{O}(J)$ individually, one gets the typical scale for the norm
$\left\|H^{(1)}_{1}\right\|^{2}$ $=\sum_{n,m=1}^{L}\,\Big{|}H^{(1)}_{1}\Big{|}^{2}_{n,m}=4\sum_{n=1}^{\ell}|\alpha|^{2}_{n}+2|\alpha_{L}|^{2}\sim L,$ (141)
$\left\|H^{(2)}_{1}\right\|^{2}$ $=\sum_{n,m=1}^{L}\,\Big{|}H^{(2)}_{1}\Big{|}^{2}_{n,m}=4\sum_{n=1}^{\ell}|\beta|^{2}_{n}+2|\beta_{L}|^{2}\sim L,$ (142)
that is diverging in the thermodynamic limit.
Appendix E Details of numerical simulations
E.1 Particle-number-nonconserving free fermions
Suppose one wants to simulate numerically the quantum driving protocol by working at the level of Dirac fermions rather than Majoranas. Both $H[\lambda]$ and $H[\lambda]+H_{1}[\lambda]$ belong to the more general family of fermionic quadratic Hamiltonians
$H=\sum_{i,j=1}^{L}c_{i}\mathcal{H}_{ij}^{--}c_{j}+\sum_{i,j=1}^{L}c_{i}\mathcal{H}_{ij}^{-+}c_{j}^{\dagger}+\sum_{i,j=1}^{L}c_{i}^{\dagger}\mathcal{H}_{ij}^{+-}c_{j}+\sum_{i,j=1}^{L}c_{i}^{\dagger}\mathcal{H}_{ij}^{++}c_{j}^{\dagger}.$ (143)
For now, it is assumed that $H$ (and thus the matrices $\mathcal{H}$) are time-independent. The time-dependent case is recovered easily, as shown later on.
Introducing the vector
$\Psi=\begin{pmatrix}c_{1}&c_{2}&\dots&c_{L}&c^{\dagger}_{1}&c^{\dagger}_{N}&\dots&c^{\dagger}_{L}\end{pmatrix}^{T},$ (144)
one can rewrite succinctly
$H=\Psi^{\dagger}\mathcal{H}\Psi,$ (145)
where the matrix $\mathcal{H}$ reads
$\mathcal{H}=\left(\begin{array}[]{c|c}\mathcal{H}^{+-}&\mathcal{H}^{++}\\ \hline\cr\mathcal{H}^{--}&\mathcal{H}^{-+}\end{array}\right).$ (146)
By Hermiticity of $H$, one can choose the four submatrices above to respect
$\mathcal{H}^{+-}=-\left(\mathcal{H}^{-+}\right)^{*},\qquad\mathcal{H}^{++}=-\left(\mathcal{H}^{--}\right)^{*},$ (147)
and
$\mathcal{H}^{+-}=\left(\mathcal{H}^{+-}\right)^{\dagger},\qquad\mathcal{H}^{++}=-\left(\mathcal{H}^{++}\right)^{T}.$ (148)
To diagonalize $H$, it is sufficient to diagonalize $\mathcal{H}$ via the introduction of Bogoliubov operators $\xi$. Define $\mathcal{U}$ to be the matrix that diagonalizes $\mathcal{H}$, i.e.
$\mathcal{H}=\mathcal{U}\,\mathcal{H}_{\mathrm{diag}}\,\mathcal{U}^{\dagger}.$ (149)
Because of the symmetry in Eq. (147), it holds [125, 126, 127, 128]
$\mathcal{U}^{\dagger}:=\left(\begin{array}[]{c|c}u^{*}&v\\ \hline\cr v^{*}&u\end{array}\right),$ (150)
with $u,v$ being two $L\times L$ matrices, and
$\mathcal{H}_{\mathrm{diag}}=\left(\begin{array}[]{c|c}\epsilon&0\\ \hline\cr 0&-\epsilon\end{array}\right),$ (151)
where $\epsilon$ is the diagonal matrix of eigenvalues (that must appear in positive/negative couples, as entailed by the symmetries in Eqs. (147)–(148)). Finally, introducing
$\Phi:=\begin{pmatrix}\xi\\ \xi^{\dagger}\end{pmatrix}:=\mathcal{U}^{\dagger}\Psi\,,$ (152)
one gets to the diagonal form
$H=\Phi^{\dagger}\mathcal{H}_{\mathrm{diag}}\,\Phi=\sum_{j=1}^{L}\epsilon_{j}\left(\xi^{\dagger}_{j}\xi^{\phantom{\dagger}}_{j}-\xi^{\phantom{\dagger}}_{j}\xi^{\dagger}_{j}\right)=\sum_{j=1}^{L}\epsilon_{j}\left(2\xi^{\dagger}_{j}\xi^{\phantom{\dagger}}_{j}-1\right).$ (153)
The Heisenberg time evolution of the operators $c,c^{\dagger}$ follows:
$\Psi(t)=e^{iHt}\Psi e^{-iHt}=e^{iHt}\mathcal{U}\Phi e^{-iHt}=\mathcal{U}\left(\begin{array}[]{c|c}e^{-2i\epsilon t}&0\\ \hline\cr 0&e^{2i\epsilon t}\end{array}\right)\Phi=\mathcal{U}\left(\begin{array}[]{c|c}e^{-2i\epsilon t}&0\\ \hline\cr 0&e^{2i\epsilon t}\end{array}\right)\mathcal{U}^{\dagger}\Psi=e^{-2i\mathcal{H}t}\Psi.$ (154)
Thus, in order to time evolve the operators $c,c^{\dagger}$, one just needs to compute the exponential of the $2L\times 2L$ matrix $\mathcal{H}$.
[…]
where $\widetilde{\mathrm{T}}$ denotes the anti-time ordering. Notice that, upon trotterization in time steps of $dt$, it is the last time to act first on the operator:
$O(t)\simeq e^{iH(0)dt}\cdots e^{iH(t-dt)dt}O(0)e^{-iH(t-dt)dt}\cdots e^{-iH(0)dt}.$ (156)
For our $\Psi$ operator, the first Trotter step to apply is thus
$e^{iH(t-dt)dt}\Psi_{\alpha}e^{-iH(t-dt)dt}=\big{(}e^{-2i\mathcal{H}(t-dt)dt}\big{)}_{\alpha\beta}\Psi_{\beta},$ (157)
where the indices of the matrices are written down explicitly for convenience. The next time step reads
$e^{iH(t-2dt)dt}e^{iH(t-dt)dt}\Psi_{\alpha}e^{-iH(t-dt)dt}e^{iH(t-2dt)dt}$ $=\big{(}e^{-2i\mathcal{H}(t-dt)dt}\big{)}_{\alpha\beta}\,e^{iH(t-2dt)dt}\Psi_{\beta}e^{-iH(t-2dt)dt}$ (158)
$=\big{(}e^{-2i\mathcal{H}(t-dt)dt}\big{)}_{\alpha\beta}\big{(}e^{-2i\mathcal{H}(t-2dt)dt}\big{)}_{\beta\gamma}\Psi_{\gamma}.$ (159)
Iterating, one gets to
[…]
This form could have been guessed directly from Eq. (154), but all the steps were performed for clarity, given the subtlety of the anti-time ordering.
E.2 Specification of the problem under consideration
The formulas above apply to the Dirac Hamiltonian, Eq. (70), upon setting
$\mathcal{H}^{+-}[\lambda]=\begin{pmatrix}1-\lambda&-\lambda J_{1}/2&&&\lambda J_{L}/2\\ -\lambda J_{1}/2&1-\lambda&-\lambda J_{2}/2&&\\ &-\lambda J_{2}/2&\ddots&\ddots&\\ &&\ddots&1-\lambda&-\lambda J_{L-1}/2\\ \lambda J_{L}/2&&&-\lambda J_{L-1}/2&1-\lambda\\ \end{pmatrix},$ (161a)
$\mathcal{H}^{++}[\lambda]=\begin{pmatrix}0&-\lambda J_{1}/2&&&-\lambda J_{L}/2\\ \lambda J_{1}/2&0&-\lambda J_{2}/2&&\\ &\lambda J_{2}/2&\ddots&\ddots&\\ &&\ddots&0&-\lambda J_{L-1}/2\\ \lambda J_{L}/2&&&\lambda J_{L-1}/2&0\\ \end{pmatrix}.$ (161b)
One can check that the conjugation properties, Eq. (148), are satisfied. The first order CD Hamiltonian, Eq. (116), reads instead
$\mathcal{H}^{+-}[\lambda]=0,\qquad\mathcal{H}^{++}[\lambda]=\begin{pmatrix}0&-i\alpha_{1}&&&-i\alpha_{L}\\ i\alpha_{1}&0&-i\alpha_{2}&&\\ &i\alpha_{1}&\ddots&\ddots&\\ &&\ddots&0&-i\alpha_{L-1}\\ i\alpha_{L}&&&i\alpha_{L-1}&0\\ \end{pmatrix},$ (162)
while the second-order one reads
$\mathcal{H}^{+-}[\lambda]=0,\qquad\mathcal{H}^{++}[\lambda]=\begin{pmatrix}0&0&-i\beta_{1}&&&-i\beta_{L-1}&\\ 0&0&0&-i\beta_{2}&&&-i\beta_{L}\\ i\beta_{1}&0&0&0&\ddots&&\\ &i\beta_{2}&0&\ddots&\ddots&\ddots&\\ &&\ddots&\ddots&\ddots&\ddots&-i\beta_{L-2}\\ i\beta_{L-1}&&&\ddots&\ddots&0&0\\ &i\beta_{L}&&&i\beta_{L-2}&0&0\\ \end{pmatrix}.$ (163)
We are using the same letter $\mathcal{H}$ for all the different Hamiltonians, bare or counterdiabatic, in order not to overload the notation.
The QBCD Hamiltonian, Eq. (35), instead, reads in the $\Gamma$ fermion representation
$H^{\mathrm{QBCD}}_{1}=\sum_{i,j=1}^{L}\Gamma^{+}_{i}\mathcal{H}^{\mathrm{QBCD}}_{ij}\Gamma^{-}_{j}.$ (164)
In order to find the $c$-fermion representation of $H^{\mathrm{QBCD}}_{1}$, it is convenient to transform back the $\Gamma^{\pm}$ operators to the Dirac representation:
$\begin{cases}\displaystyle\Gamma^{+}_{k}=\frac{i\,\gamma_{k}+\gamma_{2L-k+1}}{\sqrt{2}}\,,&k=1,3,\dots,L\\ \displaystyle\Gamma^{+}_{k}=\frac{\gamma_{k}+i\,\gamma_{2L-k+1}}{\sqrt{2}}\,,&k=2,4,\dots,L-1\end{cases}\quad\implies\quad\begin{cases}\displaystyle\Gamma^{+}_{2j-1}=i\frac{c_{j}+c^{\dagger}_{j}+c_{L-j+1}-c^{\dagger}_{L-j+1}}{2},&j=1,\dots\ell+1\\ \displaystyle\Gamma^{+}_{2j}=i\,\frac{c_{j}-c^{\dagger}_{j}+c_{L-j+1}+c^{\dagger}_{L-j+1}}{2},&j=1,\dots\ell\end{cases}$ (165)
From here, the corresponding Dirac Hamiltonian reads
$H^{\mathrm{QBCD}}_{1}=\frac{1}{4}\sum_{i,j}$ $\left[\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j-1}\,\left(c_{i}+c^{\dagger}_{i}+c_{L-i+1}-c^{\dagger}_{L-i+1}\right)\left(c^{\dagger}_{j}+c_{j}+c^{\dagger}_{L-j+1}-c_{L-j+1}\right)\right.$ (166)
$+\mathcal{H}^{\mathrm{QBCD}}_{2i,2j}\,\left(c_{i}-c^{\dagger}_{i}+c_{L-i+1}+c^{\dagger}_{L-i+1}\right)\left(c^{\dagger}_{j}-c_{j}+c^{\dagger}_{L-j+1}+c_{L-j+1}\right)$
$+\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j}\left(c_{i}+c^{\dagger}_{i}+c_{L-i+1}-c^{\dagger}_{L-i+1}\right)\left(c^{\dagger}_{j}-c_{j}+c^{\dagger}_{L-j+1}+c_{L-j+1}\right)$
$\left.+\mathcal{H}^{\mathrm{QBCD}}_{2i,2j-1}\,\left(c_{i}-c^{\dagger}_{i}+c_{L-i+1}+c^{\dagger}_{L-i+1}\right)\left(c^{\dagger}_{j}+c_{j}+c^{\dagger}_{L-j+1}-c_{L-j+1}\right)\right],$
where the summations are running until $\ell+1$ for odd indices and until $\ell$ for the even ones. Above, $\mathcal{H}_{ij}$ are the matrix elements of the single-particle QBCD Hamiltonian. Decomposing further the expression above in the particle number-conserving and non-conserving blocks, one finds four different matrices $\mathcal{H}^{++}$, $\mathcal{H}^{+-}$, $\mathcal{H}^{-+}$, and $\mathcal{H}^{--}$. For the sake of brevity, we use a non-symmetrized form of these submatrices, which does not obey Eqs. (147)–(148):
$\mathcal{H}^{++}_{ij}$ $=\frac{1}{4}\left(\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j-1}-\mathcal{H}^{\mathrm{QBCD}}_{2i,2j}+\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j}-\mathcal{H}^{\mathrm{QBCD}}_{2i,2j-1}\right),$ (167a)
$\mathcal{H}^{++}_{L-i+1,L-j+1}$ $=\frac{1}{4}\left(-\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j-1}+\mathcal{H}^{\mathrm{QBCD}}_{2i,2j}-\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j}+\mathcal{H}^{\mathrm{QBCD}}_{2i,2j-1}\right),$ (167b)
$\mathcal{H}^{++}_{L-i+1,j}$ $=\frac{1}{4}\left(-\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j-1}+\mathcal{H}^{\mathrm{QBCD}}_{2i,2j}-\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j}+\mathcal{H}^{\mathrm{QBCD}}_{2i,2j-1}\right),$ (167c)
$\mathcal{H}^{++}_{i,L-j+1}$ $=\frac{1}{4}\left(\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j-1}-\mathcal{H}^{\mathrm{QBCD}}_{2i,2j}+\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j}-\mathcal{H}^{soor\mathrm{QBCD}}_{2i,2j-1}\right),$ (167d)
for $\mathcal{H}^{++}$,
$\mathcal{H}^{+-}_{ij}$ $=\frac{1}{4}\left(\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j-1}+\mathcal{H}^{\mathrm{QBCD}}_{2i,2j}-\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j}-\mathcal{H}^{\mathrm{QBCD}}_{2i,2j-1}\right),$ (168a)
$\mathcal{H}^{+-}_{L-i+1,L-j+1}$ $=\frac{1}{4}\left(\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j-1}+\mathcal{H}^{\mathrm{QBCD}}_{2i,2j}-\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j}-\mathcal{H}^{\mathrm{QBCD}}_{2i,2j-1}\right),$ (168b)
$\mathcal{H}^{+-}_{L-i+1,j}$ $=\frac{1}{4}\left(-\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j-1}-\mathcal{H}^{\mathrm{QBCD}}_{2i,2j}+\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j}+\mathcal{H}^{\mathrm{QBCD}}_{2i,2j-1}\right),$ (168c)
$\mathcal{H}^{+-}_{i,L-j+1}$ $=\frac{1}{4}\left(-\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j-1}-\mathcal{H}^{\mathrm{QBCD}}_{2i,2j}+\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j}+\mathcal{H}^{\mathrm{QBCD}}_{2i,2j-1}\right),$ (168d)
for $\mathcal{H}^{+-}$,
$\mathcal{H}^{-+}_{ij}$ $=\frac{1}{4}\left(\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j-1}+\mathcal{H}^{\mathrm{QBCD}}_{2i,2j}+\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j}+\mathcal{H}^{\mathrm{QBCD}}_{2i,2j-1}\right),$ (169a)
$\mathcal{H}^{-+}_{L-i+1,L-j+1}$ $=\frac{1}{4}\left(\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j-1}+\mathcal{H}^{\mathrm{QBCD}}_{2i,2j}+\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j}+\mathcal{H}^{\mathrm{QBCD}}_{2i,2j-1}\right),$ (169b)
$\mathcal{H}^{-+}_{L-i+1,j}$ $=\frac{1}{4}\left(\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j-1}+\mathcal{H}^{\mathrm{QBCD}}_{2i,2j}+\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j}+\mathcal{H}^{\mathrm{QBCD}}_{2i,2j-1}\right),$ (169c)
$\mathcal{H}^{-+}_{i,L-j+1}$ $=\frac{1}{4}\left(\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j-1}+\mathcal{H}^{\mathrm{QBCD}}_{2i,2j}+\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j}+\mathcal{H}^{\mathrm{QBCD}}_{2i,2j-1}\right),$ (169d)
for $\mathcal{H}^{-+}$, and
$\mathcal{H}^{--}_{ij}$ $=\frac{1}{4}\left(\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j-1}-\mathcal{H}^{\mathrm{QBCD}}_{2i,2j}-\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j}+\mathcal{H}^{\mathrm{QBCD}}_{2i,2j-1}\right),$ (170a)
$\mathcal{H}^{--}_{L-i+1,L-j+1}$ $=\frac{1}{4}\left(-\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j-1}+\mathcal{H}^{\mathrm{QBCD}}_{2i,2j}+\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j}-\mathcal{H}^{\mathrm{QBCD}}_{2i,2j-1}\right),$ (170b)
$\mathcal{H}^{--}_{L-i+1,j}$ $=\frac{1}{4}\left(\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j-1}-\mathcal{H}^{\mathrm{QBCD}}_{2i,2j}-\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j}+\mathcal{H}^{\mathrm{QBCD}}_{2i,2j-1}\right),$ (170c)
$\mathcal{H}^{--}_{i,L-j+1}$ $=\frac{1}{4}\left(-\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j-1}+\mathcal{H}^{\mathrm{QBCD}}_{2i,2j}+\mathcal{H}^{\mathrm{QBCD}}_{2i-1,2j}-\mathcal{H}^{\mathrm{QBCD}}_{2i,2j-1}\right),$ (170d)
for $\mathcal{H}^{--}$. Notice that the block Hamiltonians above separate into four further blocks, depending on whether the Dirac fermions are to the left or right of site $\ell$.
E.3 Expectation values
Once the operators $c(t),c^{\dagger}(t)$ are obtained, expectation values can be computed (note instead that accessing the fidelity is more complicated). For instance, the instantaneous energy expectation reads
$GS\displaystyle\expectationvalue{H[\lambda](t)}{\mathrm{GS}}$ $=-\lambda\sum_{j=1}^{L}J_{j}\expectationvalue{\sigma_{j}^{z}(t)\sigma_{j+1}^{z}(t)}{\rightarrow}-(1-\lambda)\sum_{j=1}^{L}\expectationvalue{\sigma_{j}^{x}(t)}{\rightarrow}$ (171)
=$0𝑐\displaystyle=-\lambda\sum_{j=1}^{L}\tilde{J}_{j}\expectationvalue{\left(c_{j}^{\dagger}(t)c_{j+1}^{\dagger}(t)+c_{j}^{\dagger}(t)c_{j+1}(t)-c_{j}(t)c_{j+1}^{\dagger}(t)-c_{j}(t)c_{j+1}(t)\right)}{0_{c}}$
[…]
where $\Psi_{i}$ refers to $c_{i}$, and $\Psi_{i^{\dagger}}$ to $c_{i}^{\dagger}$ (and similarly for the corresponding rows and columns of $\mathcal{V}$). A similar computation leads to
$0𝑐\displaystyle\sum_{j=1}^{L}\tilde{J}_{j}\expectationvalue{c_{j}^{\dagger}(t)c_{j+1}(t)}{0_{c}}$ $=\sum_{i,j=1}^{L}\tilde{J}_{j}\mathcal{V}_{j^{\dagger},i}\mathcal{V}_{j+1,i^{\dagger}},$ (174)
$0𝑐\displaystyle\sum_{j=1}^{L}\tilde{J}_{j}\expectationvalue{c_{j}(t)c_{j+1}^{\dagger}(t)}{0_{c}}$ $=\sum_{i,j=1}^{L}\tilde{J}_{j}\mathcal{V}_{j,i}\mathcal{V}_{(j+1)^{\dagger},i^{\dagger}},$ (175)
$0𝑐\displaystyle\sum_{j=1}^{L}\tilde{J}_{j}\expectationvalue{c_{j}(t)c_{j+1}(t)}{0_{c}}$ $=\sum_{i,j=1}^{L}\tilde{J}_{j}\mathcal{V}_{j,i}\mathcal{V}_{j+1,i^{\dagger}},$ (176)
$0𝑐\displaystyle\sum_{j=1}^{L}\expectationvalue{c_{j}^{\dagger}(t)c_{j}(t)}{0_{c}}$ $=\sum_{i,j=1}^{L}\mathcal{V}_{j^{\dagger},i}\mathcal{V}_{j,i^{\dagger}}.$ (177)
Putting everything together, one finds
[…]
Similarly, the kink number reads
[…]
where
$\tilde{1}_{j}=\begin{cases}1&j=1,2,\dots,L-1\\ -1&j=L\end{cases}$ (180)
takes into account the restriction to the even parity sector.
Appendix F Derivation of the minimization equations for the Max Cut
In this section, we derive the equations for the coefficients for the first- and second-order CD expansions for the Max Cut problem. The enlisted Hamiltonians at a given $\lambda$ value read as
$H_{P}=-\lambda\sum_{c}\frac{1-\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}}{2}-(1-\lambda)\sum_{j}\sigma^{x}_{j},$ (181)
$H^{(1)}_{1}=\sum_{c}\alpha_{c}\left(\sigma^{z}_{c_{1}}\sigma^{y}_{c_{2}}+\sigma^{y}_{c_{1}}\sigma^{z}_{c_{2}}\right),$ (182)
$H^{(2)}_{1}=\sum_{c,c^{\prime}}\beta_{c,c^{\prime}}\delta_{c_{2},c^{\prime}_{1}}\left(\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{y}_{c^{\prime}_{2}}+\sigma^{y}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}\right).$ (183)
In the first-order local expansion, the action integrand reads
$G(H^{(1)}_{1})=$ $-\sum_{c=1}^{M}\frac{1+\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}}{2}+\sum_{c=1}^{M}\frac{1+2\lambda\alpha_{c}}{2}[\sigma^{x}_{c_{1}}+\sigma^{x}_{c_{2}}]$
$+\lambda\sum_{c,c^{\prime}}\alpha_{c}\left[(1-\delta_{c_{1},c^{\prime}_{1}})\delta_{c_{2},c^{\prime}_{2}}\sigma^{z}_{c_{1}}\sigma^{z}_{c^{\prime}_{1}}\sigma^{x}_{c_{2}}+\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}\right]$
$+4(1-\lambda)\sum_{c=1}^{M}\alpha_{c}(\sigma^{y}_{c_{1}}\sigma^{y}_{c_{2}}-\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}})+\lambda\sum_{c,c^{\prime}}^{M}\alpha_{c}(\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}\delta_{c_{2},c^{\prime}_{1}}+\sigma^{z}_{c^{\prime}_{1}}\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}}\delta_{c_{1},c^{\prime}_{2}})$
$=-\sum_{c=1}^{M}\frac{1+8(1-\lambda)\alpha_{c}}{2}\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}+\sum_{c=1}^{M}\frac{1+2\lambda\alpha_{c}}{2}(\sigma^{x}_{c_{1}}+\sigma^{x}_{c_{2}})+4(1-\lambda)\sum_{c=1}^{M}\alpha_{c}\sigma^{y}_{c_{1}}\sigma^{y}_{c_{2}}$
$+\lambda\sum_{c,c^{\prime}}\alpha_{c}\left[(1-\delta_{c_{1},c^{\prime}_{1}})\delta_{c_{2},c^{\prime}_{2}}\sigma^{z}_{c_{1}}\sigma^{z}_{c^{\prime}_{1}}\sigma^{x}_{c_{2}}+\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}+\delta_{c_{2},c^{\prime}_{1}}\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}+\delta_{c_{1},c^{\prime}_{2}}\sigma^{z}_{c^{\prime}_{1}}\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}}\right],$
where all terms have been ordered so that the strictly greater inequality is always understood between the subindices, $c_{1}<c_{2}$ and $c^{\prime}_{1}<c^{\prime}_{2}$. The corresponding trace norm reads
$\mathrm{Tr}\left[G^{\dagger}\,G\right]=\sum_{c}\frac{[1+8(1-\lambda)\alpha_{c}]^{2}}{4}$
$+\frac{1}{2}\sum_{c,c^{\prime}}(2\lambda^{2}\alpha_{c}\alpha_{c^{\prime}}+\lambda(\alpha_{c}+\alpha_{c^{\prime}}))\left[\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})+(1-\delta_{c_{1},c^{\prime}_{1}})\delta_{c_{2},c^{\prime}_{2}}+2\delta_{c_{1},c^{\prime}_{1}}\delta_{c_{2},c^{\prime}_{2}}+\delta_{c_{1},c^{\prime}_{2}}+\delta_{c_{2},c^{\prime}_{1}}\right]$
$+\lambda^{2}\sum_{c}(\alpha_{c}+\alpha_{c^{\prime}})^{2}\delta_{c_{1},c^{\prime}_{2}}+\lambda^{2}\sum_{c}\alpha^{2}_{c}\left[\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})+(1-\delta_{c_{1},c^{\prime}_{1}})\delta_{c_{2},c^{\prime}_{2}}\right]+16(1-\lambda)^{2}\sum_{c}\alpha^{2}_{c}.$ (185)
Taking now the derivative with respect to a given $\alpha_{c}$ the coefficient of $\lambda^{2}\alpha_{c}$ will equal the number of ways the given spins of $\alpha_{c}$ appear in other interactions, that is $6$ altogether for the two spins in $\alpha_{c}$. At the same time, all $\alpha_{c^{\prime}}$ will be coupled to all other interactions that share at least one spin with it, plus the additional correlation term of $\delta_{c_{1},c^{\prime}_{2}}+\delta_{c_{2},c^{\prime}_{1}}$. This yields
$\frac{\delta\mathrm{Tr}(G^{\dagger}G)}{\delta\alpha_{c}}$ $=\left[64(1-\lambda)^{2}+12\lambda^{2}\right]\alpha_{c}+2\lambda^{2}\sum_{c^{\prime}}\alpha_{c^{\prime}}\left[\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})+(1-\delta_{c_{1},c^{\prime}_{1}})\delta_{c_{2},c^{\prime}_{2}}+2\delta_{c_{1},c^{\prime}_{2}}+2\delta_{c_{2},c^{\prime}_{1}}\right]$ (186)
$=-\lambda\sum_{c^{\prime}}\left(\delta_{c_{1},c^{\prime}_{1}}+\delta_{c_{2},c^{\prime}_{2}}+\delta_{c_{1},c^{\prime}_{2}}+\delta_{c_{2},c^{\prime}_{1}}\right)-4(1-\lambda)=-2(\lambda-2).$
For the second order, the action integrated reads
$G(H^{(2)}_{1})=-\sum_{c=1}^{M}\frac{1+\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}}{2}+\sum_{j=1}^{L}\sigma^{x}_{j}+4(1-\lambda)\sum_{c,c^{\prime}}^{M}\delta_{c_{2},c^{\prime}_{1}}\beta_{c,c^{\prime}}\left(\sigma^{y}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{y}_{c^{\prime}_{2}}-\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}\right)$
$+\lambda\sum_{c,c^{\prime},c^{\prime\prime}}(\delta_{c_{1},c^{\prime}_{2}}\beta_{c^{\prime},c}\delta_{c^{\prime}_{1},c^{\prime\prime}_{2}}\sigma^{z}_{c^{\prime\prime}_{1}}\sigma^{x}_{c^{\prime\prime}_{2}}\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}}+\delta_{c_{2},c^{\prime}_{1}}\beta_{c,c^{\prime}}\delta_{c^{\prime\prime}_{1},c^{\prime}_{2}}\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{x}_{c^{\prime\prime}_{1}}\sigma^{z}_{c^{\prime\prime}_{2}})-\lambda\sum_{c,c^{\prime}}(\delta_{c_{2},c^{\prime}_{1}}\beta_{c,c^{\prime}}+\delta_{c^{\prime}_{2},c_{1}}\beta_{c^{\prime},c})\sigma^{y}_{c_{1}}\sigma^{y}_{c_{2}}$
$-\lambda\sum_{c,c^{\prime},c^{\prime\prime}}(\beta_{c,c^{\prime}}\delta_{c_{2},c^{\prime}_{1}}+\beta_{c^{\prime\prime},c^{\prime}}\delta_{c^{\prime\prime}_{2},c^{\prime}_{1}})(1-\delta_{c_{1},c^{\prime\prime}_{1}})\delta_{c_{2},c^{\prime\prime}_{2}}\sigma^{z}_{c^{\prime\prime}_{1}}\sigma^{z}_{c_{1}}\sigma^{y}_{c_{2}}\sigma^{y}_{c^{\prime}_{2}}+(\beta_{c,c^{\prime}}\delta_{c_{2},c^{\prime}_{1}}\delta_{c_{1},c^{\prime}_{1}}+\beta_{c,c^{\prime\prime}}\delta_{c_{2},c^{\prime\prime}_{1}}\delta_{c_{1},c^{\prime\prime}_{1}})(1-\delta_{c^{\prime}_{2},c^{\prime\prime}_{2}})\sigma^{y}_{c_{1}}\sigma^{y}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}\sigma^{z}_{c^{\prime\prime}_{2}}$
$+\lambda\sum_{c,c^{\prime},c^{\prime\prime}}\beta_{c,c^{\prime}}\delta_{c_{2},c^{\prime}_{1}}\left[\sigma^{z}_{c_{1}}(1-\delta_{c^{\prime\prime}_{1},c_{2}})\delta_{c^{\prime\prime}_{2},c^{\prime}_{2}}(\sigma^{z}_{c^{\prime\prime}_{1}}\sigma^{x}_{c_{2}}+\sigma^{x}_{c_{2}}\sigma^{z}_{c^{\prime\prime}_{1}})\sigma^{x}_{c^{\prime}_{2}}-\sigma^{z}_{c_{1}}\delta_{c^{\prime\prime}_{1},c_{2}}(1-\delta_{c^{\prime\prime}_{2},c^{\prime}_{2}})\sigma^{y}_{c_{2}}(\sigma^{z}_{c^{\prime\prime}_{2}}\sigma^{y}_{c^{\prime}_{2}}+\sigma^{y}_{c^{\prime}_{2}}\sigma^{z}_{c^{\prime\prime}_{2}})\right]$
$-\lambda\sum_{c,c^{\prime},c^{\prime\prime}}\beta_{c,c^{\prime}}\delta_{c_{2},c^{\prime}_{1}}\left[(1-\delta_{c^{\prime\prime}_{1},c_{1}})\delta_{c^{\prime\prime}_{2},c_{2}}(\sigma^{z}_{c^{\prime\prime}_{1}}\sigma^{y}_{c_{1}}+\sigma^{y}_{c_{1}}\sigma^{z}_{c^{\prime\prime}_{1}})\sigma^{y}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}-\delta_{c^{\prime\prime}_{1},c_{1}}(1-\delta_{c^{\prime\prime}_{2},c_{2}})\sigma^{x}_{c_{1}}(\sigma^{z}_{c^{\prime\prime}_{2}}\sigma^{x}_{c_{2}}+\sigma^{x}_{c_{2}}\sigma^{z}_{c^{\prime\prime}_{2}})\sigma^{z}_{c^{\prime}_{2}}\right],$
where the ordering in the indices, $c_{1}<c_{2}<c^{\prime}_{2}$ of the connected interaction edges is fixed, and those denoted by $c^{\prime\prime}$ refer to the ones of the second order CD term. Here, the first term originates simply from the fact that in every three-body terms, the corresponding $\sigma^{x}_{j}$ will lead to either $\sigma^{z}_{j}$ or $\sigma^{y}_{j}$. The factor of three in the second line comes from the freedom of choosing one of the $c^{\prime\prime}_{i=1,2}$ spins in the construction. The third and fourth lines correspond to the cases when the $\sigma^{z}\sigma^{z}$ terms can potentially have non-zero commutator with the $\sigma^{y}\sigma^{x}$ and $\sigma^{x}\sigma^{y}$ terms.
The total second order term thus reads as
$G(H^{(2)}_{1}+H^{(1)}_{1})=-\sum_{c=1}^{M}\frac{1+8(1-\lambda)\alpha_{c}}{2}\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}+\sum_{c=1}^{M}\frac{1+2\lambda\alpha_{c}}{2}(\sigma^{x}_{c_{1}}+\sigma^{x}_{c_{2}})$ (188)
$+\lambda\sum_{c,c^{\prime}}\alpha_{c}\left[(1-\delta_{c_{1},c^{\prime}_{1}})\delta_{c_{2},c^{\prime}_{2}}\sigma^{z}_{c_{1}}\sigma^{z}_{c^{\prime}_{1}}\sigma^{x}_{c_{2}}+\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}\right]$
$+4(1-\lambda)\sum_{c,c^{\prime}}^{M}\delta_{c_{2},c^{\prime}_{1}}\beta_{c,c^{\prime}}\sigma^{y}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{y}_{c^{\prime}_{2}}-\sum_{c,c^{\prime}}\delta_{c_{2},c^{\prime}_{1}}\left[4(1-\lambda)\beta_{c,c^{\prime}}-\lambda(\alpha_{c}+\alpha_{c^{\prime}})\right]\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}$
$+\lambda\sum_{c,c^{\prime},c^{\prime\prime}}(\delta_{c_{1},c^{\prime}_{2}}\beta_{c^{\prime},c}\delta_{c^{\prime}_{1},c^{\prime\prime}_{2}}\sigma^{z}_{c^{\prime\prime}_{1}}\sigma^{x}_{c^{\prime\prime}_{2}}\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}}+\delta_{c_{2},c^{\prime}_{1}}\beta_{c,c^{\prime}}\delta_{c^{\prime\prime}_{1},c^{\prime}_{2}}\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{x}_{c^{\prime\prime}_{1}}\sigma^{z}_{c^{\prime\prime}_{2}})-\sum_{c}\left(\lambda\sum_{c^{\prime}}\delta_{c_{2},c^{\prime}_{1}}\beta_{c,c^{\prime}}+\delta_{c^{\prime}_{2},c_{1}}\beta_{c^{\prime},c}-4(1-\lambda)\alpha_{c}\right)\sigma^{y}_{c_{1}}\sigma^{y}_{c_{2}}$
$-\lambda\sum_{c,c^{\prime},c^{\prime\prime}}(\beta_{c,c^{\prime}}\delta_{c_{2},c^{\prime}_{1}}+\beta_{c^{\prime\prime},c^{\prime}}\delta_{c^{\prime\prime}_{2},c^{\prime}_{1}})(1-\delta_{c_{1},c^{\prime\prime}_{1}})\delta_{c_{2},c^{\prime\prime}_{2}}\sigma^{z}_{c^{\prime\prime}_{1}}\sigma^{z}_{c_{1}}\sigma^{y}_{c_{2}}\sigma^{y}_{c^{\prime}_{2}}+(\beta_{c,c^{\prime}}\delta_{c_{2},c^{\prime}_{1}}\delta_{c_{2},c^{\prime\prime}_{1}}+\beta_{c,c^{\prime\prime}}\delta_{c_{2},c^{\prime\prime}_{1}}\delta_{c_{2},c^{\prime}_{1}})(1-\delta_{c^{\prime}_{2},c^{\prime\prime}_{2}})\sigma^{y}_{c_{1}}\sigma^{y}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}\sigma^{z}_{c^{\prime\prime}_{2}}$
$+\lambda\sum_{c,c^{\prime},c^{\prime\prime}}\beta_{c,c^{\prime}}\delta_{c_{2},c^{\prime}_{1}}\left[\sigma^{z}_{c_{1}}(1-\delta_{c^{\prime\prime}_{1},c_{2}})\delta_{c^{\prime\prime}_{2},c^{\prime}_{2}}(\sigma^{z}_{c^{\prime\prime}_{1}}\sigma^{x}_{c_{2}}+\sigma^{x}_{c_{2}}\sigma^{z}_{c^{\prime\prime}_{1}})\sigma^{x}_{c^{\prime}_{2}}-\sigma^{z}_{c_{1}}\delta_{c^{\prime\prime}_{1},c_{2}}(1-\delta_{c^{\prime\prime}_{2},c^{\prime}_{2}})\sigma^{y}_{c_{2}}(\sigma^{z}_{c^{\prime\prime}_{2}}\sigma^{y}_{c^{\prime}_{2}}+\sigma^{y}_{c^{\prime}_{2}}\sigma^{z}_{c^{\prime\prime}_{2}})\right]$
$-\lambda\sum_{c,c^{\prime},c^{\prime\prime}}\beta_{c,c^{\prime}}\delta_{c_{2},c^{\prime}_{1}}\left[(1-\delta_{c^{\prime\prime}_{1},c_{1}})\delta_{c^{\prime\prime}_{2},c_{2}}(\sigma^{z}_{c^{\prime\prime}_{1}}\sigma^{y}_{c_{1}}+\sigma^{y}_{c_{1}}\sigma^{z}_{c^{\prime\prime}_{1}})\sigma^{y}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}-\delta_{c^{\prime\prime}_{1},c_{1}}(1-\delta_{c^{\prime\prime}_{2},c_{2}})\sigma^{x}_{c_{1}}(\sigma^{z}_{c^{\prime\prime}_{2}}\sigma^{x}_{c_{2}}+\sigma^{x}_{c_{2}}\sigma^{z}_{c^{\prime\prime}_{2}})\sigma^{z}_{c^{\prime}_{2}}\right].$
Now, the trace of the square reads
$\mathrm{Tr}\left[G^{\dagger}G\right]=\sum_{c}\frac{[1+8(1-\lambda)\alpha_{c}]^{2}}{4}$ (189)
$+\frac{1}{2}\sum_{c,c^{\prime}}(2\lambda^{2}\alpha_{c}\alpha_{c^{\prime}}+\lambda(\alpha_{c}+\alpha_{c^{\prime}}))\left[\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})+(1-\delta_{c_{1},c^{\prime}_{1}})\delta_{c_{2},c^{\prime}_{2}}+2\delta_{c_{1},c^{\prime}_{1}}\delta_{c_{2},c^{\prime}_{2}}+\delta_{c_{1},c^{\prime}_{2}}+\delta_{c_{2},c^{\prime}_{1}}\right]$
$+\lambda^{2}\sum_{c,c^{\prime}}\alpha^{2}_{c}\left[\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})+(1-\delta_{c_{1},c^{\prime}_{1}})\delta_{c_{2},c^{\prime}_{2}}\right]$
+16($+16(1-\lambda)^{2}\sum_{c,c^{\prime}}\delta_{c_{2},c^{\prime}_{1}}\beta^{2}_{c,c^{\prime}}+\sum_{c,c^{\prime}}\delta_{c_{2},c^{\prime}_{1}}\left[4(1-\lambda)\beta_{c,c^{\prime}}-\lambda(\alpha_{c}+\alpha_{c^{\prime}})\right]^{2}+\lambda^{2}\sum_{c,c^{\prime},c^{\prime\prime}}\left(\delta_{c^{\prime}_{2},c_{1}}\beta^{2}_{c^{\prime},c}\delta_{c^{\prime}_{1},c^{\prime\prime}_{2}}+\delta_{c_{2},c^{\prime}_{1}}\beta^{2}_{c,c^{\prime}}\delta_{c^{\prime\prime}_{1},c^{\prime}_{2}}\right)$
$+\sum_{c}\left(\sum_{c^{\prime}}\lambda\delta_{c_{2},c^{\prime}_{1}}\beta_{c,c^{\prime}}+\lambda\delta_{c^{\prime}_{2},c_{1}}\beta_{c^{\prime},c}-4(1-\lambda)\alpha_{c}\right)^{2}$
$+\lambda^{2}\sum_{c,c^{\prime},c^{\prime\prime}}(\beta_{c,c^{\prime}}\delta_{c_{2},c^{\prime}_{1}}+\beta_{c^{\prime\prime},c^{\prime}}\delta_{c^{\prime\prime}_{2},c^{\prime}_{1}})^{2}(1-\delta_{c^{\prime\prime}_{1},c_{1}})\delta_{c_{2},c^{\prime\prime}_{2}}+(\beta_{c,c^{\prime}}\delta_{c_{2},c^{\prime}_{1}}\delta_{c_{2},c^{\prime\prime}_{1}}+\beta_{c,c^{\prime\prime}}\delta_{c_{2},c^{\prime\prime}_{1}}\delta_{c_{2},c^{\prime}_{1}})^{2}(1-\delta_{c^{\prime}_{2},c^{\prime\prime}_{2}})$
$+2\lambda^{2}\sum_{c,c^{\prime},c^{\prime\prime}}\beta^{2}_{c,c^{\prime}}\delta_{c_{2},c^{\prime}_{1}}\left[(1-\delta_{c^{\prime\prime}_{1},c_{2}})\delta_{c^{\prime\prime}_{2},c^{\prime}_{2}}+\delta_{c^{\prime\prime}_{1},c_{2}}(1-\delta_{c^{\prime\prime}_{2},c^{\prime}_{2}})+(1-\delta_{c^{\prime\prime}_{1},c_{1}})\delta_{c^{\prime\prime}_{2},c_{2}}+\delta_{c^{\prime\prime}_{1},c_{1}}(1-\delta_{c^{\prime\prime}_{2},c_{2}})\right].$
For computing the derivatives we again use combinatorial arguments. As all terms in $G\left(H^{(1)}_{1}+H^{(2)}_{1}\right)$ originating from different ordering of the connected spins are orthogonal to each other the constant prefactor in front of $\lambda^{2}\beta_{c,c^{\prime}}$ is given by four times the number of ways how a given spin can be involved in other interactions, i.e., by $12$. This can be seen easily by the fact, that one gets a non-zero operator for the commutator whenever the first two operators of $\sigma^{z}_{c^{\prime}_{1}}\sigma^{x}_{c^{\prime}_{2}}\sigma^{y}_{c^{\prime}_{3}}$ match any of the three in, $\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{3}$. As both of the resulting commutators give orthogonal operators, eventually one obtains two times the number times the given $c_{1},\,c_{2},\,c_{3}$ spins appear in the interactions. The off-diagonal, terms connecting different $\beta_{c,c^{\prime\prime}}$-s are obtained similarly as in the first order case, by enlisting all possible connections to other interaction pairs of $c,c^{\prime\prime}$ or $c^{\prime}c^{\prime\prime}$. One finds
$\frac{\delta\mathrm{Tr}(G^{\dagger}G)}{\delta\alpha_{c}}=0$
$\Rightarrow\left[64(1-\lambda)^{2}+12\lambda^{2}\right]\alpha_{c}+2\lambda^{2}\sum_{c^{\prime}}\alpha_{c^{\prime}}\left[\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})+(1-\delta_{c_{1},c^{\prime}_{1}})\delta_{c_{2},c^{\prime}_{2}}+2\delta_{c_{1},c^{\prime}_{2}}+2\delta_{c_{2},c^{\prime}_{1}}\right]$ (190)
$=16\lambda(1-\lambda)\sum_{c^{\prime}}(\beta_{c,c^{\prime}}\delta_{c^{\prime}_{1},c_{2}}+\beta_{c^{\prime},c}\delta_{c^{\prime}_{2},c_{1}})-2(\lambda-2),$
$\frac{\delta\mathrm{Tr}(G^{\dagger}G)}{\delta\beta_{c,c^{\prime}}}=0$ (191)
$\Rightarrow\left[64(1-\lambda)^{2}+24\lambda^{2}\right]\beta_{c,c^{\prime}}\delta_{c_{2},c^{\prime}_{1}}$
$+2\lambda^{2}\sum_{c^{\prime\prime}}(\delta_{c^{\prime}_{2},c^{\prime\prime}_{1}}\beta_{c^{\prime},c^{\prime\prime}}+\delta_{c^{\prime\prime}_{2},c_{1}}\beta_{c^{\prime\prime},c})+2\delta_{c^{\prime\prime}_{2},c^{\prime}_{1}}\beta_{c^{\prime\prime},c^{\prime}}+\delta_{c_{1},c^{\prime\prime}_{1}}(1-\delta_{c_{2},c^{\prime\prime}_{2}})+2\delta_{c_{2},c^{\prime\prime}_{1}}\beta_{c,c^{\prime\prime}}(1-\delta_{c^{\prime}_{2},c^{\prime\prime}_{2}})$
$=16\lambda(1-\lambda)\delta_{c_{2},c^{\prime}_{1}}(\alpha_{c}+\alpha_{c^{\prime}}).$
Appendix G Derivation of the variational equations with COLD for the Max Cut
In this section, we provide a detailed derivation of how the COLD approach can be involved in the local variational expansion for the Max Cut problem. We apply a longitudinal field as the optimizable operator with time-dependent coefficient given by the Fourier expansion as
$f=\sum_{k=1}^{N_{k}}\gamma_{k}\sin(k\pi t/T),$ (192)
with $N_{k}$ denoting the number of considered Fourier modes. In this case, the full Hamiltonian first and second-order CD corrections read as
$H^{(2)}_{\mathrm{COLD}}=-\lambda\sum_{c}\frac{1-\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}}{2}-(1-\lambda)\sum_{j}\sigma^{x}_{j}+\sum_{j}\epsilon_{j}\sigma^{y}_{j}+\sum_{c}\alpha_{c}(\sigma^{z}_{c_{1}}\sigma^{y}_{c_{2}}+\sigma^{y}_{c_{1}}\sigma^{z}_{c_{2}})+f\sum_{j}\sigma^{z}_{j}.$ (193)
The integral action reads
$G(H^{(2)}_{\mathrm{COLD}})=-\sum_{c=1}^{M}\frac{1+8(1-\lambda)\alpha_{c}}{2}\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}+\sum_{c=1}^{M}\frac{1+2\lambda\alpha_{c}}{2}(\sigma^{x}_{c_{1}}+\sigma^{x}_{c_{2}})+4(1-\lambda)\sum_{c=1}^{M}\alpha_{c}\sigma^{y}_{c_{1}}\sigma^{y}_{c_{2}}$
$+\lambda\sum_{c,c^{\prime}}\alpha_{c}\left[(1-\delta_{c_{1},c^{\prime}_{1}})\delta_{c_{2},c^{\prime}_{2}}\sigma^{z}_{c_{1}}\sigma^{z}_{c^{\prime}_{1}}\sigma^{x}_{c_{2}}+\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}+\delta_{c_{2},c^{\prime}_{1}}\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}+\delta_{c_{1},c^{\prime}_{2}}\sigma^{z}_{c^{\prime}_{1}}\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}}\right]$
$-f\sum_{c}\alpha_{c}(\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}+\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}})-f\sum_{j}\epsilon_{j}\sigma^{x}_{j}+\partial_{\lambda}f\sum_{j}\sigma^{z}_{j}-(1-\lambda)\sum_{j}\epsilon_{j}\sigma^{z}_{j}-\frac{\lambda}{2}\sum_{c}(\epsilon_{c_{2}}\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}+\epsilon_{c_{1}}\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}})$
$=-\sum_{c=1}^{M}\frac{1+8(1-\lambda)\alpha_{c}}{2}\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}+\sum_{c=1}^{M}\left(\frac{1+2\lambda\alpha_{c}}{2}+f\frac{\epsilon_{c_{1}}}{3}\right)\sigma^{x}_{c_{1}}+\left(\frac{1+2\lambda\alpha_{c}}{2}+f\frac{\epsilon_{c_{2}}}{3}\right)\sigma^{x}_{c_{2}}+4(1-\lambda)\sum_{c=1}^{M}\alpha_{c}\sigma^{y}_{c_{1}}\sigma^{y}_{c_{2}}$
$+\partial_{\lambda}f\sum_{j}\sigma^{z}_{j}+\lambda\sum_{c,c^{\prime}}\alpha_{c}\left[(1-\delta_{c_{1},c^{\prime}_{1}})\delta_{c_{2},c^{\prime}_{2}}\sigma^{z}_{c_{1}}\sigma^{z}_{c^{\prime}_{1}}\sigma^{x}_{c_{2}}+\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}+\delta_{c_{2},c^{\prime}_{1}}\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}+\delta_{c_{1},c^{\prime}_{2}}\sigma^{z}_{c^{\prime}_{1}}\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}}\right]$
$-\sum_{c}\left((f\alpha_{c}+\lambda\epsilon_{c_{1}}/2)\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}+(f\alpha_{c}+\lambda\epsilon_{c_{2}}/2)\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}}\right)+\sum_{j}\left[\partial_{\lambda}f-(1-\lambda)\epsilon_{j}\right]\sigma^{z}_{j}.$ (194)
By similar techniques to those in the previous section, the action to be minimized becomes
$\mathrm{Tr}\left[G^{\dagger}\,G\right]=\sum_{c}\frac{[1+8(1-\lambda)\alpha_{c}]^{2}}{4}$
$+\sum_{c,c^{\prime}}\left(\frac{1+2\lambda\alpha_{c}}{2}+f\frac{\epsilon_{c_{1}}}{3}\right)\left(\frac{1+2\lambda\alpha_{c^{\prime}}}{2}+f\frac{\epsilon_{c^{\prime}_{1}}}{3}\right)\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})+\left(\frac{1+2\lambda\alpha_{c}}{2}+f\frac{\epsilon_{c_{2}}}{3}\right)\left(\frac{1+2\lambda\alpha_{c^{\prime}}}{2}+f\frac{\epsilon_{c^{\prime}_{2}}}{3}\right)(1-\delta_{c_{1},c^{\prime}_{1}})\delta_{c_{2},c^{\prime}_{2}}$
$+\sum_{c,c^{\prime}}\left[\left(\frac{1+2\lambda\alpha_{c}}{2}+f\frac{\epsilon_{c_{1}}}{3}\right)^{2}+\left(\frac{1+2\lambda\alpha_{c}}{2}+f\frac{\epsilon_{c_{2}}}{3}\right)^{2}\right]\delta_{c_{1},c^{\prime}_{1}}\delta_{c_{2},c^{\prime}_{2}}$
$+\sum_{c,c^{\prime}}\left(\frac{1+2\lambda\alpha_{c}}{2}+f\frac{\epsilon_{c_{1}}}{3}\right)\left(\frac{1+2\lambda\alpha_{c^{\prime}}}{2}+f\frac{\epsilon_{c^{\prime}_{2}}}{3}\right)\delta_{c_{1},c^{\prime}_{2}}+\left(\frac{1+2\lambda\alpha_{c}}{2}+f\frac{\epsilon_{c_{2}}}{3}\right)\left(\frac{1+2\lambda\alpha_{c^{\prime}}}{2}+f\frac{\epsilon_{c^{\prime}_{1}}}{3}\right)\delta_{c_{2},c^{\prime}_{1}}$
$+\lambda^{2}\sum_{c,c^{\prime}}(\alpha_{c}+\alpha_{c^{\prime}})^{2}\delta_{c_{1},c^{\prime}_{2}}+\lambda^{2}\sum_{c,c^{\prime}}\alpha^{2}_{c}\left[\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})+(1-\delta_{c_{1},c^{\prime}_{1}})\delta_{c_{2},c^{\prime}_{2}}\right]+16(1-\lambda)^{2}\sum_{c}\alpha^{2}_{c}+\sum_{j}\left[\partial_{\lambda}f-(1-\lambda)\epsilon_{j}\right]^{2}$
$+\sum_{c}(f\alpha_{c}+\lambda\epsilon_{c_{1}}/2)^{2}+(f\alpha_{c}+\lambda\epsilon_{c_{2}}/2)^{2}.$ (195)
The corresponding minimization equations are obtained by taking the variations with respect to both $\epsilon_{j}$ and $\alpha_{c}$
$\frac{\delta\mathrm{Tr}\left[G^{\dagger}\,G\right]}{\delta\epsilon_{n}}=$ (196)
$\frac{2}{3}f\sum_{c,c(n)}\left(\frac{1+2\lambda\alpha_{c}}{2}+f\frac{\epsilon_{c_{1}}}{3}\right)\delta_{n,c_{1}}(1-\delta_{c_{2}(n),c_{2}})+\left(\frac{1+2\lambda\alpha_{c}}{2}+f\frac{\epsilon_{c_{2}}}{3}\right)(1-\delta_{c_{1}(n),c_{1}})\delta_{n,c_{2}}$
$+\sum_{c(n)}\frac{4}{3}f\left(\frac{1+2\lambda\alpha_{c(n)}}{2}+f\frac{\epsilon_{n}}{3}\right)+\frac{2}{3}f\sum_{c,c(n)}\left(\frac{1+2\lambda\alpha_{c}}{2}+f\frac{\epsilon_{c_{2}}}{3}\right)\delta_{n,c_{2}}+\left(\frac{1+2\lambda\alpha_{c}}{2}+f\frac{\epsilon_{c_{1}}}{3}\right)\delta_{n,c_{1}}$
$+\lambda\sum_{c(n)}(f\alpha_{c(n)}+\lambda\epsilon_{n}/2)-2(1-\lambda)(\partial_{\lambda}f-(1-\lambda)\epsilon_{n})=0,$
where $c(n)$ denotes the edge in which the $n$-th spin appears and the summation over holds for every edge in which $n=c_{1}$ or $n=c_{2}$. Rearranging it so that all $\epsilon_{n}$-s appear on the l.h.s., one obtains
$\frac{2}{9}f^{2}\sum_{c,c(n)}\epsilon_{c_{1}}(\delta_{n,c_{2}}+\delta_{n,c_{1}}(1-\delta_{c_{2}(n),c_{2}}))+\epsilon_{c_{2}}(\delta_{n,c_{1}}+(1-\delta_{c_{1}(n),c_{1}})\delta_{n,c_{2}})+\frac{2}{3}\left[2f^{2}+\lambda^{2}+6(1-\lambda)^{2}\right]\epsilon_{n}$ (197)
$+\frac{2}{3}f\lambda\sum_{c,c(n)}\alpha_{c}\left[(1-\delta_{c_{1}(n),c_{1}})\delta_{n,c_{2}}+\delta_{n,c_{1}}(1-\delta_{c_{2}(n),c_{2}})+\delta_{n,c_{1}}+\delta_{n,c_{2}}\right]+\frac{7}{3}\lambda f\sum_{c(n)}\alpha_{c(n)}$
$=2\partial_{\lambda}f(1-\lambda)-(\frac{4}{3}fL+4+2fL).$
The minimization with respect to $\alpha_{c}$ reads as
$\frac{\delta\mathrm{Tr}(G^{\dagger}G)}{\delta\alpha_{c}}$ $=64(1-\lambda)^{2}\alpha_{c}+2\lambda^{2}\sum_{c^{\prime}}\alpha_{c}\left[\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})+(1-\delta_{c_{1},c^{\prime}_{1}})\delta_{c_{2},c^{\prime}_{2}}+2\delta_{c_{1},c^{\prime}_{2}}+2\delta_{c_{2},c^{\prime}_{1}}\right]$ (198)
$+2\lambda\sum_{c^{\prime}}\left(\frac{1+2\lambda\alpha_{c^{\prime}}}{2}+f\frac{\epsilon_{c^{\prime}_{1}}}{3}\right)\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})+\left(\frac{1+2\lambda\alpha_{c^{\prime}}}{2}+f\frac{\epsilon_{c^{\prime}_{2}}}{3}\right)\delta_{c_{2},c^{\prime}_{2}}(1-\delta_{c_{1},c^{\prime}_{1}})$
$+2\lambda\sum_{c^{\prime}}\left(\frac{1+2\lambda\alpha_{c^{\prime}}}{2}+\frac{\epsilon_{c^{\prime}_{2}}}{3}\right)\delta_{c_{1},c^{\prime}_{2}}+\left(\frac{1+2\lambda\alpha_{c^{\prime}}}{2}+\frac{\epsilon_{c^{\prime}_{1}}}{3}\right)\delta_{c^{\prime}_{1},c_{2}}$
$+2\lambda\left[\left(\frac{1+2\lambda\alpha_{c}}{2}+\frac{\epsilon_{c_{1}}}{3}\right)+\left(\frac{1+2\lambda\alpha_{c}}{2}+\frac{\epsilon_{c_{2}}}{3}\right)\right]+2f(2\alpha_{c}+\lambda(\epsilon_{c_{1}}+\epsilon_{c_{2}})/2)=-4(1-\lambda).$
Splitting out also the constants, one obtains
$\frac{\delta\mathrm{Tr}(G^{\dagger}G)}{\delta\alpha_{c}}$ $=\left[64(1-\lambda)^{2}+12\lambda^{2}+4f\lambda\right]\alpha_{c}+2\lambda^{2}\sum_{c^{\prime}}\alpha_{c^{\prime}}\left[\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})+(1-\delta_{c_{1},c^{\prime}_{1}})\delta_{c_{2},c^{\prime}_{2}}+2\delta_{c_{1},c^{\prime}_{2}}+2\delta_{c_{2},c^{\prime}_{1}}\right]$
$=$ $-2\lambda\sum_{c^{\prime}}\left(\frac{1}{2}+f\frac{\epsilon_{c^{\prime}_{1}}}{3}\right)\delta_{c_{1},c^{\prime}_{1}}+\left(\frac{1}{2}+f\frac{\epsilon_{c^{\prime}_{2}}}{3}\right)\delta_{c_{2},c^{\prime}_{2}}-2\lambda\sum_{c^{\prime}}\left(\frac{1}{2}+\frac{\epsilon_{c^{\prime}_{2}}}{3}\right)\delta_{c_{1},c^{\prime}_{2}}+\left(\frac{1}{2}+\frac{\epsilon_{c^{\prime}_{1}}}{3}\right)\delta_{c^{\prime}_{1},c_{2}}$
$-f\lambda(\epsilon_{c_{1}}+\epsilon_{c_{2}})-4(1-\lambda)$
$=$ $-\frac{2}{3}\lambda f\sum_{c^{\prime}}\epsilon_{c^{\prime}_{1}}(\delta_{c_{1},c^{\prime}_{1}}+\delta_{c_{2},c^{\prime}_{1}})+\epsilon_{c^{\prime}_{2}}(\delta_{c_{2},c^{\prime}_{2}}+\delta_{c_{1},c^{\prime}_{2}})-f\lambda(\epsilon_{c_{1}}+\epsilon_{c_{2}})-2(\lambda-2).$ (199)
Appendix H First order CD ansatz for the $3$-regular $3$-XORSAT
In this section, we show the steps for the derivation of the minimization equation for the $\alpha_{c}$ coefficients in the XORSAT problem. The first-order CD term already contains three spin interactions written as
$H^{(1)}_{1}=\sum_{c}\,\alpha_{c}\left(\sigma^{y}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{c_{3}}+\sigma^{z}_{c_{1}}\sigma^{y}_{c_{2}}\sigma^{z}_{c_{3}}+\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{y}_{c_{3}}\right).$ (200)
with the ordering $c_{1}<c_{2}<c_{3}$. The argument of the action thus takes the form of
$G(H^{(1)}_{1})=-\sum_{c}\frac{1+12(1-\lambda)}{2}\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{c_{3}}+\sum_{c}\frac{1+3\lambda\alpha_{c}}{3}(\sigma^{x}_{c_{1}}+\sigma^{x}_{c_{2}}+\sigma^{x}_{c_{3}})$ (201)
$+\lambda\sum_{c,c^{\prime}}\alpha_{c}\left(\delta_{c_{1},c^{\prime}_{2}}\delta_{c_{2},c^{\prime}_{3}}\sigma^{z}_{c^{\prime}_{1}}(\sigma^{x}_{c_{1}}+\sigma^{x}_{c_{2}})\sigma^{z}_{c_{3}}+\delta_{c_{2},c^{\prime}_{1}}\delta_{c_{3},c^{\prime}_{2}}\sigma^{z}_{c_{1}}(\sigma^{x}_{c_{2}}+\sigma^{x}_{c_{3}})\sigma^{z}_{c^{\prime}_{3}}\right)$
$+\lambda\sum_{c,c^{\prime}}\alpha_{c}\left(\delta_{c_{1},c^{\prime}_{1}}\delta_{c_{2},c^{\prime}_{2}}(1-\delta_{c_{3},c^{\prime}_{3}})(\sigma^{x}_{c_{1}}+\sigma^{x}_{c_{2}})\sigma^{z}_{c_{3}}\sigma^{z}_{c^{\prime}_{3}}+\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})\delta_{c_{3},c^{\prime}_{3}}(\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}+\sigma^{z}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}\sigma^{x}_{c_{3}})\right)$
$+\lambda\sum_{c}\alpha_{c}\left((1-\delta_{c_{1},c^{\prime}_{1}})\delta_{c_{2},c^{\prime}_{2}}\delta_{c_{3},c^{\prime}_{3}}\sigma^{z}_{c_{1}}\sigma^{z}_{c^{\prime}_{1}}(\sigma^{x}_{c_{2}}+\sigma^{x}_{c_{3}})\right)$
$+\lambda\sum_{c,c^{\prime}}\alpha_{c}\left(\delta_{c_{1},c^{\prime}_{3}}\sigma^{z}_{c^{\prime}_{1}}\sigma^{z}_{c^{\prime}_{2}}\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{c_{3}}+\delta_{c_{3},c^{\prime}_{1}}\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{x}_{c_{3}}\sigma^{z}_{c^{\prime}_{2}}\sigma^{z}_{c^{\prime}_{3}}+(1-\delta_{c_{1},c^{\prime}_{1}})\delta_{c_{2},c^{\prime}_{2}}(1-\delta_{c_{3},c^{\prime}_{3}})\sigma^{z}_{c_{1}}\sigma^{z}_{c^{\prime}_{1}}\sigma^{x}_{c_{2}}\sigma^{z}_{c_{3}}\sigma^{z}_{c^{\prime}_{3}}\right)$
$+\lambda\sum_{c,c^{\prime}}\alpha_{c}\left(\delta_{c_{1},c^{\prime}_{2}}(1-\delta_{c_{2},c^{\prime}_{3}})\sigma^{z}_{c^{\prime}_{1}}\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{c^{\prime}_{3}}\sigma^{z}_{c_{3}}+\delta_{c_{2},c^{\prime}_{1}}(1-\delta_{c_{3},c^{\prime}_{2}})\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}\sigma^{z}_{c_{3}}\sigma^{z}_{c^{\prime}_{3}}\right)$
$+\lambda\sum_{c,c^{\prime}}\alpha_{c}\left((1-\delta_{c_{1},c^{\prime}_{2}})\delta_{c_{2},c^{\prime}_{3}}\sigma^{z}_{c^{\prime}_{1}}\sigma^{z}_{c_{1}}\sigma^{z}_{c^{\prime}_{2}}\sigma^{x}_{c_{2}}\sigma^{z}_{c_{3}}+(1-\delta_{c_{2},c^{\prime}_{1}})\delta_{c_{3},c^{\prime}_{2}}\sigma^{z}_{c_{1}}\sigma^{z}_{c^{\prime}_{1}}\sigma^{z}_{c_{2}}\sigma^{x}_{c_{3}}\sigma^{z}_{c^{\prime}_{3}}\right)$
$+\lambda\sum_{c,c^{\prime}}\alpha_{c}\left(\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})(1-\delta_{c_{3},c^{\prime}_{3}})\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}\sigma^{z}_{c_{3}}\sigma^{z}_{c^{\prime}_{3}}+(1-\delta_{c_{1},c^{\prime}_{1}})(1-\delta_{c_{2},c^{\prime}_{2}})\delta_{c_{3},c^{\prime}_{3}}\sigma^{z}_{c_{1}}\sigma^{z}_{c^{\prime}_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}\sigma^{x}_{c_{3}}\right)$
$+2(1-\lambda)\sum_{c}\alpha_{c}\left(\sigma^{y}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{z}_{c_{3}}+\sigma^{y}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{x}_{c_{3}}+\sigma^{x}_{c_{1}}\sigma^{y}_{c_{2}}\sigma^{z}_{c_{3}}+\sigma^{z}_{c_{1}}\sigma^{y}_{c_{2}}\sigma^{x}_{c_{3}}+\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{y}_{c_{3}}+\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{y}_{c_{3}}\right).$
Note that the last five-body term in the $5$-th line is orthogonal to the first two due to the underlying graph structure. This is the consequence of the rule that every spin can only appear in three interactions, while matching the indices of $(1-\delta_{c_{1},c^{\prime}_{1}})\delta_{c_{2},c^{\prime}_{2}}(1-\delta_{c_{3},c^{\prime}_{3}})\sigma^{z}_{c_{1}}\sigma^{z}_{c^{\prime}_{1}}\sigma^{x}_{c_{2}}\sigma^{z}_{c_{3}}\sigma^{z}_{c^{\prime}_{3}}$ and of $\delta_{c_{1},c^{\prime}_{3}}\sigma^{z}_{c^{\prime}_{1}}\sigma^{z}_{c^{\prime}_{2}}\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{c_{3}}$ in the trace norm requires the matching of the third spin by four different interactions. Similar consideration imply that also the last term in the third line, $\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})\delta_{c_{3},c^{\prime}_{3}}(\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}+\sigma^{z}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}\sigma^{x}_{c_{3}})$ is orthogonal to the first two, $\delta_{c_{1},c^{\prime}_{1}}\delta_{c_{2},c^{\prime}_{2}}(1-\delta_{c_{3},c^{\prime}_{3}})(\sigma^{x}_{c_{1}}+\sigma^{x}_{c_{2}})\sigma^{z}_{c_{3}}\sigma^{z}_{c^{\prime}_{3}}$, as matching the proper indices in the trace norm would require that the $c_{1}$ spin appears in four different interactions. As a result, the trace of the square reads
$\mathrm{Tr}\left[G^{\dagger}G\right]=\sum_{c}\frac{(1+12(1-\lambda)\alpha_{c})^{2}}{4}$ (202)
$+\frac{1}{3}\sum_{c,c^{\prime}}\left(3\lambda^{2}\alpha_{c}\alpha_{c^{\prime}}+\lambda(\alpha_{c}+\alpha_{c^{\prime}})\right)\left[\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})(1-\delta_{c_{3},c^{\prime}_{3}})+\delta_{c_{1},c^{\prime}_{1}}\delta_{c_{2},c^{\prime}_{2}}(1-\delta_{c_{3},c^{\prime}_{3}})+\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})\delta_{c_{3},c^{\prime}_{3}}\right]$
$+\frac{1}{3}\sum_{c,c^{\prime}}\left(3\lambda^{2}\alpha_{c}\alpha_{c^{\prime}}+\lambda(\alpha_{c}+\alpha_{c^{\prime}})\right)\left[(1-\delta_{c_{1},c^{\prime}_{1}})\delta_{c_{2},c^{\prime}_{2}}(1-\delta_{c_{3},c^{\prime}_{3}})+(1-\delta_{c_{1},c^{\prime}_{1}})\delta_{c_{2},c^{\prime}_{2}}\delta_{c_{3},c^{\prime}_{3}}+(1-\delta_{c_{1},c^{\prime}_{1}})(1-\delta_{c_{2},c^{\prime}_{2}})\delta_{c_{3},c^{\prime}_{3}}\right]$
$+\frac{1}{3}\sum_{c,c^{\prime}}\left(3\lambda^{2}\alpha_{c}\alpha_{c^{\prime}}+\lambda(\alpha_{c}+\alpha_{c^{\prime}})\right)\left[\delta_{c_{1}c^{\prime}_{2}}+\delta_{c_{1}c^{\prime}_{3}}+\delta_{c_{2}c^{\prime}_{3}}+\delta_{c_{2}c^{\prime}_{1}}+\delta_{c_{3}c^{\prime}_{1}}+\delta_{c_{3}c^{\prime}_{2}}+3\delta_{c_{1},c^{\prime}_{1}}\delta_{c_{2},c^{\prime}_{2}}\delta_{c_{3},c^{\prime}_{3}}\right]$
$+2\lambda^{2}\sum_{c,c^{\prime}}(\alpha_{c}+\alpha_{c^{\prime}})^{2}\delta_{c_{1},c^{\prime}_{2}}\delta_{c_{2},c^{\prime}_{3}}+2\lambda^{2}\sum_{c,c^{\prime}}\alpha^{2}_{c}\left(\delta_{c_{1},c^{\prime}_{1}}\delta_{c_{2},c^{\prime}_{2}}(1-\delta_{c_{3},c^{\prime}_{3}})+\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})\delta_{c_{3},c^{\prime}_{3}}+(1-\delta_{c_{1},c^{\prime}_{1}})\delta_{c_{2},c^{\prime}_{2}}\delta_{c_{3},c^{\prime}_{3}}\right)$
$+\lambda^{2}\sum_{c,c^{\prime}}\left(\alpha_{c}+\alpha_{c^{\prime}}\right)^{2}\delta_{c_{1},c^{\prime}_{3}}+\lambda^{2}\sum_{c,c^{\prime}}(\alpha_{c}+\alpha_{c^{\prime}})^{2}\delta_{c_{1},c^{\prime}_{2}}(1-\delta_{c_{2},c^{\prime}_{3}})+\lambda^{2}\sum_{c,c^{\prime}}(\alpha_{c}+\alpha_{c^{\prime}})^{2}\delta_{c_{2},c^{\prime}_{3}}(1-\delta_{c_{1},c^{\prime}_{2}})$
$+\lambda^{2}\sum_{c,c^{\prime}}\alpha^{2}_{c}\left[\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})(1-\delta_{c_{3},c^{\prime}_{3}})+(1-\delta_{c_{1},c^{\prime}_{1}})(1-\delta_{c_{2},c^{\prime}_{2}})\delta_{c_{3},c^{\prime}_{3}}+(1-\delta_{c_{1},c^{\prime}_{1}})\delta_{c_{2},c^{\prime}_{2}}(1-\delta_{c_{3},c^{\prime}_{3}})\right]$
+12($+12(1-\lambda)^{2}\sum_{c}\alpha^{2}_{c},$
where we have enlisted in the first three lines all the terms that distinguish between the $c\neq c^{\prime}$ and $c=c^{\prime}$ cases.
Although these expressions might appear slightly complicated at first glance, the terms $\alpha_{c}\sum_{c^{\prime}}\delta_{\dots}$ simplify to a constant for all $c$ as it counts how many times the given edges of an interaction $c$ appear in other interactions. Consequently, it gives a factor of $9$. Additional differentiations also induce an additional factor of $2$. The coupled $\alpha_{c^{\prime}}$ follow the same rule, those $\alpha_{c}$ and $\alpha_{c^{\prime}}$ are connected that share at least one spin.
$18\lambda^{2}\alpha_{c}+96(1-\lambda)^{2}+2\lambda^{2}\sum_{c^{\prime}}\alpha_{c^{\prime}}\left[3\delta_{c_{1},c^{\prime}_{2}}+3\delta_{c_{2},c^{\prime}_{1}}+2\delta_{c_{1},c^{\prime}_{3}}+2\delta_{c_{3},c^{\prime}_{1}}\right]+2\lambda^{2}\sum_{c^{\prime}\neq c}\alpha_{c^{\prime}}\left[\delta_{c_{1},c^{\prime}_{1}}+\delta_{c_{2},c^{\prime}_{2}}+\delta_{c_{3},c^{\prime}_{3}}\right]$
$=-6(1-\lambda)-\frac{2}{3}\lambda\sum_{c^{\prime}\neq c}\left[\delta_{c_{1},c^{\prime}_{1}}+\delta_{c_{2},c^{\prime}_{2}}+\delta_{c_{3},c^{\prime}_{3}}+\delta_{c_{1}c^{\prime}_{2}}+\delta_{c_{1}c^{\prime}_{3}}+\delta_{c_{2}c^{\prime}_{3}}+\delta_{c_{2}c^{\prime}_{1}}+\delta_{c_{3}c^{\prime}_{1}}+\delta_{c_{3}c^{\prime}_{2}}\right].$ (203)
As a final remark, the last term counts how many interactions have at least one shared spin with $c$. This final result is reported in the main text via simplified notation. The constraint $c\neq c^{\prime}$ provides a more compact notation and ensures that only non-equal $\alpha_{c^{\prime}}$-s are considered.